GTA Technical Reports

The technical reports presented here are available via FTP at: ftp://ftp.gta.ufrj.br/pub/gta/TechReports

Last update: Feb, 14th 2017

Technical Reports — 2017

GTA-17-01 Doctor Science Thesis Author - Diogo Menezes Ferrazani Mattos Title - “Segurança em Redes Definidas por Software: Autenticação, Controle de Acesso e Consistência com Plano de Controle Eficientemente Distribuído" Advisors - Otto Carlos Muniz Bandeira Duarte and Guy Pujolle COPPE/PEE/UFRJ - January 2017 English - A4 size - 119 p. File: Diogo17.pdf

GTA-17-02 Undergraduate Final Project Author - Felipe da Rocha Moraes Title - "Sistema de Detecção de Passageiros Utilizando Técnicas de Tratamento de Imagem" Advisor - Miguel Elias Mitre Campista DEL/POLI/UFRJ - Fevereiro 2017 Portuguese - A4 size - 63 p. - File: FeMoraes17.pdf

GTA-17-03 Andreoni Lopez, M., Lobato, A. G. P., Mattos, D. M. F., Alvarenga, I. D., Duarte, O. C. M. B., Pujolle, G. - "Um Algoritmo Não Supervisionado e Rápido para Seleção de Características em Classificação de Tráfego", in XXXV Simpósio Brasileiro de Redes de Computadores e Sistemas Distribud́os - SBRC’2017, Belém- Pará, PA, Brazil, May 2017. Portuguese, A4 size, 14 p., file: ALM17.pdf

GTA-17-04 Mattos, D. M. F., Duarte, O. C. M. B., Pujolle, G. - "Um Protocolo Simples e Eficiente para Atualização Consistente de Políticas em Redes Definidas por Software com Controle Distribuído", in XXXV Simpósio Brasileiro de Redes de Computadores e Sistemas Distribud́os - SBRC’2017, Belém- Pará, PA, Brazil, May 2017. Portuguese, A4 size, 14 p., file: MDP17.pdf

GTA-17-05 Andreoni Lopez, M., Silva, R. S., Alvarenga, I. D., Mattos, D. M. F., Duarte, O. C. M. B. - "Coleta e Caracterização de um Conjunto de Dados de Tráfego Real de Redes de Acesso em Banda Larga", in XXII Workshop de Gerência e Operação de Redes e Serviços (WGRS’2017) - SBRC’2017, Belém- Pará, PA, Brazil, May 2017. Portuguese, A4 size, 14 p., file: ASA17.pdf

Technical Reports — 2016

GTA-16-01 Doctor Science Thesis Author - Fabio Antonio Vieira Pinto Title - “Um Modelo para Dimensionamento de Postos de Recarga Rápida para Veículos Elétricos baseado no Perfil de Carga das Baterias de Íons de Lítio” Advisors - Luís Henrique Maciel Kosmalski Cos COPPE/PEE/UFRJ - January 2016 Portuguese, A4 size, 117 p. - File: Fabio16.pdf

GTA-16-02 Mattos, D. M. F., Duarte, O. C. M. B., and Pujolle, G.

“A Resilient Distributed Controller for Software Defined Networking" in IEEE International Conference on Communications - ICC 2016, pp. 1-6, May 2016. English, A4 size, 6 p. - File: MDP16.pdf

GTA-16-03 Master Science Thesis Author - Fernando Fraga Machado Title - “Análise da Estabilidade do Ranqueamento de Grau em Redes Livre de Escala” Advisor - Miguel Elias Mitre Campista COPPE/PEE/UFRJ - January 2016 Portuguese - A4 size - 53 p. - File: Fernando16.pdf

GTA-16-04 Mattos, D. M. F., and Duarte, O. C. M. B. - “AuthFlow: authentication and access control mechanism for software defined networking”, to be published in Annals of Telecommunications, Springer, ISSN 0003-4347, DOI 10.1007/s12243-016-0505-z, 2016. English, A4 size, 9 p., file: MaDu16.pdf

GTA-16-05 Andreoni Lopez, M., Mattos, D. M. F., and Duarte, O. C. M. B. - “An elastic intrusion detection system for software networks”, to be published in Annals of Telecommunications, Springer, ISSN 0003-4347, DOI 10.1007/s12243-016-0506-y, 2016. English, A4 size, 11 p., file: AMD16.pdf

GTA-16-06 Mattos, D. M. F., Duarte, O. C. M. B., and Pujolle, G. - “Reverse Update: A Consistent Policy Update Scheme for Software Defined Networking”, to be published in IEEE Communications Letters, vol.PP, no.99, pp.1-1, IEEE, ISSN 1089-7798, DOI 10.1109/LCOMM.2016.2546240, 2016. English, Letter size, 4 p., file: MDP16a.pdf

GTA-16-07 Undergraduate Final Project Author - Antonio Gonzalez Pastana Lobato Title - “Um Sistema Acurado de Detec? de Amea? em Tempo Real por Processamento de Fluxos” Advisors - Otto Carlos Muniz Bandeira Duarte and Martin Andreoni Lopez DEL/POLI/UFRJ - April 2016 Portuguese - A4 size - 69 p. - File: Antonio16.pdf

GTA-16-08 Lobato, A. G. P., Andreoni Lopez, M. and Duarte, O. C. M. B. - “An Accurate Threat Detection System through Real-Time Stream Processing”, Technical Report, Electrical Engineering Program, COPPE/UFRJ, April 2016. English, A4 size, 6 p., file: LAD16.pdf

GTA-16-09 Baron, B., Campista, M. E. M., Spathis, P, Costa, L. H. M. K., Amorim, M. D., Duarte, O. C. M. B., Pujolle, G., Viniotis, Y. - “Virtualizing vehicular node resources: Feasibility study of virtual machine migration”, in Vehicular Communications, Elsevier, ISSN 2214-2096, DOI 10.1016/j.vehcom.2016.04.001, vol. 4, pp. 39-46, April 2016 English, A4 size, 8 p., file: BCS16.pdf GTA-16-10 Lobato, A. G. P., Andreoni Lopez, M. and Duarte, O. C. M. B. -

“Um Sistema Acurado de Detec? de Amea? em Tempo Real por Processamento de Fluxos”, in XXXIV Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?s - SBRC’2016, Salvador- Bahia, BA, Brazil, May 2016. Portuguese, A4 size, 14 p., file: LAD16b.pdf

GTA-16-11 Andreoni Lopez, M., Lobato, A. G. P., and Duarte, O. C. M. B. - “Monitoramento de Tr?go e Detec? de Amea? por Sistemas Distribu?s de Processamento de Fluxos: uma An?se de Desempenho”, in XXI Workshop de Ger?ia e Opera? de Redes e Servi? (WGRS 2016) - SBRC’2016, Salvador-Bahia, BA, Brazil, May 2016. Portuguese, A4 size, 14 p., file: ALD16.pdf

GTA-16-12 Torres, J.V., Boutaba, R., and Duarte - “An Autonomous and Efficient Controller-based Routing Scheme for Networking Named-Data Mobility”, Technical Report (submitted to Computer Communications 2016), Electrical Engineering Program, COPPE/UFRJ, April 2016. English, A4 size, 12 p., file: TBD16.pdf

GTA-16-13 Torres, J.V., Boutaba, R., and Duarte - “Evaluating CRoS-NDN: A comparative performance analysis of the Controller-based Routing Scheme for Named-Data Networking”, Technical Report (submitted to IEEE/ACM Transactions on Networking 2016), Electrical Engineering Program, COPPE/UFRJ, May 2016. English, A4 size, 14 p., file: TBD16a.pdf

GTA-16-14 Bari, Md. F., Chowdhury, S. R., Ahmed, R., Boutaba, R., and Duarte, O. C. M. B. - “Orchestrating Virtualized Network Functions”, to appear on IEEE Transactions on Network and Service Management, ISSN: 1932-4537, May 2016. English, Letter size, 14 p., file: BCA16.pdf

GTA-16-15 Pinto Neto, João B., Gomes, Lucas C., Castanho, Eduardo M., Campista. Miguel Elias M. , and Costa, Luís Henrique M. K. - “Um Algoritmo para Cálculo de Distância para Prevenção de Colisão Dianteira em Redes Veiculares”, in XXI Workshop de Gerência e Operação de Redes e Serviços - WGRS’2016, Salvador, BA, Brazil, pp. 59-72, May 2016. Portuguese, A4 size, 14 p., file: NGC16.pdf

GTA-16-16 Pinto Neto, João B. and Costa, Luís Henrique M. K. - “Enderea̧mento em redes ad hoc móveis: conceitos, protocolos e tendências", in Revista Brasileira de Computação Aplicada, vol. 8, no. 1, pp. 26-39, June, 2016, ISSN: 2176-6649. Portuguese, A4 size, 11 p., file: jb-rbca2016.pdf

GTA-16-17 Doctor Science Thesis Author - Alyson de Jesus dos Santos Title - “Caracterizaçã o de Redes Tolerantes a Atrasos e Desconexõ es na Bacia Hidrográfica Amazônica”

Advisors - Luís Henrique Maciel Kosmalski Costa and Pedro Braconnot Velloso COPPE/PEE/UFRJ - January 2016 Portuguese, A4 size, 89 p. - File: Alyson16.pdf

GTA-16-18 Medeiros, D. S. V., Campista, M. E. M., Mitton, N., Amorim, M. D., and Pujolle, G. - “Intermediação por Espalhamento: Caminhos Quase Mais Curtos Também Importam", in XXXIV Simpósio Brasileiro de Redes de Computadores e Sistemas Distribuídos - SBRC’2016, Salvador, BA, Brazil, pp. 967-980, May 2016. Portuguese, A4 size, 14 p., file: MCM16.pdf

GTA-16-19 Sciammarella, T., Couto, R. S., Rubinstein, M. G., Campista, M. E. M., and Costa, L. H. M. K. - “Uma Análise do Tráfego de Controle de uma Nuvem IaaS Distribuída", in XXXIV Simpósio Brasileiro de Redes de Computadores e Sistemas Distribuídos - SBRC’2016, Salvador, BA, Brazil, pp. 674-687, May 2016. Portuguese, A4 size, 14 p., file: SCR16.pdf

GTA-16-20 Machado, F. F., Figueiredo, D. R., and Campista, M. E. M. - “Análise da Estabilidade de Ranqueamento de Grau na Rede de Sistemas Autônomos da Internet", in XXXIV Simpósio Brasileiro de Redes de Computadores e Sistemas Distribuídos - SBRC’2016, Salvador, BA, Brazil, pp. 396-409, May 2016. Portuguese, A4 size, 14 p., file: MFC16.pdf

GTA-16-21 Silvério, A. J., Campista, M. E. M., and Costa, L. H. M. K. - “Roteamento por Segmentos: Conceitos, Desafios e Aplicações Práticas", in Minicursos do XXXIV Simpósio Brasileiro de Redes de Computadores e Sistemas Distribuídos - SBRC’2016, Salvador, BA, Brazil, Chapter 5, pp. 210-265, ISSN 2177-4978, May 2016. Portuguese, A4 size, 56 p., file: SCC16.pdf

GTA-16-22 Doctor Science Thesis Author - Marcus de Lima Braga Title - “Análise e Planejamento de Rotas em Redes Viárias” Advisors - Luís Henrique Maciel Kosmalski Costa and Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - January 2016 Portuguese, A4 size, 126 p. - File: Marcus16.pdf

GTA-16-23 Dias, D. S., Costa, L. H. M. K. - “Análise da Capacidade de Dados de uma Rede de Ônibus Urbanos”, in XXXIV Simpósio Brasileiro de Telecomunicações e Processamento de Sinais - SBrT 2016, Santarém, PA, Brazil, August 2016. Portuguese, A4 size, 5 p. - File: DC16.pdf

GTA-16-24 Andreoni Lopez, M., Lobato, A. G. P., and Duarte, O. C. M. B.- “Performance Comparison of Open-Source Stream Processing Platforms”, in IEEE Global Communications Conference - GLOBECOM’2016,Washington, DC USA, December 2016. English, A4 size, 6 p- File: ALD16b.pdf

GTA-16-25 Master Science Thesis Author - Eduardo Gomes de Oliveira Title - "Uma Biblioteca de Auxílio ao Desenvolvimento de Aplicações de Redes Móveis Advisor - Miguel Elias M. Campista and Luís Henrique M. K. Costa COPPE/PEE/UFRJ - Março 2016 Portuguese - A4 size - 60 p. - File: Eduardo16.pdf GTA-16-26 Pinto Neto, J. B., Gomes, L. C., Castanho, E. M., Campista, M. E. M., and

Costa, L. H. M. K. - “Um Sistema Redundante de Frenagem Segura para Trens de Levitação Magnética”, in XXXIV Simpósio Brasileiro de Telecomunicações e Processamento de Sinais - SBRT’2016, Santarém, PA, Brazil, Aug 2016. Portuguese, A4 size, 5 p., file: NGC16a.pdf

GTA-16-27 Pinto Neto, J. B., Gomes, L. C., Castanho, E. M., Campista, M. E. M., and Costa, L. H. M. K. - “An Error Correction Algorithm for Forward Collision Warning Applications”, in 19th International Conference on Inteligent Transportation Systems - IEEE ITSC’2016, pp. 1926-1931, Rio de Janeiro, RJ, Brazil, Nov 2016. English, A4 size, 6 p., file: NGC16b.pdf

GTA-16-28 Sciammarella, T., Couto, R. S., Rubinstein, M. G., Campista, M. E. M., and Costa, L. H. M. K. - “Analysis of Control Traffic in a Geo-distributed Collaborative Cloud”, in 5th IEEE International Conference on Cloud Networking - IEEE CloudNet’2016, Pisa, Itália, pp. 1-6, October 2016. English, A4 size, 6 p., file: SCR16a.pdf

GTA-16-29 Silva, V. B. C., Campista, M. E. M., and Costa, L. H. M. K. - “TraC: A Trajectoryaware Content distribution strategy for vehicular networks”, in Vehicular Communications, Elsevier, ISSN 2214-2096, DOI 10.1016/j.vehcom.2016.09.005, vol. 5, pp. 18-34, July 2016 English, A4 size, 57 p., file: SCC16a.pdf

GTA-16-30 Mattos, D. M. F., Duarte, O. C. M. B., and Pujolle, G. - “Profiling Software Defined Networks for Dynamic Distributed-Controller Provisioning”, in 7th International Conference Network of the Future IEEE- NoF’2016, B??s-RJ, Brazil, November 2016. English, A4 size, 5 p., file: MDP16b.pdf

GTA-16-31 Andreoni Lopez, M., Mattos, D. M. F., and Duarte, O. C. M. B. - “Evaluating Allocation Heuristics for an Efficient Virtual Network Function Chaining”, in 7th International Conference Network of the Future IEEE - NoF’2016, B??s-RJ, Brazil, November 2016. English, A4 size, 5 p., file: AMD16b.pdf

GTA-16-33 Medeiros, D. S. V., Campista, M. E. M., Mitton, N., Amorim, M. D., and Pujolle, G. - “Weighted Betweenness for Multipath Networks”, in Global Information Infrastructure and Networking Symposium 2016 - GIIS’2016, Porto, Portugal, pp. 1-6, October 2016. English, A4 size, 6 p., file: MCM16b.pdf

GTA-16-34 Andreoni Lopez, M., Lobato, A. G. P., Duarte, O. C. M. B., and Pujolle, G. - “Design and Performance Evaluation of a Virtualized Network Function for Real-Time Threat Detection using Stream Processing”, COPPE/UFRJ, November 2016. English, A4 size, 6 p., file: ALDP16.pdf

GTA-16-35 Doctor of Science Thesis Author - Jo?Vitor Torres Title - "An Autonomous and Efficient Controller-based Routing Scheme for Networking Named-data Mobility" Advisor -Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - June 2016 English - A4 size - 79 p. JoaoVitor16.pdf GTA-16-36 Santos, A. J., Costa, L. H. M. K., Braga, M. L., Velloso, P. B., and Ghamri-

Doudane, Y. - "Characterization of a Delay and Disruption Tolerant Network in the Amazon Basin", in Vehicular Communications, Elsevier, ISSN 2214-2096, DOI 10.1016/j.vehcom.2016.09.002, vol. 5, pp. 35-43, July 2016 English, A4 size, 9 p., file: SCB16.pdf

GTA-16-37 Braga, M. L., Santos, A. J., Pedroza, A. C. P., Costa, L. H. M. K., Amorim, M. D., and Ghamri-Doudane, Y. - "Anytime Route Planning with Constrained Devices", in Computers and Electrical Engineering, Elsevier, ISSN 0045-7906, DOI 10.1016/j.compeleceng.2016.07.011, vol. 52, pp. 53-67, August 2016 English, A4 size, 15 p., file: BSP16.pdf

GTA-16-38 Undergraduate Final Project Author - Hugo de Freitas Siqueira Sadok Menna Barreto Title - “Compartilhamento de Recursos em Nuvens Colaborativas” Advisors - Miguel Elias Mitre Campista and Lu?Henrique Maciel Kosmalski Costa DEL/POLI/UFRJ - September 2016 Portuguese - A4 size - 62 p. - File: Sadok16.pdf

GTA-16-39 Undergraduate Final Project Author - Pedro Henrique Cruz Caminha Title - “Implementa? de Autentica? Federada em uma Nuvem Comunit?a Geodistribu?” Advisors - Lu?Henrique Maciel Kosmalski Costa and Rodrigo de Souza Couto DEL/POLI/UFRJ - April 2016 Portuguese - A4 size - 72 p. - File: Cruz16.pdf

GTA-16-40 Cruz, P., Pinto Neto, João B., Campista, M. E. M., Costa, L. H. M. K. "On the Accuracy of Data Sensing In the Presence of Mobility", in 7th International Conference Network of the Future - NoF’2016, Buzios-RJ, Brazil, November 2016. English, A4 size, 3 p., file: CCPC16.pdf

GTA-16-41 Dias, D. S., Costa, L. H. M. K., Amorim, M. D. - "Capacity Analysis of a City-Wide Hybrid V2V Network", in 7th International Conference Network of the Future IEEE

- NoF’2016, Buzios-RJ, Brazil, November 2016. English, A4 size, 3 p., file: DCA16.pdf

GTA-16-42 Pinto, F. A. V., Costa, L. H. M. K., Menasché, D. S., and Amorim, M. D. - "Space-Aware Modeling of Two-Phase Electric Charging Stations", in IEEE Transactions on Intelligent Transportation Systems, DOI. 10.1109/TITS.2016.2580127, vol. PP, no. 99, July 2016. English, A4 size, 10 p., file: PCM16.pdf

GTA-16-43 Oliveira, E. G., Ribeiro Júnior, J. G., Campista, M. E. M., Costa, L. H. M. K. -“ComuniCAR: Uma Biblioteca de Auxílio ao Desenvolvimento de Aplicações para Redes Mó”, in VI Brazilian Symposium on Computing Systems Engineering (SBESC’16), João Pessoa, PB, Brazil, November 2016 Portuguese, A4 size, 8 p., file: ORC16.pdf

GTA-16-44 Undergraduate Final Project Author - João Emílio Medina Lag Title - "PLATT: Um Novo Algoritmo de Compressão de Dados de Processo com Parametrização Automát Advisor - Miguel Elias Mitre Campista DEL/POLI/UFRJ - September 2016 Portuguese - A4 size - 95 p. - File: Emilio16.pdf

GTA-16-45 Alvarenga, I. D. and Duarte, O. C. M. B - "RIO: A Denial of Service Experimentation Platform in a Future Internet Testbed", in 7th International Conference Network of the Future - NoF’2016, Buzios-RJ, Brazil, November 2016. English, A4 size, 5 p., file: AD16.pdf

GTA-16-46 Torres, J. V., Alvarenga, I. D., Pedroza, A. C. P. and Duarte, O. C. M. B - "Proposing, Specifying, and Validating a Controller-based Routing Protocol for a Clean-Slate Named-Data Networking", in 7th International Conference Network of the Future - NoF’2016, Buzios-RJ, Brazil, November 2016. English, A4 size, 5 p., file: TAPD16.pdf

GTA-16-48 Medeiros, D. S. V., Campista, M. E. M., Mitton, N., Amorim, M. D., Pujolle, G. - “Spread Betweenness Centrality: The Power of Quasi-Shortest Paths”, Technical Report, Electrical Engineering Program, COPPE/UFRJ, September 2016. English, A4 size, 13 p., file: MCM16c.pdf

GTA-16-49 Mauricio, Leopoldo A. F. Mauricio, Rubinstein, Marcelo G. and Duarte, Otto C. M. B. - “Proposing, Specifying, and Validating a Controller-based Routing Protocol for a Clean-Slate Named-Data Networking”, in 7th International Conference Network of the Future - NoF’2016, Buzios-RJ, Brazil, November 2016. English, A4 size, 3 p., file: MRD16.pdf

Technical Reports — 2015

GTA-15-01 Ribeiro, F. J. L., Pedroza, A. C. P. and Costa, L. H. M. K. - "Underwater Monitoring System For Oil Exploration Using Acoustic Sensor Networks", in Telecommunication Systems Journal, Springer, ISSN 1018-4864, DOI 10.1007/s11235-014-9948-6, vol. 58, no. 1, pp. 91-106, January 2015. English, A4 size, 16 p., file: RPC15.pdf

GTA-15-02 Doctor Science Thesis Author - Rodrigo de Souza Couto Title - “Estrat?as e An?se de Resili?ia em Redes de Centros de Dados” Advisors - Lu?Henrique Maciel Kosmalski Costa and Miguel Elias Mitre Campista COPPE/PEE/UFRJ - January 2015 Portuguese, A4 size, 126 p. - File: Couto15.pdf

GTA-15-03 Andreoni Lopez, M., and Duarte, O. C. M. B.

"Providing Elasticity to Intrusion Detection Systems in Virtualized Software Defined Networks" in IEEE International Conference on Communications - ICC 2015, pp. 1-6, June 2015. English, A4 size, 6 p. - File: AnDu15.pdf

GTA-15-04 Undergraduate Final Project Author - Bernardo de Campos Vidal Camilo Title - “An?se da Sobrecarga do IPsec no Mapeamento de Redes Virtuais” Advisors - Lu?Henrique Maciel Kosmalski Costa and Rodrigo de Souza Couto DEL/POLI/UFRJ - February 2015 Portuguese - A4 size - 76 p. - File: Camilo15.pdf GTA-15-05 Couto, R. S., Secci, S., Campista, M. E. M., and Costa, L. H. M. K. -

“Otimiza? do Posicionamento de Servidores F?cos em Centros de Dados Resilientes a Desastres”, in XXXIII Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?s - SBRC’2015, Vit??, ES, Brazil, May 2015. Portuguese, A4 size, 14 p., file: CSC15.pdf

GTA-15-06 Silva, V. B. C., Campista, M. E. M., and Costa, L. H. M. K. - “Uma Estrat?a de Cache Proativo para Distribui? de Conte??em Redes Veiculares”, Accepted for publication in XXXIII Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?s - SBRC’2015, Vit??, ES, Brazil, May 2015. Portuguese, A4 size, 14 p., file: SCC15.pdf

GTA-15-07 Master Science Thesis Author - Dayro Augusto Barahona Hernandez Title - “Uma Avalia? da Influ?ia da Velocidade dos N??o Estabelecimento de Caminhos em Redes Ad Hoc Veiculares” Advisor - Miguel Elias M. Campista and Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - March 2015 Portuguese - A4 size - 66 p. - File: Dayro15.pdf

GTA-15-08 Hernandez, D. A. B., Medeiros, D. S. V., Campista, M. E. M., and Pedroza, A. C. P. - "Uma Avalia? da Influ?ia da Velocidade dos N??o Estabelecimento de Caminhos em Redes Ad Hoc Veiculares", in XXXIII Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?os - SBRC’2015, pp.431-444, Vit??, ES, Brazil, May 2015. Portuguese, A4 size, file: DDCA15.pdf

GTA-15-09 Mattos, D. M. F., Andreoni Lopez, M., Ferraz, L. H. G., and Duarte, O. C. M. B. - “Controlador Resiliente com Distribui? Eficiente para Redes Definidas por Software”, in XXXIII Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?s - SBRC’2015, Vit??, ES, Brazil, May 2015. Portuguese, A4 size, 14 p., file: MAFD15.pdf

GTA-15-10 Andreoni Lopez, M., Mattos, D. M. F., Ferraz, L. H. G., and Duarte, O. C. M. B. - “Localiza? Eficiente de Sensores Colaborativos para Detec? e Preven? de Intrus?em Ambientes Virtualizados”, in XX Workshop de Ger?ia e Opera? de Redes e Servi? (WGRS 2015) - SBRC’2015, Vit??, ES, Brazil, May 2015. Portuguese, A4 size, 14 p., file: AMFD15.pdf

GTA-15-11 Mattos, D. M. F.; Ferraz, L. H. G., and Duarte, O. C. M. B - “Virtual Machine Migration”, Chapter 3 in Cloud Services, Networking and Management book. Edited by Nelson L. S. da Fonseca and Raouf Boutaba, ISBN: 978-1-118-84594-3, pp. 49-72, Wiley-IEEE Press, April 2015. English, Letter size, 24 p., file: MFD15.pdf GTA-15-12 Couto, R. S., Secci, S., Campista, M. E. M., Costa, L. H. M. K. -

“Disaster-Resilient IaaS Cloud Networks: Requirements and Research Directions”, in International Workshop Software Networks - SOFTNET’2015, Rio de Janeiro, RJ, Brazil, March 2015. English, A4 size, 1 p., file: CSC15a.pdf

GTA-15-13 Medeiros, D. S. V., Campista, M. E. M. - "C-RAN: Concepts, Challenges, and Potential Solutions", in International Workshop Software Networks - SOFTNET’2015, Rio de Janeiro, RJ, Brazil, March 2015. English, A4 size, 1 p., file: MeCa15.pdf

GTA-15-14 Silva, V. B. C., Sciammarella,T., Campista, M. E. M., and Costa, L. H. M. K. - “Time of Arrival Prediction with Dynamic Route Tracking for Public Transportation Systems”, in 14th IFIP Annual Mediterranean Ad Hoc Networking Workshop - Med-Hoc-Net 2015, Vilamoura, Algarve, Portugal, June 2015. English, A4 size, 8 p., file: SSC15.pdf

GTA-15-15 Master Science Thesis Author - Vitor Borges Coutinho da Silva Title - "Distribui? de Conte??Orientada ?rajet?? em Redes Veiculares" Advisor - Miguel Elias M. Campista COPPE/PEE/UFRJ - June 2015 Portuguese - A4 size - 102 p. - File: Borges15.pdf

GTA-15-16 Sammarco, M., Belblidia, N., Amorim, M. D., Costa, L. H. M. K., and Conan, V. - “Burst Size Matters: When Node Degree Helps Opportunistic Distribution”, in 14th IFIP Annual Mediterranean Ad Hoc Networking Workshop - Med-Hoc-Net 2015, Vilamoura, Algarve, Portugal, June 2015. English, A4 size, 8 p., file: SBA15.pdf

GTA-15-17 Campista, M. E. M., Sammarco, M., Amorim, M. D., and Razafindralambo, T. - “Monitoramento Colaborativo de Redes Sem-fio: Acurácia do Sistema e Denúncia de Farejadores Maliciosos”, in XIV Workshop em Desempenho de Sistemas Computacionais e de Comunicação - WPerformance 2015, Recife, PE, Brazil, July 2015. Portuguese, A4 size, 14 p., file: CSA15.pdf

GTA-15-18 Couto, R. S. C., Sciammarella, T., Barreto, H. F. S. S. M., Campista, M. E. M., Rubinstein, M. G., Costa, L. H. M. K., Vetter F., Marins, A. L. A. -

“GT-PID: Uma Nuvem IaaS Universitária Geograficamente Distribuída”, in Quinta Conferencia de Directores de Tecnología de Informaci?? TICAL 2015, Vin?del Mar, Chile, July 2015. Portuguese, A4 size, 19 p., file: CSB15.pdf

GTA-15-19 Costa, V. T. and Costa, L. H. M. K. - “Vulnerabilities and solutions for isolation in FlowVisor-based virtual network environments”, in Journal of Internet Services and Applications, ISSN 1869-0238, DOI 10.1186/s13174-015-0034-4, vol. 6, no. 18, pp. 1-9, August 2015. English, A4 size, 9 p., file: CoCo15.pdf GTA-15-20 Belblidia, N., Sammarco, M., Amorim, M. D., and Costa, L. H. M. K. -

“EPICS: Fair Opportunistic Multi-Content Dissemination”, in IEEE Transactions on Mobile Computing, ISSN 1536-1233, DOI 10.1109/TMC.2014.2368558, vol. 14, no. 9, pp. 1847-1860, September 2015. English, A4 size, 14 p., file: BSA15.pdf

GTA-15-21 Piedrahita, A. F. M., Rueda, S., Mattos, D. M. F., and Duarte, O. C. M. B. - “FlowFence: A Denial of Service Defense System for Software Defined Networking”, in Global Information Infrastructure and Networking Symposium 2014 (GIIS’15), Gudalajara, Mexico, October 2015 (to appear). English, Letter size, 6 p., file: PMR15.pdf

GTA-15-22 Cardoso, L. P., Mattos, D. M. F., Duarte, O. C. M. B., and Pujolle, Guy - “An Efficient Energy-Aware Mechanism for Virtual Machine Migration”, in Global Information Infrastructure and Networking Symposium 2014 (GIIS’15), Gudalajara, Mexico, October 2015 (to appear). English, Letter size, 6 p., file: CMD15.pdf

GTA-15-23 Mattos, D.M.F., and Duarte, O. C. M. B. - “Atualização Reversa: Garantindo Consistência de Estados em Redes Definidas por Software”, in Anais do XV Simpósio Brasileiro em Segurança da Informação e de Sistemas Computacionais - SBSeg’15, Florian??is, SC, Brazil, November 2015 (to appear). Portuguese, A4 size, 14 p., file: MaDu15.pdf

GTA-15-24 Sammarco, M., Campista, M. E. M., and Amorim, M. D. - “Scalable Wireless Traffic Capture Through Community Detection and Trace Similarity”, in IEEE Transactions on Mobile Computing, ISSN 1536-1233, DOI 10.1109/TMC.2015.2477809, vol. 15, no. 7, pp. 1757-1769, July 2015. English, A4 size, 13 p., file: SCA15.pdf

GTA-15-25 Master Science Thesis Author - Margarida Maria Bastos Vieira de Souza Title - “Análise de uma Rede Sem Fio Ad Hoc em um Cenário de Missão Militar Tática ” Advisor - Aloysio de Castro Pinto Pedroza and Miguel Elias Mitre Campista COPPE/PEE/UFRJ - March 2015 Portuguese - A4 size - 73 p. - File: Margarida15.pdf

GTA-15-26 Author - Ferraz, L. H. G., Mattos, D. M. F., Laufer, R. P., Duarte, O. C. M. B., and Pujolle, G. Title - “A High-Performance Two-Phase Multipath Scheme for Data-Center Networks” Accept to be published in Compute Network, ISSN 1389-1286, 2016. English, Letter size, 35 p., file: FLD15.pdf

GTA-15-27 Piedrahita, A. F. M., Rueda , S., Mattos, D. F. M., and Duarte, O. C. M. B. - “FlowFence: Um Sistema de Defesa contra Ataques de Negação de Serviço para Redes Definidas por Software”, in Anais do XV Simpósio Brasileiro em Segurança da Informação e de Sistemas Computacionais - SBSeg’15, Florian??is, SC, Brazil, November 2015). Portuguese, A4 size, 4 p., file: PRM15b.pdf

GTA-15-28 Couto, R. S., Secci, S., Campista, M. E. M., Costa, L. H. M. K. - “Reliability and Survivability Analysis of Data Center Network Topologies”, in Journal of Network and Systems Management, Springer, ISSN 1064-7570, DOI. 10.1007/s10922-015-9354-8, vol. 24, no. 2, pp. 346-392, April 2015. English, A4 size, 46 p., file: CSC15b.pdf

GTA-15-29 Couto, R. S., Secci, S., Campista, M. E. M., Costa, L. H. M. K. - “Server Placement with Shared Backups for Disaster-Resilient Clouds”, to appear in Computer Networks, Elsevier, ISSN 1389-1286, DOI. 10.1016/j.comnet.2015.09.039, 2015. English, A4 size, 21 p., file: CSC15c.pdf

GTA-15-30 Undergraduate Final Project Author - Tatiana Sciammarella Title - “Análise do Tráfego de Rede em Cenários de Computação em Nuvem Geodistribuídos” Advisors - Miguel Elias Mitre Campista and Luís Henrique Maciel Kosmalski Costa DEL/POLI/UFRJ - Agosto 2015 Portuguese - A4 size - 59 p. - File: Tatiana15.pdf

GTA-15-31 Doctor Science Thesis Author - Lyno Henrique Gon?ves Ferraz Title - “A High-Performance Two-Phase Multipath Scheme for Data-Center Networks” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - January 2015 English, A4 size, 79 p. - File: Lyno15.pdf

GTA-15-32 Undergraduate Final Project Author - Igor Drummond Alvarenga Title - “RIO: Uma Plataforma para Experimentação de Ataques de Negação de Serviço em uma Rede de Testes para Internet do Futuro” Advisor - Otto Carlos Muniz Bandeira Duarte DEL/POLI/UFRJ - November 2015 Portuguese - A4 size - 68 p. - File: Igor15.pdf

Technical Reports — 2014

GTA-14-01 Moraes, I. M., Mattos, D. M. F., Ferraz, L. H. G., Campista, M. E. M., Rubinstein, M. G., Costa, L. H. M. K., Amorim, M. D., Velloso, P. B., Duarte, O. C. M. B., and Pujolle, G. - “FITS: A flexible virtual network testbed architecture”, in Elsevier Computer Networks, DOI. 10.1016/j.bjp.2014.01.002, vol. 63, pp. 221-237, April 2014. English, A4 size, 17 p., file: MMF14.pdf

GTA-14-02 Ribeiro J??r, J. G., Campista, M. E. M., and Costa, L. H. M. K. - “COTraMS: A Collaborative and Opportunistic Traffic Monitoring System”, in IEEE Transactions on Intelligent Transportation Systems, DOI. 10.1109/TITS.2013.2291996, vol. 15, no. 3, pp. 949-958, June 2014. English, A4 size, 10 p., file: RCC14.pdf GTA-14-03 Couto, R. S., Campista, M. E. M., and Costa, L. H. M. K. - “Network Resource

Control for Xen-based Virtualized Software Routers”, in Computer Networks, Elsevier, ISSN 1389-1286, DOI. 10.1016/j.comnet.2014.02.003, vol. 64, pp. 71-88, May 2014. English, A4 size, 53 p., file: CCC14.pdf

GTA-14-04 Undergraduate Final Project Author - Victor Torres da Costa Title - "Controle e Isolamento de Recursos em Ambientes de Redes Virtuais OpenFlow" Advisors - Luis Henrique Maciel Kosmalski Costa DEL/POLI/UFRJ - February 2014 Portuguese - A4 size - 71 p.

GTA-14-05 Silva, V. B. C., Silva, Sciammarella,T., Campista, M. E. M., and Costa, L. H. M. K. - “WiBus: Um Sistema de Monitoramento de Transportes P??cos Usando Redes IEEE 802.11”, to appear in XXXII Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?s - SBRC’2014, Florian??is, SC, Brazil, May 2014. Portuguese, A4 size, 14 p., file: SSC14.pdf

GTA-14-06 Barreto, H. F. S. S. M., Campista, M. E. M., and Costa, L. H. M. K. - “Spammers no Twitter: Quando Contatos Deixam de ser Bem-vindos”, to appear in Workshop de Redes P2P, Din?cas, Sociais e Orientadas a Conte??(Wp2p+ 2014) - SBRC’2014, Florian??is, SC, Brazil, May 2014. Portuguese, A4 size, 14 p., file: BCC14.pdf

GTA-14-07 Lobato, A. G. P., Figueiredo, U. da R., Andreoni L., M., and Duarte, O. C. M. B. - "Uma Arquitetura El?ica para Preven? de Intrus?em Redes Virtuais usando Redes Definidas por Software", in XXXII Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?s - SBRC’2014, Florian??is, SC, Brazil, May 2014. Portuguese, A4 size, 14 p., file: LFA14.pdf

GTA-14-08 Mattos, D.M.F., Ferraz, L.H.G., and Duarte, O. C. M. B. - "AuthFlow: Authentication and Access Control Mechanism for Software Defined Networking", Technical Report, Electrical Engineering Program, COPPE/UFRJ, April 2014. English, A4 size, 7 p., file: MFD14.pdf

GTA-14-09 Ferraz, L.H.G., Mattos, D.M.F., and Duarte, O. C. M. B. - "A Two-Phase Multipathing Scheme with Genetic Algorithm for Data Center Network", Tto appear in IEEE Global Communications Conference - GLOBECOM’2012, Austin, Texas, USA, December 2014. English, A4 size, 7 p., file: FMD14.pdf

GTA-14-10 Mattos, D.M.F., and Duarte, O. C. M. B. - "XenFlow: Seamless Migration Primitive and Quality of Service for Virtual Networks", to appear in IEEE Global Communications Conference - GLOBECOM’2012, Austin, Texas, USA, December 2014. English, A4 size, 7 p., file: MaDu14.pdf

GTA-14-11 Ramos, F. B., Alvarenga, I. D., Caldas, P. M., Duarte, O. C. M. B. - "Distribui? de V?o sob Demanda Baseada em Redes de Distribui? de Conte??utilizando Redes Orientadas a Conte??, in XXXII Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?s - SBRC’2014, Florian??is, SC, Brazil, May 2014. Portuguese, A4 size, 14 p., file: RAC14.pdf

GTA-14-12 Mattos, D. M. F. e Duarte, O. C. M. B. - "AuthFlow: Um Mecanismo de Autentica? e Controle de Acesso para Redes Definidas por Software", in XXXII Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?s - SBRC’2014, Florian??is, SC, Brazil, May 2014. Portuguese, A4 size, 14 p., file: MaDu14a.pdf GTA-14-13 Ferraz, L. H. G., Duarte, Mattos, D. M. F., O. C. M. B. - "Um Esquema de Mul-

ticaminhos com Algoritmos Gen?cos para Redes de Centro de Dados", in XXXII Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?s - SBRC’2014, Florian??is, SC, Brazil, May 2014. Portuguese, A4 size, 14 p., file: FMD14a.pdf

GTA-14-14 Cardoso, L. P., Ferraz, L. H. G., Duarte, O. C. M. B. - "Migra? de m?inas virtuais para economia de energia", in XXXII Workshop de Ger?ia e Opera? de Redes e Servi? (WGRS 2014) - SBRC’2014, Florian??is, SC, Brazil, May 2014. Portuguese, A4 size, 14 p., file: CFD14.pdf

GTA-14-15 Bezerra, G. M. G., Mattos, D. M. F., Ferraz, L. H. G., Duarte, O. C. M. B. - "Sistema Automatizado de Ger?ia de Recursos para Ambientes Virtualizados", in XXXII Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?s - SBRC’2014, Florian??is, SC, Brazil, May 2014. Portuguese, A4 size, 14 p., file: BMF14.pdf

GTA-14-16 Santos, A. J., Braga, M. L., Velloso, P. B., Ribeiro J??r, J. G., Costa, L. H. M. K. - "An?se da Capacidade de uma Rede Tolerante a Atrasos e Desconex??na Bacia Hidrogr?ca Amazonense", to appear in XXXII Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?s - SBRC’2014, Florian??is, SC, Brazil, May 2014. Portuguese, A4 size, 14 p., file: SBV14.pdf

GTA-14-17 Couto, R. S., Secci, S., Campista, M. E. M., and Costa, L. H. M. K. - "Lat?ia Versus Sobreviv?ia no Projeto de Centros de Dados Geograficamente Distribu?s", in XXXII Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?s - SBRC’2014, pp. 809-822, Florian??is, SC, Brazil, May 2014. Portuguese, A4 size, 14 p., file: CSC14.pdf

GTA-14-18 Master Science Thesis Author - Mart?Esteban Andreoni Lopez Title - “Uma Arquitetura de Detec? e Preven? de Intrus?para Redes Definidas por Software" Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - May 2014 Portuguese - A4 size - 93 p. - File: Martin14.pdf

GTA-14-19 Master Science Thesis Author - Andr?Felipe Murillo Piedrahita Title - "Ferramenta de Avalia? de Ataques de Nega? de Servi?em uma Plataforma de Testes" Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - May 2014 Portuguese - A4 size - 96 p. - File: Andres14.pdf

GTA-14-20 Ferraz, L. H. G., Velloso, P. B., and Duarte, O. C. M. B. - “An accurate and precise malicious node exclusion mechanism for ad hoc networks”, Ad Hoc Networks, Volume 19, August 2014, Pages 142-155. English, A4 size, 14 p., file: FVD14.pdf

GTA-14-21 Torres, J.V., and Duarte, O. C. M. B. - "CRoS-NDN: Controller-based Routing Strategy for Named Data Networking", Technical Report, Electrical Engineering Program, COPPE/UFRJ, April 2014. English, A4 size, 7 p., file: ToDu14a.pdf GTA-14-22 Torres, J. V., and Duarte, O. C. M. B. - "Uma Estrat?a R?da e Eficiente de Lo-

caliza? e Encaminhamento em Redes Orientadas a Conte??, in XXXII Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?s - SBRC’2014, Florian??is, SC, Brazil, May 2014. Portuguese, A4 size, 14 p., file: ToDu14b.pdf

GTA-14-23 Undergraduate Final Project Author - Leonardo Pais Cardoso Title - "Migra? de M?inas Virtuais para Economia de Energia" Advisors - Lu?Henrique M. K. Costa, Otto Carlos M. B. Duarte and Lyno Henrique G. Ferraz DEL/POLI/UFRJ - June 2014 Portuguese - A4 size - 73 p. - File: Cardoso14.pdf

GTA-14-24 Silva, V. B. C., Gomes, L. C., Lanza, M. L. D., Campista, M. E. M., and Costa, L. H. M. K. - "Um Prot??o de Rede Sem-Fio para o Trem de Levita? Magn?ca MAGLEV COBRA", in XII Rio de Transportes - RDT 2014, Rio de Janeiro, RJ, Brazil, August 2014. Portuguese, A4 size, 1p., file: SGL14.pdf

GTA-14-25 Couto, R. S, Secci, S., Campista, M. E. M., Costa, L. H. M. K - "Latency Versus Survivability in Geo-Distributed Data Center Design", accepted for publication in IEEE Global Communications Conference - GLOBECOM’2014, Austin, Texas, USA, December 2014. English, A4 size, 7 p., file: CSC14a.pdf

GTA-14-26 Couto, R. S, Secci, S., Campista, M. E. M., Costa, L. H. M. K - "Network Design Requirements for Disaster Resilience in IaaS Clouds", in IEEE Communications Magazine, Vol. 53, No. 10, pp 52-58, October 2014. English, A4 size, 11 p., file: CSC14b.pdf

GTA-14-27 Santos, A. J., Braga, M. L., Velloso, P. B., Ribeiro J??r, J. G., Costa, L. H. M. K. - "Capacity Analysis of a Delay and Disruption Tolerant Network in the Amazon Basin", in Global Information Infrastructure and Networking Symposium 2014 (GIIS’14), Montreal, Canada, September 2014. English, A4 size, 6 p., file: SBV14a.pdf

GTA-14-28 Silva, V. B. C., Sciammarella,T., Campista, M. E. M., and Costa, L. H. M. K. - "WiBus: A Wi-Fi based Monitoring System for Public Transportation with Dynamic Route Tracking", in 2014 IFIP Wireless Days - WD’2014, Rio de Janeiro, RJ, Brazil, November 2014. English, A4 size, 3 p., file: SSC14b.pdf

GTA-14-29 Braga, M. L., Santos, A. J., Pedroza, A. C. P., Costa, L. H. M. K. - "Planejamento de Rotas com Algoritmos Anytime em Redes Veiculares na Plataforma Raspberry Pi", in IV Simpósio Brasileiro de Engenharia de Sistemas Computacionais 2014 (SBESC’14), Manaus, AM, Brazil, November 2014. Portuguese, A4 size, 6 p., file: BSP14.pdf

Technical Reports — 2013

GTA-13-01 Silva, V. B. C., Silva, F. O. B., Campista, M. E. M., and Costa, L. H. M. K. - “A Trajectory-Based Approach to Improve Delivery in Drive-Thru Internet Scenarios”, in Workshop on Emerging Vehicular Networks: V2V/V2I and Railroad Communications - IEEE International Conference on Communications - ICC 2013, pp. 489-494, June 2013. English, A4 size, 6 p., file: SSC13.pdf

GTA-13-02 Silva, V. B. C., Silva, F. O. B., Campista, M. E. M., and Costa, L. H. M. K. - “Roteamento Baseado na Trajet?? para Redes Veiculares Desconectadas com M??plos Gateways”, in XXXI Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?s - SBRC’2013, pp. 1038-1051, Bras?a, DF, Brazil, May 2013. Portuguese, A4 size, 14 p., file: SSC13a.pdf GTA-13-03 Ribeiro J??r, J. G., Quintanilha. I. M., Campista, M. E. M., Costa, L. H. M. K.

- “Sistema para Monitoramento Descentralizado de Tr?ito baseado em Redes Veiculares Infraestruturadas”, in XXXI Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?s - SBRC’2013, Bras?a, DF, pp. 863-876, Brazil, May 2013. Portuguese, A4 size, 14 p., file: JCC13.pdf

GTA-13-04 Phe-Neau, T., Campista, M. E. M., de Amorim, M. D., Conan, V. - “Padr??de Mobilidade de Vizinhan?em Redes de Contato Intermitente”, in XXXI Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?s - SBRC’2013, pp. 397-410, Bras?a, DF, Brazil, May 2013. Portuguese, A4 size, 14 p., file: PCA13.pdf

GTA-13-05 Master Science Thesis Author - Daniel de Souza Dias Title - “Posicionamento Online Baseado no Tr?go de M?inas Virtuais em Rede de Data Center" Advisor - Lu?Henrique Maciel Kosmalski Costa COPPE/PEE/UFRJ - March 2013 Portuguese - A4 size - 74 p. - File: Daniel13.pdf

GTA-13-06 Guimar?, P. H. V., Murillo, A., Andreoni, M., Mattos, D. M. F., Ferraz, L. H. G., Pinto, F. A. V., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Comunica? em Redes El?icas Inteligentes: Efici?ia, Confiabilidade, Seguran?e Escalabilidade”, in Minicursos do Simp?? Brasileiro de Redes de Computadores - SBRC’2013, pp. 101-164, Bras?a, DF, Brazil, May 2013. Portuguese, A4 size, 64 p., file: GMAM13.pdf

GTA-13-07 Torres, J. V., Ferraz, L. H. G., and Duarte, O. C. M. B. - “Redes Orientadas a Conte??Baseadas em Controladores Hier?uicos”, in XXXI Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?s, 14 p., Bras?a, pp. 717-730, DF, Brazil, May 2013. Portuguese, A4 size, 14 p., file: TFD13.pdf

GTA-13-08 Figueiredo, U. da R., Lobato, A. G. P., Mattos, D. M. F., Ferraz L. H. G., and Duarte, O. C. M. B. - “An?se de Desempenho de Mecanismos de Encaminhamento de Pacotes em Redes Virtuais”, in XXXI Workshop de Ger?ia e Opera? de Redes e Servi? (WGRS 2013) - SBRC’2013, pp. 61-84, Bras?a, DF, Brazil, May 2013. Portuguese, A4 size, 14 p., file: FLMF13.pdf

GTA-13-09 Guimar?, P. H. V., Ferraz, L. H. G., Torres, J. V., Mattos, D. M. F., Murillo P., A. F., Andreoni L., M. E., Alvarenga, I. D., Rodrigues, C. S. C., and Duarte, O. C. M. B. - “Experimenting Content-Centric Networks in the Future Internet Testbed Environment”, in Workshop of Cloud Convergence: Challenges for Future Infrastructures and Services (WCC-02) - ICC’2013, Budapest, Hungary, June 2013. English, A4 size, 5 p., file: GFTM13a.pdf

GTA-13-10 Sammarco, M., Campista, M. E. M., and Amorim, M. D. - “Trace Selection for Improved WLAN Monitoring”, in 5th ACM HotPlanet Workshop, pp. 9-14, Hong Kong, China, August 2013. English, A4 size, 6 p., file: SCA13.pdf

GTA-13-11 Doctor Science Thesis Author - Jos?eraldo Ribeiro J??r Title - “Sistema Oportun?ico e Colaborativo para Monitoramento de Tr?ito baseado em Redes Veiculares Infraestruturadas” Advisors - Lu?Henrique Maciel Kosmalski Costa COPPE/PEE/UFRJ - June 2013 Portuguese, A4 size, 114 p. - File: JGRJunior13.pdf

GTA-13-12 Undergraduate Final Project Author - Vitor Borges Coutinho da Silva Title - “Um Sistema de Localiza? e Previs?de Chegada dos Ve?los de Transporte P??co Usando Redes IEEE 802.11 Advisors - Miguel Elias Mitre Campista and Lu?Henrique Maciel Kosmalski Costa DEL/POLI/UFRJ - August 2013 Portuguese - A4 size - 95 p. - File: Borges13.pdf

GTA-13-19 Campista, M. E. M., Rubinstein, M. G., Moraes, I. M., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Challenges and Research Directions for the Future Internetworking”, in IEEE Communications Surveys & Tutorials, v. 16, no. 2, p. 1050-1079, Second Quarter 2014. English, A4 size, 31 p., file: CRM13.pdf

GTA-13-20 Ribeiro J??r, J. G., Campista, M. E. M., and Costa L. H. M. K. - “A Decentralized Traffic Monitoring System Based on Vehicle-to-Infrastructure Communications”, in Wireless Days Conference, Valencia, Spain, p. 1-6, November 2013. English, A4 size, 6 p., file: RCC13.pdf

GTA-13-21 Phe-Neau, T., de Amorim, M. D., Campista, M. E. M., Conan, V. - “Examining Vicinity Dynamics in Opportunistic Networks”, in 8th ACM Workshop on Performance Monitoring and Measurement of Heterogeneous Wireless and Wired Networks (ACM PM2HW2N’13), Presented at the ACM MSWiM’13 Poster Session, Barcelona, Spain, pp. 153-160, November 2013. English, A4 size, 8 p., file: PAC13.pdf

GTA-13-22 Murillo, A., Mattos, D. M. F., Ferraz, L. H. G., Duarte, O. C. M. B. - "FlowFence: Um Sistema para Preven? de Ataques de Nega? de Servi?usando Controle de Banda", Technical Report, Electrical Engineering Program, COPPE/UFRJ, December 2013. Portuguese, A4 size, 14 p., file: MMFD13.pdf

GTA-13-23 Ramos, F. B., Alvarenga, I. D., Caldas, P. M., Duarte, O. C. M. B. - "Distribuicao de V?o sob Demanda Baseada em Redes de Distribui? de Conte??utilizando Redes Orientadas a Conte??, Technical Report, Electrical Engineering Program, COPPE/UFRJ, December 2013. Portuguese, A4 size, 14 p., file: RACD13.pdf

GTA-13-24 Mattos, D. M. F. e Duarte, O. C. M. B. - "AuthFlow: Um Mecanismo de Autentica? e Controle de Acesso para Redes Definidas por Software", Technical Report, Electrical Engineering Program, COPPE/UFRJ, December 2013. Portuguese, A4 size, 14 p., file: MaDu13.pdf

GTA-13-25 Andreoni L., M., Figueiredo, U. da R., Lobato, A. G. P. e Duarte, O. C. M. B. - "BroFlow: Um Sistema Eficiente de Detec? e Preven? de Intrus?em Redes Definidas por Software", Technical Report, Electrical Engineering Program, COPPE/UFRJ, December 2013. Portuguese, A4 size, 14 p., file: AFLD13.pdf GTA-13-26 Ferraz, L. H. G., Mattos, D. M. F. e Duarte, O. C. M. B. - "Um Esquema

de Multicaminhos com Algoritmos Gen?cos para Redes de Centro de Dados", Technical Report, Electrical Engineering Program, COPPE/UFRJ, December 2013. Portuguese, A4 size, 14 p., file: MFD13.pdf

GTA-13-27 Lobato, A. G. P., Figueiredo, U. da R., Andreoni L., M. e Duarte, O. C. M. B. - "Uma Arquitetura El?ica para Preven? de Intrus?em Redes Virtuais usando Redes Definidas por Software", Technical Report, Electrical Engineering Program, COPPE/UFRJ, December 2013. Portuguese, A4 size, 14 p., file: LFLD13a.pdf

GTA-13-28 Torres, J. V. e Duarte, O. C. M. B. - "Uma Estrat?a R?da e Eficiente de Localiza? e Encaminhamento em Redes Orientadas a Conte??, Technical Report, Electrical Engineering Program, COPPE/UFRJ, December 2013. Portuguese, A4 size, 14 p., file: ToDu13.pdf

GTA-13-29 Ferraz, L. H. G., and Duarte, O. C. M. B. - "Analysis of Multipathing Techniques in Data Center Networks", in Second Workshop on Network Virtualization and Intelligence for the Future Internet - WNetVirt’2013, Angra dos Reis, RJ, Brazil, October 2013. English, A4 size, 1p., file: FeDu13.pdf

GTA-13-30 Murillo, A., and Duarte, O. C. M. B. - "Virtual Networks for Cyber Security Testing of Future Internet Technologies", in Second Workshop on Network Virtualization and Intelligence for the Future Internet - WNetVirt’2013, Angra dos Reis, RJ, Brazil, October 2013. English, A4 size, 1p., file: MuDu13.pdf

GTA-13-31 Andreoni, M., and Duarte, O. C. M. B. - "Detecting and Preventing Intruders for Cyber-Security", in Second Workshop on Network Virtualization and Intelligence for the Future Internet - WNetVirt’2013, Angra dos Reis, RJ, Brazil, October 2013. English, A4 size, 1p., file: AnDu13.pdf

GTA-13-32 Mattos, D. M. F., and Duarte, O. C. M. B. - "Authentication and Access Control Architecture for Software Defined Networks", in Second Workshop on Network Virtualization and Intelligence for the Future Internet - WNetVirt’2013, Angra dos Reis, RJ, Brazil, October 2013. English, A4 size, 1p., file: MaDu13a.pdf

GTA-13-33 da Costa, V. T, and Costa, L. H. M. K. - "Vulnerability Study of FlowVisor-based Virtualized Network Environments", in Second Workshop on Network Virtualization and Intelligence for the Future Internet - WNetVirt’2013, Angra dos Reis, RJ, Brazil, October 2013. English, A4 size, 1p., file: CoCo13.pdf

GTA-13-34 Lobato, A. G. P., Ulisses Figueiredo, and Duarte, O. C. M. B. - "An Architecture for Intrusion Prevention using Software Defined Networks", in Second Workshop on Network Virtualization and Intelligence for the Future Internet - WNetVirt’2013, Angra dos Reis, RJ, Brazil, October 2013. English, A4 size, 1p., file: LUD13.pdf GTA-13-35 Sammarco, M., Campista, M. E. M., and Amorim, M. D. - "Towards a Scalable

WiFi Monitoring System", in Second Workshop on Network Virtualization and Intelligence for the Future Internet - WNetVirt’2013, Angra dos Reis, RJ, Brazil, October 2013. English, A4 size, 1p., file: SCA13a.pdf

GTA-13-36 Cinelli, L. P. and Duarte, O. C. M. B. - "Unified Authentication with Smart Cards via Near Field Communication", in Second Workshop on Network Virtualization and Intelligence for the Future Internet - WNetVirt’2013, Angra dos Reis, RJ, Brazil, October 2013. English, A4 size, 1p., file: CiDu13.pdf

GTA-13-37 Caldas, P. M. and Duarte, O. C. M. B. - "Performance Evaluation of Content Centric Networks", in Second Workshop on Network Virtualization and Intelligence for the Future Internet - WNetVirt’2013, Angra dos Reis, RJ, Brazil, October 2013. English, A4 size, 1p., file: CaDu13.pdf

GTA-13-38 Ramos, F. B. and Duarte, O. C. M. B. - "A Network Architecture to Distribute Video on Demand based on Content Centric Networking", in Second Workshop on Network Virtualization and Intelligence for the Future Internet - WNetVirt’2013, Angra dos Reis, RJ, Brazil, October 2013. English, A4 size, 1p., file: RaDu13.pdf

GTA-13-39 Bezerra, G. M. G. and Duarte, O. C. M. B. - "Autonomic Resource Management System for Virtualized Environments", in Second Workshop on Network Virtualization and Intelligence for the Future Internet - WNetVirt’2013, Angra dos Reis, RJ, Brazil, October 2013. English, A4 size, 1p., file: BeDu13.pdf

GTA-13-40 Figueiredo, U. R., Lobato, A. G. P., and Duarte, O. C. M. B. - "Performance Evaluation of Packet Forwarding Mechanism in Virtual Networks", in Second Workshop on Network Virtualization and Intelligence for the Future Internet - WNetVirt’2013, Angra dos Reis, RJ, Brazil, October 2013. English, A4 size, 1p., file: FLD13.pdf

GTA-13-41 Garcia, R. G. and Duarte, O. C. M. B. - "Performance and Overhead Analysis of Security Mechanisms in Virtual Private Networks", in Second Workshop on Network Virtualization and Intelligence for the Future Internet - WNetVirt’2013, Angra dos Reis, RJ, Brazil, October 2013. English, A4 size, 1p., file: GaDu13.pdf

GTA-13-42 Camilo, B.C.V, and Costa, L. H. M. K. - "Implementation of a Customizable Security-Aware Virtual Network", in Second Workshop on Network Virtualization and Intelligence for the Future Internet - WNetVirt’2013, Angra dos Reis, RJ, Brazil, October 2013. English, A4 size, 1p., file: CaCo13.pdf

GTA-13-43 Campista, M. E. M., Costa, L. H. M. K. Duarte, O. C. M. B., Fonseca, N. L. S., Madeira, E. R. M., Moraes, I. M., and Pujolle, G. - “Virtual Networks - Pluralist Approach for the Next Generation of the Internet”, Editors: Duarte, O. C. M. B. and Pujolle. G., ISTE Ltd and John Wiley and Sons, Inc., 1st ed., 2013. English, 291 p.

GTA-13-44 Phe-Neau, T., Campista, M. E. M., Amorim, M. D., Conan, V. - “An?se e Gera? de Vizinhan?em Redes de Contato Intermitente”, in Revista Brasileira de Redes de Computadores e Sistemas Distribu?s - RB-RESD, v. 6, no. 1, p. 31-46, December 2013 Portuguese, A4 size, 14 p., file: PCA13a.pdf

Technical Reports — 2012

GTA-12-01 Carvalho, H. E. T., Fernandes, N. C., Duarte, O. C. M. B., Pujolle, G. - "SLAPv: A Service Level Agreement Enforcer for Virtual Networks”, in International Conference on Computing, Networking and Communications, Internet Services and Applications Symposium (ICNC12 - ISA), January 2012. English, A4 size, 5 p., file: CFDP12.pdf

GTA-12-02 Fernandes, N. C., Duarte, O. C. M. B. - "VIPER: Fine Control of Resource Sharing in Virtual Networks”, Technical Report, Electrical Engineering Program, COPPE/UFRJ, January 2012. English, A4 size, 12 p., file: FeDu12.pdf

GTA-12-03 Ribeiro, F. J. L., Pedroza, A. C. P. and Costa, L. H. M. K. - “Deepwater Monitoring System in Underwater Delay/Disruption Tolerant Network”, in IEEE Latin America Transactions, IEEE, ISSN 1548-0992, DOI 10.1109/TLA.2012.6142480, vol. 10, no. 1, pp. 1324-1331, January 2012. Portuguese, A4 size, 8 p., file: RPC12.pdf

GTA-12-04 Campista, M. E. M., Costa, L. H. M. K., and Duarte, O. C. M. B. - “A routing protocol suitable for backhaul access in wireless mesh networks”, in Computer Networks, Elsevier, ISSN 1389-1286, DOI 10.1016/j.comnet.2011.10.016, Vol. 56, No. 2, pp. 703-718, February 2012. English, A4 size, 16 p., file: CCD12.pdf

GTA-12-05 Ribeiro J??r, J. G., Campista, M. E. M., Costa, L. H. M. K - “Opportunistic System for Collaborative Traffic Monitoring Using Existing IEEE 802.11 Networks”, in in Intelligent Vehicular Networking: V2V/V2I Communications and Applications - IEEE International Conference on Communications - ICC 2012, June 2012. English, A4 size, 16 p., file: RJCC12.pdf

GTA-12-06 Couto, R. S., Campista, M. E. M., Costa, L. H. M. K. - “Uma Avalia? da Robustez Intra Data Centers Baseada na Topologia da Rede”, in XXX Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?os - SBRC’2012, pp. 610-623, Ouro Preto, MS, Brazil, May 2012. Portuguese, A4 size, 14 p., file: CCC12.pdf

GTA-12-07 Ribeiro, F. J. L., Pedroza, A. C. P. and Costa, L. H. M. K. - “Deepwater Applications for Brazilian Pre-Salt Exploration Using Underwater Sensor Networks”, accepted in The 22nd International Offshore (Ocean) and Polar Engineering Conference - ISOPE-2012, Rhodes, Greece, June 2012. English, A4 size, 7 p., file: RPC12a.pdf GTA-12-08 Mattos, D. M. F., Ferraz, L. H. G., Costa, L. H. M. K., and Duarte, O. M. C. B. -

“Evaluating Virtual Router Performance for a Pluralist Future Internet”, accepted in The 3rd International Conference on Information and Communications Systems - ICICS-2012, Irbid, Jordan, April 2012. English, A4 size, 6 p., file: MFCD12.pdf

GTA-12-09 Carvalho, H. E. T., and Duarte, O. M. C. B. - “VOLTAIC : Volume Optimization Layer To AssIgn Cloud resources”, accepted in The 3rd International Conference on Information and Communications Systems - ICICS-2012, Irbid, Jordan, April 2012. English, A4 size, 6 p., file: CD12.pdf

GTA-12-10 Fragni, C. and Costa, L. H. M. K. - “ECO-ALOC: Energy-Efficient Resource Allocation for Cluster-Based Software Routers”, in Computer Networks, Elsevier, ISSN 1389-1286, DOI 10.1016/j.comnet.2012.03.007, March 2012. English, A4 size, 12 p., file: FC12.pdf

GTA-12-11 M. Sammarco, N. Belblidia, Y. Loppez, M. D. de Amorim, L. H. M. K. Costa, and J. Leguay - “PePiT: Opportunistic Dissemination of Large Contents on Android Mobile Devices”, in ACM MobiOpp 2012 Demos, Zurich, Switzerland, March 2012. English, A4 size, 2 p., file: SBL12.pdf

GTA-12-12 Costa, L. H. M. K., Amorim, M. D., Campista, M. E. M., Rubinstein, M. G., Florissi, P., and Duarte, O. C. M. B. - “Grandes Massas de Dados na Nuvem: Desafios e T?icas para Inova?”, in Minicursos do Simp?? Brasileiro de Redes de Computadores - SBRC’2012, Ouro Preto, MG, Brazil, May 2012. Portuguese, A4 size, 58 p., file: CAC12.pdf

GTA-12-13 Mattos, D. M. F., Mauricio, L. H., Cardoso, L. P., Alvarenga, I. D, Ferraz, L. H. G., and Duarte, O. C. M. B. - “Uma Rede de Testes Interuniversit?a a com T?icas de Virtualizacao H?ridas”, in Sal?de Ferramentas do XXX Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?os - SBRC’2012, Ouro Preto, MG, Brazil, May 2012. Portuguese, A4 size, 8 p., file: MMC12.pdf

GTA-12-14 Mattos, D. M. F., and Duarte, O. C. M. B. - “QFlow: Um Sistema com Garantia de Isolamento e Oferta de Qualidade de Servi?para Redes Virtualizadas”, in XXX Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?os - SBRC’2012, Ouro Preto, MG, Brazil, May 2012. Portuguese, A4 size, 14 p., file: MaDu12.pdf

GTA-12-15 Campista, M. E. M., Dias de Amorim, M., and Costa, L. H. M. K. - “Big Wireless Measurement Campaigns: Are They Really Worth the Price?”, in 4th ACM International Workshop on Hot Topics in Planet-Scale Measurements - ACM HotPlanet 2012, pp. 27-32, Low Wood Bay, Lake District, UK, June 2012. English, Letter size, 6 p., file: CDC12.pdf

GTA-12-16 Mattos, D. M. F., Ferraz, L. H. G., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Virtual Network Performance Evaluation for Future Internet Architectures”, in Journal of Emerging Technologies in Web Intelligence - JETWI (to appear), Academy Publisher, ISSN 1798-0461. English, A4 size, 11 p., file: MFCD12.pdf GTA-12-17 Carvalho, H. E. T., and Duarte, O. C. M. B. - “Elastic Allocation and Automatic

Migration Scheme for Virtual Machines”, in Journal of Emerging Technologies in Web Intelligence - JETWI (to appear), Academy Publisher, ISSN 1798-0461. English, A4 size, 11 p., file: CD12a.pdf

GTA-12-18 Couto, R. S, Campista, M. E. M., Costa, L. H. M. K - “A Reliability Analysis of Datacenter Topologies”, in IEEE Global Communications Conference - GLOBE-COM’2012, Anaheim, California, USA, December 2012 English, A4 size, 6 p., file: CCC12a.pdf

GTA-12-19 Belblidia, N., Amorim, M. D., Costa, L. H. M. K., Leguay, J., and Conan, V. - “Part-whole dissemination of large multimedia contents in opportunistic networks”, Computer Communications, Elsevier, ISSN 0140-3664, DOI http://dx.doi.org/10.1016/j.comcom.2012.03.006, Available Online March 2012. English, A4 size, 12 p., file: BAC12.pdf

GTA-12-20 Ribeiro J??r, J. G., Quintanilha, I. M., Campista, M. E. M., Costa, L. H. M. K. - “Evaluation of an Opportunistic Collaborative Traffic Monitoring System”, in IFIP/IEEE Wireless Days Conference 2012, November 2012. English, A4 size, 6 p., file: RJQCC12.pdf

GTA-12-21 Alves, R. S., Campista, M. E. M., Costa, L. H. M. K., and Duarte, O.C. M. B. - “Towards a Pluralist Internet Using a Virtual Machine Server for Network Customization”, in Asian Internet Engineering Conference - AINTEC’2012, pp. 9-16, Bangkok, Thailand, November 2012. English, Letter size,8 p., file: ACC12.pdf

GTA-12-22 Dias, D. S., Costa, L. H. M. K., “Online Traffic-aware Virtual Machine Placement in Data Center Networks"., in Global Information Infrastructure and Networking Symposium 2012 (GIIS’12), Choroni, Venezuela, December 2012. English, A4 size, 8 p., file: DC12.pdf

GTA-12-23 Doctor Science Thesis Author - Fabr?o Jorge Lopes Ribeiro Title - “Sistema de Monitoramento Subaqu?co para Explora? de Petr?? Usando Redes de Sensores Ac??cos" Advisors - Aloysio de Castro Pinto Pedroza and Lu?Henrique Maciel Kosmalski Costa COPPE/PEE/UFRJ - November 2012 Portuguese, A4 size, 114 p. - File: FabricioJorge12.pdf

GTA-12-24 Undergraduate Final Project Author - Pedro Henrique Valverde Guimaraes Title - “Arquitetura de Gerenciamento de Identidades Usando OpenID e Cartoes Inteligentes" Advisors - Otto Carlos M. B. Duarte DEL/POLI/UFRJ - November 2012 Portuguese - A4 size - 84 p. - File: Guimaraes12.pdf

GTA-12-25 Master Science Thesis Author - Diogo Menezes Ferrazani Mattos Title - “Uma Arquitetura de Virtualiza? de Redes Orientada ?Migra? com Qualidade de Servi? Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - December 2012 Portuguese - A4 size - 74 p. - File: Diogo12.pdf

GTA-12-26 Master Science Thesis Author - Hugo Eiji Tibana Carvalho Title - “VOLTAIC: Um Sistema de Ger?ia Automatizada de Recursos de N??irtuais Para Computa? em Nuvens" Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - December 2012 Portuguese - A4 size - 94 p. - File: Hugo12.pdf

GTA-12-27 Pisa, P. S., and Abdalla, M., and Duarte, O. C. M. B. - “Somewhat Homomorphic Encryption Scheme for Arithmetic Operations on Large Integers”, in 4th Global Information Infrastructure and Networking Symposium - GIIS-2012, Choroni, Venezuela, IEEE, December 2012. English, A4 size, 8 p., file: PAD12.pdf

GTA-12-28 Torres, J. V., Ferraz, L. H. G., and Duarte, O. C. M. B. - "Controller-based Routing Scheme for Named Data Network", Technical Report, Electrical Engineering Program, COPPE/UFRJ, December 2012. English, A4 size, 6 p., file: TFD12.pdf

GTA-12-29 Mattos, D. M. F., Mauricio, L. H., Ferraz, L. H. G., and Duarte, O. C. M. B. - "Um Mecanismo de Comuta? de Pacotes com M??plos Caminhos Baseado em OpenFlow", Technical Report, Electrical Engineering Program, COPPE/UFRJ, December 2012. Portugues, A4 size, 14 p., file: MMFD12.pdf

GTA-12-30 Mattos, D. M. F., Ferraz, L. H. G., Duarte, O. C. M. B. - "Uma Arquitetura para Isolamento de Redes Virtuais usando uma Abordagem H?ida Xen e OpenFlow", Technical Report, Electrical Engineering Program, COPPE/UFRJ, December 2012. Portugues, A4 size, 14 p., file: MFD12.pdf

GTA-12-31 Moreira, M. D. D., Laufer, R. P., Velloso, P. B., and Duarte, O. C. M. B. - "Capacity and Robustness Tradeoffs in Bloom Filters for Distributed Applications", in IEEE Transactions on Parallel and Distributed Systems, vol. 23, no. 12, pp. 2219-2230, December, 2012. English, A4 size, 12 p., file: DLVD12.pdf

GTA-12-32 Mattos, D. M. F., and Duarte, O. C. M. B. - "Isolamento e Oferta de Qualidade de Servi?em Redes Virtualizadas", in Revista Brasileira de Redes de Computadores e Sistemas Distribuidos (RESD), vol. 5, no. 1, pp. 45-55, June, 2012, ISSN: 1983-4217. Portuguese, A4 size, 11 p., file: MaDu12a.pdf

Technical Reports — 2011

GTA-11-01 Augusto, C. H. P., Carvalho, C. B., Silva, M. W. R., and de Rezende, J. F. - “REUSE: A Combined Routing and Link Scheduling Mechanism for Wireless Mesh Networks”, in Computer Communications (COMCOM), 2011. ISSN: 0140-3664, vol. 34, no. 18, pp. 2207-2216, December 2011. English, A4 size, 10 p., file: ACS11.pdf

GTA-11-02 Ribeiro, F. J. L., Pedroza, A. C. P. and Costa, L. H. M. K. - “Deepwater Monitoring System Using Logistic-Support Vessels in Underwater Sensor Networks”, accepted in The 21st International Offshore (Ocean) and Polar Engineering Conference, Maui, Hawaii, USA, June 2011. Portuguese, A4 size, 7 p., file: RPC11.pdf

GTA-11-03 Campista, M. E. M., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Improving the multiple access method of home networks over the electrical wiring”, in Annals of Telecommunications, Springer, ISSN 0003-4347, DOI 10.1007/s12243-011-0247-x, Vol. 66, n. 11-12, pages 683-694, 2011. English, A4 size, 18 p., file: CCD11.pdf

GTA-11-04 Couto, R. S., Campista, M. E. M., Costa, L. H. M. K. - “XTC: Um Controlador de Vaz?para Roteadores Virtuais Baseados em Xen”, in XXIX Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?os - SBRC’2011, pp. 17-30, Campo Grande, MS, Brazil, May 2011. Portuguese, A4 size, 14 p., file: CCC11.pdf

GTA-11-05 Carvalho, H. E. T., Fernandes, N. C., Duarte, O. C. M. B. - “Um Controlador Robusto de Acordos de N?el de Servi?para Redes Virtuais Baseado em L??a Nebulosa”, to appear in XXIX Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?os - SBRC’2011, Campo Grande, MS, Brazil, May 2011. Portuguese, A4 size, 14 p., file: CFD11.pdf

GTA-11-06 Mattos, D. M. F., Fernandes, N. C., Duarte, O. C. M. B. - “XenFlow: Um Sistema de Processamento de Fluxos Robusto e Eficiente para Migra? em Redes Virtuais”, to appear in XXIX Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?os - SBRC’2011, Campo Grande, MS, Brazil, May 2011. Portuguese, A4 size, 14 p., file: MFD11.pdf

GTA-11-07 Fernandes, N. C., Duarte, O. C. M. B. - “Provendo Isolamento e Qualidade de Servi?em Redes Virtuais”, to appear in XXIX Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?os - SBRC’2011, Campo Grande, MS, Brazil, May 2011. Portuguese, A4 size, 14 p., file: FD11.pdf

GTA-11-08 Ferraz, L., Fernandes, N. C., Velloso, P. B., Duarte, O. C. M. B. - “Um Mecanismo de Exclus?Acurado baseado em Confian?para Controle de Acesso em Redes Ad Hoc”, to appear in XXIX Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?os - SBRC’2011, Campo Grande, MS, Brazil, May 2011. Portuguese, A4 size, 14 p., file: FFVD11.pdf

GTA-11-09 Fragni, C., Costa, L. H. M. K., Duarte, O. C. M. B. - “ECO-ALOC: Aloca? Energeticamente Eficiente de Recursos para Roteadores de Software em Cluster”, to appear in XXIX Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?os - SBRC’2011, Campo Grande, MS, Brazil, May 2011. Portuguese, A4 size, 14 p., file: FCD11.pdf

GTA-11-10 Moreira, M. D. D., Laufer, R. P., Fernandes, N. C., Duarte, O. C. M. B. - “A Stateless Traceback Technique for Identifying the Origin of Attacks from a Single Packet”, to appear in IEEE International Conference on Communications (ICC 2011 - Communication and Information System Security Symposium (CISS)), Kyoto, Japan, June 2011. English, A4 size, 5 p., file: MLFD11.pdf GTA-11-11 Fernandes, N. C., Duarte, O. C. M. B. - “XNetMon: A Network Monitor for

Securing Virtual Networks”, to appear in IEEE International Conference on Communications (ICC 2011 - Next Generation Networking and Internet Symposium (NGNI)), Kyoto, Japan, June 2011. English, A4 size, 5 p., file: FeDu11.pdf

GTA-11-12 Master Science Thesis Author - Carlo Fragni Title - “ECO-ALOC: Aloca? Energeticamente Eficiente de Recursos para Roteadores de Software em Cluster” Advisor - Lu? Henrique Maciel Kosmalski Costa COPPE/PEE/UFRJ - March 2011 Portuguese - A4 size - 93 p. - File: Carlo11.pdf

GTA-11-13 Ribeiro, F. J. L., Pedroza, A. C. P. and Costa, L. H. M. K. - “Sistema de Monitoramento Subaqu?co Usando Navios de Suporte Log?tico em uma Rede de Sensores Ac??cos Tolerante a Atrasos e Interrup?s”, accepted in XXIX Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?os - SBRC’2011, Campo Grande, MS, Brazil, May 2011. Portuguese, A4 size, 14 p., file: RPC11a.pdf

GTA-11-14 Ribeiro Junior, J. G, Campista, M. E. M., Costa, L. H. M. K. - “Monitoramento Colaborativo de Tr?ito utilizando Redes IEEE 802.11 em Cidades Inteligentes”, in XVI Workshop de Ger?ia e Opera? de Redes e Servi? (WGRS 2011) - SBRC’2011, pp. 91-104, Campo Grande, MS, Brazil, May 2011. Portuguese, A4 size, 14 p., file: RJCC11.pdf

GTA-11-15 Moraes, I. M., Pisa, P. S., Carvalho, H. E. T., Alves, R. S., Ferraz, L. H. G., Ferreira, T. N., Couto, R. S., da Silva Neto, D. J., da Costa, V. P., Lage, R. A., dos Santos, L. V., Fernandes, N. C., Campista, M. E. M., Costa, L. H. M. K., and Duarte, O. C. M. B.

- “VNEXT: Uma Ferramenta de Controle e Gerenciamento para Redes Virtuais Baseadas em Xen”, in Sal?de Ferramentas do XXIX Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?os - SBRC’2011, pp. 981-988, Campo Grande, MS, Brazil, May 2011. Portuguese, A4 size, 8 p., file: MPC11.pdf

GTA-11-16 Mattos, D. M. F., Fernandes, N. C., Cardoso, L. P., da Costa, V. T, Mauricio, L. H., Barreto, F. P. B. M., Portella, A. Y., Moraes, I. M., Campista, M. E. M., Costa, L. H. M. K., and Duarte, O. C. M. B. - “OMNI: Uma Ferramenta para Gerenciamento Aut??o de Redes OpenFlow”, in Sal?de Ferramentas do XXIX Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?os - SBRC’2011, pp. 957-964, Campo Grande, MS, Brazil, May 2011. Portuguese, A4 size, 8 p., file: MFC11.pdf

GTA-11-17 Couto, R. S., Campista, M. E. M., Costa, L. H. M. K. and Duarte, O. C. M. B. - “An Experimental Analysis of Routing Inconsistency in Indoor Wireless Mesh Networks”, in IEEE Symposium on Computers and Communications - ISCC’2011 , pp. 609-614, Corfu, Greece, June 2011. English, A4 size, 6 p., file: CCC11a.pdf

GTA-11-18 Alves, R. S., Campista, M. E. M. and Costa, L. H. M. K. - “Um Servidor de M?inas Virtuais Adaptado a M??plas Pilhas de Protocolos”, in XVI Workshop de Ger?ia e Opera? de Redes e Servi? (WGRS 2011) - SBRC’2011, pp. 3-16, Campo Grande, MS, Brazil, May 2011. Portuguese, A4 size, 14 p., file: ACC11.pdf

GTA-11-19 Guedes, R. M. and de Rezende, J. F. - “Sele? Din?ca de Canais de Controle em R?os Cognitivos Utilizando Reinforcement Learning”, XXIX Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?os - SBRC’2011, pp. 75-88, Campo Grande, MS, Brazil, May 2011. Portuguese, A4 size, 14 p., file: GueRe11.pdf

GTA-11-20 Carvalho, C. B. and de Rezende, J. F. - “Roteamento, Atribui? e Adapta? Conjunta de Largura de Canais em Redes em Malha sem Fio IEEE 802.11”, XXIX Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?os - SBRC’2011, pp. 381-394, Campo Grande, MS, Brazil, May 2011. Portuguese, A4 size, 14 p., file: CaRe11.pdf

GTA-11-21 Silva, M. W. R. and de Rezende, J. F. - “Redes de R?os Cognitivos com Disponibilidade Din?ca de Oportunidades”, XXIX Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?os - SBRC’2011, pp. 61-74, Campo Grande, MS, Brazil, May 2011. Portuguese, A4 size, 14 p., file: SiRe11.pdf

GTA-11-22 Mendes, A. C., Augusto, C. H. P., Silva, M. W. R., Guedes, R. M. and de Rezende, J. F. - “Sele? da Ordem de Sensoreamento de Canais em uma Rede Cognitiva Oportunista”, I Workshop on Broadband Access Networks (WRA 2011)- SBRC’2011, pp. 75-88, Campo Grande, MS, Brazil, May 2011. Portuguese, A4 size, 14 p., file: MASRR11.pdf

GTA-11-23 Coutinho, P. S. and de Rezende, J. F. - “Detec? de Energia para R?os Cognitivos Usando GNU Radio e USRP2”, I Workshop on Broadband Access Networks (WRA 2011)- SBRC’2011, pp. 61-74, Campo Grande, MS, Brazil, May 2011. Portuguese, A4 size, 14 p., file: CoRe11.pdf

GTA-11-24 Pires, A. A. and de Rezende, J. F. - “Aumento do Reuso Espacial em Redes Ad Hoc IEEE 802.11 com o uso de Enlaces Independentes”, XXIX Simp?? Brasileiro de Redes de Computadores e Sistemas Distribu?os - SBRC’2011, pp. 353-366, Campo Grande, MS, Brazil, May 2011. Portuguese, A4 size, 14 p., file: PiRe11.pdf

GTA-11-25 Laufer, R. P., Velloso, P. B., and Duarte, O. C. M. B. - “A Generalized Bloom Filter to Secure Distributed Network Applications”, Computer Networks, Vol. 55, No. 8, pp. 1804-1819, June 2011. English, Letter, 16 p., file: RVD11.pdf

GTA-11-26 Undergraduate Final Project Author - Hugo Eiji Tibana Carvalho Title - “Controle de Acordos de N?el de Servi?em Redes Virtuais Baseado em L??a Nebulosa" Advisors - Otto Carlos M. B. Duarte and Natalia Castro Fernandes DEL/POLI/UFRJ - March 2011 Portuguese - A4 size - 113 p. - File: Hugo11.pdf

GTA-11-27 Undergraduate Final Project Author - Pedro Silveira Pisa Title - “ADAGA: sistemA de Detec? de Anomalias para o Gerenciamento Aut??o de redes virtuais" Advisors - Otto Carlos M. B. Duarte and Natalia Castro Fernandes DEL/POLI/UFRJ - March 2011 Portuguese - A4 size - 102 p. - File: Pisa11.pdf

GTA-11-29 Undergraduate Final Project Author - Pedro Smith Coutinho Title - “Detec? de Energia para R?os Cognitivos usando GNU Radio e USRP2" Advisors - Jos?erreira de Rezende DEL/POLI/UFRJ - March 2011 Portuguese - A4 size - 88 p. - File: Coutinho11.pdf

GTA-11-30 Mendes, A.C, Augusto, C. H. P., Silva, M. W. R., Guedes, R. M., and de Rezende, J. F. - “Channel Sensing Order for Cognitive Radio Networks Using Reinforcement Learning”, accepted in 36th IEEE Conference on Local Computer Networks - LCN’2011, Bonn, Germany, October 2011. English, A4 size, 8 p., file: MASG11.pdf

GTA-11-31 Pisa, P. S., Couto, R. S., Carvalho, H. E. T., Neto, D. J. S., Fernandes, N. C., Campista, M. E. M., Costa, L. H. M. K., Duarte, O. C. M. B., and Pujolle, G. - “VNEXT: Virtual NEtwork management for Xen-based Testbeds”, in 2nd IFIP International Conference Network of the Future - NoF’2011, Paris, France, November 2011. English, A4 size, 5 p., file: PCC11.pdf

GTA-11-32 Undergraduate Final Project Author - Manuela Ferreira de Lima Title - “Utiliza? de Rede de Sensores Zigbee para Monitora? e Controle Dom?ico" Advisor - Lu? Henrique Maciel Kosmalski Costa DEL/POLI/UFRJ - August 2011 Portuguese - A4 size - 54 p. - File: Lima11.pdf

GTA-11-33 Mattos, D. M. F., Fernandes, N. C., da Costa, V. T, Cardoso, L. P., Campista, M. E. M., Costa, L. H. M. K., and Duarte, O. C. M. B. - “OMNI: OpenFlow MaNagement Infrastructure”, in 2nd IFIP International Conference Network of the Future - NoF’2011, Paris, France, November 2011. English, A4 size, 5 p., file: MFC11-2.pdf

GTA-11-34 Undergraduate Final Project Author - Rodrigo de Souza Couto Title - “XTC: Um Controlador de Vaz?para Roteadores Virtuais Baseados em Xen" Advisors - Lu? Henrique Maciel Kosmalski Costa and Miguel Elias Mitre Campista DEL/POLI/UFRJ - August 2011 Portuguese - A4 size - 70 p. - File: Couto11.pdf

GTA-11-35 Couto, R. S, Campista, M. E. M., Costa, L. H. M. K - “XTC: A Throughput Control Mechanism for Xen-based Virtualized Software Routers”, in IEEE Global Communications Conference - GLOBECOM’2011, Houston, Texas, USA, December 2011. English, A4 size, 7 p., file: CCC11b.pdf

GTA-11-36 Doctor Science Thesis Author - Carlos Henrique Pereira Augusto Title - “Controle de Inundação em Redes Ad Hoc Móveis" Advisor - José Ferreira de Rezende COPPE/PEE/UFRJ - June 2011 Portuguese, A4 size, 86 p. - Directory: Augusto11

GTA-11-37 Doctor of Science Thesis Author - Natalia Castro Fernandes Title - “T?icas de Virtualiza? e Autoconfigura? para o Projeto de Redes de Nova Gera?" Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - September 2011 Portuguese, A4 size, 162 p. - Directory: Natalia11

GTA-11-38 Ribeiro, F. J. L., Pedroza, A. C. P. and Costa, L. H. M. K. - “Deepwater Monitoring System in Underwater Delay/Disruption Tolerant Network”, accepted in 10th International Information and Telecommunication Technologies Symposium - I2TS’2011, Florian??is, SC, Brazil, Dec. 2011. Portuguese, A4 size, 8 p., file: RPC11b.pdf

GTA-11-39 Master Science Thesis Author - Tib?o Menezes de Oliveira Title - “Virtualiza? de Redes IEEE 802.11 para Balanceamento de Carga” Advisor - Jos?erreira de Rezende COPPE/PEE/UFRJ - September 2011 Portuguese - A4 size - 84 p. - Directory: Oliveira11

GTA-11-40 Meirelles, A. L. S., Cardoso, K. V. and de Rezende, J. F. - “A Strategy to Improve Sensing Accuracy of Energy Detection for Distributed Spectrum Management Systems”, in 7th Latin American Network Operations and Management Symposium - LANOMS’2011, Quito, Ecuador, October 2011. English, Letter size, 8 p., file: MCR11.pdf

GTA-11-41 Carvalho, C., and de Rezende, J. F. - “Roteamento, Atribui? e Adapta? Conjunta de Largura de Canais em Redes em Malha sem Fio IEEE 802.11”, in Revista Brasileira de Redes de Computadores e Sistemas Distribu?os (RESD), Vol. 4, no. 1, pp. 9-17, Junho 2011, ISSN: 1983-4217. Portuguese, A4 size, 9 p., file: CaRe11a.pdf

GTA-11-42 Doctor of Science Thesis Author - Marcel William Rocha da Silva Title - “Redes Cognitivas com Oportunidades Din?cas de Acesso ao Espectro” Advisor - Jos?erreira de Rezende COPPE/PEE/UFRJ - November 2011 Portuguese, A4 size, 82 p. - Directory: Marcel11

Technical Reports — 2010

GTA-10-01 Campista, M. E. M., Ferraz, L. H. G., Moraes, I. M., Lanza, M. L. D., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Interconex?de Redes na Internet do Futuro: Desafios e Solu?s”, in Minicursos do Simp?? Brasileiro de Redes de Computadores - SBRC’2010, pp. 47-101, Gramado, RS, Brazil, May 2010. Portuguese, A4 size, 55 p., file: CFM10.pdf, presentation: CFM10-slides.pdf GTA-10-02 Moraes, I. M. and Duarte, O. C. M. B. - “Using the Network Load for Admission

Control in OBS Networks: a Multilink Approach”, Journal of Optical Communications and Networking, ISSN 1943-0620, vol. 2, no. 3, pp. 137-147, March 2010. English, Letter size, 10 p., file: MoDu10a.pdf

GTA-10-03 Moraes, I. M. and Duarte, O. C. M. B. - “Sele? de Parceiros em Sistemas Para-Par de V?eo sob Demanda”, XXVIII Simp?? Brasileiro de Redes de Computadores - SBRC’2010, pp. 219-232, Gramado, RS, Brazil, May 2010. Portuguese, A4 size, 14 p., file: MoDu10b.pdf

GTA-10-04 Undergraduate Final Project Author - Lyno Henrique Gon?ves Ferraz Title - “Uma Avalia? do Protocolo HIP para Provis?de Mobilidade na Internet" Advisor - Otto Carlos M. B. Duarte DEL/POLI/UFRJ - March 2010 Portuguese - A4 size - 91 p. - Directory: Lyno10

GTA-10-05 Undergraduate Final Project Author - Rafael dos Santos Alves Title - “Um Servidor de M?inas Virtuais Adaptado a M??plas Pilhas de Protocolos" Advisors - Lu? Henrique M. K. Costa and Miguel Elias M. Campista DEL/POLI/UFRJ - March 2010 Portuguese - A4 size - 60 p. - file: Rafael10.pdf

GTA-10-06 Master Science Thesis Author - Diorgenes Penteado Title - “Redes Ac??cas Subaqu?cas na Monitora? de Correntes Mar?imas” Advisor - Aloysio de Castro Pinto Pedroza e Lu? Henrique Maciel Kosmalski Costa COPPE/PEE/UFRJ - March 2010 Portuguese - A4 size - 70 p. - File: Diorgenes10.pdf

GTA-10-07 Penteado, D., Costa, L. H. M. K., and Pedroza, A. C. P. - “Deep-ocean Data Acquisition Using Underwater Sensor Networks”, to appear in The International Society of Offshore and Polar Engineers - ISOPE-2010, Beijing, China, June 2010. English, A4 size, 7 p., file: PCP10.pdf

GTA-10-08 Cavalcanti, S. R., Campista, M. E. M., Abdesslem, F. B., Costa, L. H. M. K., Amorim, M. D., and Duarte, O. C. M. B. - “Veer: A Trajectory-Based Peer Selection Algorithm for Networks of Vehicles”, to appear in The 9th IFIP Annual Mediterranean Ad Hoc Networking Workshop - MedHocNet, Juan les Pins, France, June 2010. English, A4 size, 7 p., file: CCA10.pdf

GTA-10-09 Pisa, P. S., Fernandes, N. C., Carvalho, H. E. T., Moreira, M. D. D., Campista, M. E. M., Costa, L. H. M. K., and Duarte, O. C. M. B. - “OpenFlow and Xen-Based Virtual Network Migration”, to appear in The World Computer Congress 2010 - Network of the Future Conference, Brisbane, Australia, September 2010. English, A4 Size, 12p., file: PFC10.pdf

GTA-10-10 Fernandes, N. C., Moreira, M. D. D., Moraes, I. M., Ferraz, L. H. G., Couto, R. S., Carvalho, H. E. T., Campista, M. E. M., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Virtual Networks: Isolation, Performance, and Trends”, in the Annals of Telecommunications, Springer Paris, ISSN 0003-4347, DOI 10.1007/s12243-010-0208-9, Vol. 66, n. 5-6, pages 339-355, 2011. English, A4 size, 23 p., file: FMM10.pdf

GTA-10-11 Mattos, D. M. F., Fragni, C., Moreira, M. D. D., Ferraz, L. H. G., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Evaluating Virtual Router Performance for the Future Internet”, Technical Report, Electrical Engineering Program, COPPE/UFRJ, June 2010. English, A4 size, 8 p., file: MFM10.pdf

GTA-10-12 Augusto, C. H. P. and de Rezende, J. F. - “Escalonamento Distribu??o de Inunda?s em Redes Ad Hoc M??s”, XXVIII Simp?? Brasileiro de Redes de Computadores - SBRC’2010, pp. 857-870, Gramado, RS, Brazil, May 2010. Portuguese, A4 size, 14 p., file: AuRe10.pdf

GTA-10-13 Augusto, C. H. P., Carvalho, C. B., da Silva, M. W. R. and de Rezende, J. F. - “Impacto do Roteamento no Escalonamento de Enlaces em Redes em Malha Sem Fio”, XXVIII Simp?? Brasileiro de Redes de Computadores - SBRC’2010, pp. 189-202, Gramado, RS, Brazil, May 2010. Portuguese, A4 size, 14 p., file: ACS10.pdf

GTA-10-14 Carvalho, C. B. and de Rezende, J. F. - “Roteamento em Redes em Malha Sem Fio IEEE 802.11 com Adapta? de Largura de Canal”, XXVIII Simp?? Brasileiro de Redes de Computadores - SBRC’2010, pp. 161-174, Gramado, RS, Brazil, May 2010. Portuguese, A4 size, 14 p., file: CaRe10.pdf

GTA-10-15 Master Science Thesis Author - Andr?haves Mendes Title - “Detec? Cooperativa em Redes Cognitivas para o Controle da Interefrência em Sistemas Licenciados” Advisor - Jos?erreira de Rezende COPPE/PEE/UFRJ - June 2010 Portuguese - A4 size - 129 p. - Directory: Mendes10

GTA-10-16 Augusto, C. H. P. and de Rezende, J. F. - - “Distributed Broadcast Scheduling in Mobile Ad Hoc Networks”, 9th IFIP Annual Mediterranean Ad Hoc Networking Workshop - MedHocNet, Juan les Pins, France, June 2010. English, A4 size, 8 p., file: AuRe10a.pdf

GTA-10-17 Carvalho, C. B. and de Rezende, J. F. - “Routing in IEEE 802.11 Wireless Mesh Networks with Channel Width Adaptation”, 9th IFIP Annual Mediterranean Ad Hoc Networking Workshop - MedHocNet, Juan les Pins, France, June 2010. English, A4 size, 8 p., file: CaRe10a.pdf

GTA-10-18 Pires, A. A., and de Rezende, J. F. - “Mecanismo de Difus?de Padr??de Interfer?ia para o Aumento da Capacidade de Redes Ad Hoc IEEE 802.11”, in IX Workshop em Desempenho de Sistemas Computacionais e de Comunica? - WPerformance’2010 (XXX Congresso da Sociedade Brasileira de Computa? - CSBC 2010), pp. 1859-1872, Belo Horizonte, MG, Brazil, July 2010. Portuguese, A4 size, 14p.

GTA-10-19 Carvalho, C. B., and de Rezende, J. F. - “Geographic Routing in IEEE 802.11 Wireless Mesh Networks with Channel Width Adaptation”, in 7th International Telecommunications Symposium - ITS’2010, Manaus, AM, Brazil, September 2010. English, A4 size, 5 p.

GTA-10-20 Augusto, C. H. P., Carvalho, C. B., Silva, M. W. R., and de Rezende, J. F. - “The Impact of Joint Routing and Link Scheduling on the Performance of Wireless Mesh Networks”, in 35th IEEE Conference on Local Computer Networks - LCN’2010, Denver, CO, USA, October 2010. English, A4 size, 8 p., file: ACS10a.pdf

GTA-10-21 Silva, M. W. R., and de Rezende, J. F. - “Routing in Hybrid Mesh Networks of Cognitive and IEEE 802.11 Radios”, in IEEE Globecom 2010 - Selected Areas in Communications Symposium (GC’10 SAC), Miami, MI, USA, December 2010. English, A4 size, 5 p., file: SiRe10.pdf

GTA-10-22 Velloso, P. B., Laufer, R. P., Cunha, D. O., Duarte, O. C. M. B., and Pujolle, G. - "Trust Management in Mobile Ad Hoc Networks Using a Scalable Maturity-Based Model”, in IEEE Transactions on Network and Service Management, ISSN 1932-4537, Vol. 7, no. 3, pp. 172-185, September 2010. English, Letter size, 14 p. file: VLC10.pdf

GTA-10-23 Fernandes, N. C. and Duarte, O. C. M. B. - “XNetMon: Uma Arquitetura com Seguran?para Redes Virtuais”, in Anais do X Simp?? Brasileiro em Seguran?da Informa? e de Sistemas Computacionais - SBSeg’10, pp. 339-352, Fortaleza, CE, Brazil, October 2010. Portuguese, A4 size, 14 p. file: FeDu10.pdf

GTA-10-24 Oliveira, T. M., Silva, M. W. R., Cardoso, K. V., and de Rezende, J. F. - “Virtualization for Load Balancing on IEEE 802.11 Networks”, in 7th International ICST Conference on Mobile and Ubiquitous Systems: Computing, Networking and Services - MobiQuitous’2010, Sydney, AU, December 2010. English, A4 size, 11 p., file: OSC10.pdf

GTA-10-25 Fernandes, N. C. and Duarte, O. C. M. B. - “A Lightweight Group-Key Management Protocol for Secure Ad-Hoc-Network Routing”, in Computer Networks, ISSN 1389-1286, Vol. 55, Number 3, pages 759-778, Elsevier North-Holland, Inc., New York, NY, USA, February 2011. English, A4 size, 44 p., file: FeDu10-02.pdf

GTA-10-26 Ribeiro, F. J. L., Pedroza, A. C. P. and Costa, L. H. M. K. - “Protocolo de Roteamento para Sistemas Subaqu?cos de Comunica?”, in 9th International Information and Telecommunication Technologies Symposium - I2TS’2010, Rio de Janeiro, Brazil, December, 2010. Portuguese, Letter size, 7 p., file: RPC10.pdf

GTA-10-27 Moreira, M. D. D., Laufer, R. P., Fernandes, N. C. and Duarte, O. C. M. B. - “Uma T?ica de Rastreamento sem Estado para Identificar a Origem de Ataques a Partir de um ?ico Pacote”, in Anais do X Simp?? Brasileiro em Seguran?da Informa? e de Sistemas Computacionais - SBSeg’10, pp. 325-338, Fortaleza, CE, Brazil, October 2010. Portuguese, A4 size, 14 p. file: MLF10.pdf

GTA-10-28 Belblidia, N., Amorim, M. D., Costa, L. H. M. K., Leguay, J., and Conan, V. - “PACS: Chopping and shuffling large contents for faster opportunistic dissemination”, in The 8th International Conference on Wireless On-demand Network Systems and Services - WONS 2011, pages 9-16, Bardonecchia, Italy, January 2011. English, A4 size, 8 p. file: BAC11.pdf GTA-10-29 Passos, D. G., de Albuquerque, C. V. N., Campista, M. E. M., Costa, L. H. M.

K., Duarte, O. C. M. B. - “Minimum loss multiplicative routing metrics for wireless mesh networks”, in Journal of Internet Services and Applications, Springer, ISSN: 1867-4828, vol. 1, no. 3, pp. 201-214, February 2011. English, A4 size, 15 p. file: PAC11.pdf

GTA-10-30 Augusto, C. H. P., Carvalho, C. B., Silva, M. W. R., and de Rezende, J. F. - “Escalonamento de Enlaces com Reuso Espacial em Redes em Malha Sem Fio”, in Revista Brasileira de Redes de Computadores e Sistemas Distribu?os (RESD), SBC, ISSN: 1983-4217, vol. 3, no. 2, pp. 9-19, December 2010. Portuguese, A4 size, 11 p., file: ACS10b.pdf

Technical Reports — 2009

GTA-09-01 Undergraduate Final Project Author - Marcelo Duffles Donato Moreira Title - “Implementa? e Avalia? de um Sistema de Rastreamento de Pacotes IP” Advisor - Otto Carlos M. B. Duarte and Rafael P. Laufer DEL/POLI/UFRJ - February 2009 Portuguese - A4 size - 68 p. - File: marcelo09a.pdf

GTA-09-02 Doctor of Science Thesis Author - Kleber Vieira Cardoso Title - “Controle Autom?co de Taxa de Transmiss?em Redes IEEE 802.11” Advisor - Jos?erreira de Rezende PEE/COPPE/UFRJ - February 2009 Portuguese - A4 size - 152 p. - Directory: Cardoso09

GTA-09-03 Undergraduate Final Project Author - Carlo Fragni Title - “Estudo Comparativo de T?icas de Virtualiza?” Advisor - Otto Carlos M. B. Duarte DEL/POLI/UFRJ - February 2009 Portuguese - A4 size - 60 p. - File: carlo09.pdf

GTA-09-04 Moreira, M. D. D., Fernandes, N. C., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Internet do Futuro: Um Novo Horizonte”, in Minicursos do Simp?? Brasileiro de Redes de Computadores - SBRC’2009, pp. 1-59, Rio de Janeiro, RJ, Brazil, May 209. Portuguese, A4 size, 59 p., file: MFCD09.pdf

GTA-09-05 Fernandes, N. C., Moreira, M. D., and Duarte, O. C. M. B. - “An Efficient Filterbased Addressing Protocol for Autoconfiguration of Mobile Ad Hoc Networks”, in IEEE INFOCOM 2009, Rio de Janeiro, RJ, Brazil, April 2009. English, A4 size, 9p., file: FMD09.pdf

GTA-09-06 Silva, M. W. R., and de Rezende, J. F. - “Roteamento em Redes em Malha H?ridas de R?os Cognitivos e IEEE 802.11”, in XXVII Simp?? Brasileiro de Redes de Computadores - SBRC’2009, pp. 959-972, Recife, PE, Brazil, May 2009. Portuguese, A4 size, 14 p., file: SR09.pdf

GTA-09-07 Augusto, C. H. P., Silva, M. W. R., Cardoso, K. V., Mendes, A. C., Guedes, R. M., and de Rezende, J. F. - “Implementa? e Avalia? de Encadeamento de Conex??TCP em Redes de Grande Mem?? e Altas Perdas”, in XXVII Simp?? Brasileiro de Redes de Computadores - SBRC’2009, pp. 89-102, Recife, PE, Brazil, May 2009. Portuguese, A4 size, 14 p., file: ASC09.pdf GTA-09-08 Oliveira, T. M., Silva, M. W. R., Cardoso, K. V., and de Rezende, J. F. - “Bal-

anceamento de Carga Utilizando Virtualiza? em Redes IEEE 802.11”, in XXVII Simp?? Brasileiro de Redes de Computadores - SBRC’2009, pp. 511-524, Recife, PE, Brazil, May 2009. Portuguese, A4 size, 14 p., file: OSC09.pdf

GTA-09-09 Augusto, C. H. P., and de Rezende, J. F. - “Inunda? Probabil?tica em Redes Ad Hoc Utilizando Sum?o de Vizinhan?Atrav?de Filtros de Bloom”, in XXVII Simp?? Brasileiro de Redes de Computadores - SBRC’2009, pp. 699-712, Recife, PE, Brazil, May 2009. Portuguese, A4 size, 14 p., file: AR09.pdf

GTA-09-10 Silva, M. W. R., and de Rezende, J. F. - “TDCS: A New Mechanism for Automatic Channel Assignment for Independent IEEE 802.11 Networks”, in The 8th IFIP Annual Mediterranean Ad Hoc Networking Workshop - Med-Hoc-Net’2009, pp. 27-33, Haifa, Israel, June 2009. Portuguese, A4 size, 7 p., file: SiRe09a.pdf

GTA-09-11 Cardoso, K. V., and de Rezende, J. F. - “Accurate Hidden Markov Modeling of Packet Losses in Indoor 802.11 Networks”, in IEEE Communication Letters, IEEE Press, ISSN: 1089-7798. vol. 13, no. 6, pp. 417-419, June 2009. English, Letter size, 3p.

GTA-09-12 Moraes, I. M and Duarte, O. C. M. B. - “Um Mecanismo de Sele? de Pares Baseado no Tempo de Vida dos Participantes em Sistemas de V?eo Par-a-Par”, in XXVII Simp?? Brasileiro de Redes de Computadores - SBRC’2009, pp. 715-728, Recife, PE, Brazil, May 2009. Portuguese, A4 size, 14 p., file: MoDu09.pdf

GTA-09-13 Alves, R. S. A. and Campbell, I. V. and Couto, R. S. and Campista, M. E. M. and Moraes, I. M. and Rubinstein, M. G. and Costa, L. H. M. K. and Duarte, O. C. M. B. and Abdalla, M. - “Redes Veiculares: Princ?ios, Aplica?s e Desafios”, in Minicursos do Simp?? Brasileiro de Redes de Computadores - SBRC’2009, pp. 199-254, Recife, PE, Brazil, May 2009. Portuguese, A4 size, 56 p., file:ACC09.pdf

GTA-09-14 Couto, R. S., Campista, M. E. M., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Uma Avalia? Experimental do Roteamento das Redes em Malha Sem Fio em Ambientes Fechados”, (published as Poster) XXVII Simp?? Brasileiro de Redes de Computadores - SBRC’2009, pp. 983, Recife, PE, Brazil, May 2009. Portuguese, A4 size, 1 p., file: CCC09.pdf

GTA-09-15 Campbell, I. V., Alves, R. S. A., Campista, M. E. M., Costa, L. H. M. K., Duarte, O. C. M. B., and Rubinstein, M. G. - “Um algoritmo para o suporte ?ultrapassagem de ve?ulos usando redes ad-hoc veiculares”, (published as Poster) XXVII Simp?? Brasileiro de Redes de Computadores - SBRC’2009, pp. 982, Recife, PE, Brazil, May 2009. Portuguese, A4 size, 1 p., file: CAC09.pdf

GTA-09-16 Couto, R. S., Campista, M. E. M., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Uma Avalia? Experimental do Roteamento das Redes em Malha Sem-Fio em Ambientes Fechados”, XXVII Simp?? Brasileiro de Telecomunica?s - SBrT’2009, Blumenau, SC, Brazil, October 2009. Portuguese, A4 size, 6 p., file: CCC09a.pdf GTA-09-17 Augusto, C. H. P., and de Rezende, J. F. - “FloorB: Mecanismo de Controle de

Inunda? para Redes Ad Hoc M??”, in Revista Brasileira de Redes de Computadores e Sistemas Distribu?os (RESD), ISSN: 1983-4217, vol. 2, no. 1, pp. 35-49, June 2009. Portuguese, A4 size, 15 p., file: AuRe09a.pdf

GTA-09-18 Mendes, A. C., and de Rezende, J. F. - “Detec? Cooperativa em Redes Cognitivas para o Controle da Interfer?ia em Sistemas Licenciados”, in XXVII Simp?? Brasileiro de Telecomunica?s - SBrT’2009, Blumenau, SC, Brazil, October 2009. Portuguese, A4 size, 6 p., file: MeRe09.pdf

GTA-09-19 Master Science Thesis Author - Marcelo Duffles Donato Moreira Title - “Usando Pontos de Verifica? para Rastrear Ataques de Nega? de Servi?Maci?ente Distribu?os” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - September 2009 Portuguese - A4 size - 79 p. - File: marcelo09b.pdf

GTA-09-20 Rubinstein, M. G., Abdesslem, F. B., Cavalcanti, S. R., Campista, M. E. M. C., Alves, R. S., Costa, L. H. M. K., Amorim, M. D., and Duarte, O. C. M. B. - “Measuring the capacity of in-car to in-car vehicular networks”, in IEEE Communications Magazine, vol. 47, no. 11, pp.128-136, November 2009. English, A4 size, 9 p., file: RAC09.pdf

GTA-09-21 Doctor of Science Thesis Author - Jos?alentim dos Santos Filho Title - “RECOLLVE: Um Ambiente Virtual voltado para a Representa? de Atividades Colaborativas" Advisor - Aloysio de Castro Pinto Pedroza and Jean-Pierre Courtiat COPPE/PEE/UFRJ - March 2009 Portuguese - A4 size - 118 p. File: Santos09.pdf

GTA-09-22 Santos Fo., J. V., Gomes, R. L., Bacurau, R. M., Pedroza, A. C. P. and Cortiat,J.- P. - “Coordinating Collaborative Work with RECOLLVE”, in Proceedings of the 2009 International Symposium on Collaborative Technologies and Systems - CTS’2009, Baltimore, USA, May 2009. English, A4 size, 10 p., file: SGBPC09.pdf

GTA-09-23 Doctor of Science Thesis Author - Filippe Coury Jabour Neto Title - “Um esquema em dois n?eis para suporte a mobilidade em redes de sensores sem fio" Advisor - Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - June 2009 Portuguese - A4 size - 162 p. File: Jabour09.pdf

GTA-09-24 Doctor of Science Thesis Author - Eug?a Cristina M?ler Giancoli Jabour Title - “Um Protocolo de Transporte Colaborativo para Redes de Sensores Sem Fio" Advisor - Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - March 2009 Portuguese - A4 size - 139 p. File: Giancoli09.pdf GTA-09-25 Giancoli, E., Jabour, F. C., and Pedroza, A. C. P. - “Uma Calibragem do Algo-

ritmo de Karn para Redes de Sensores Sem Fio”, In XXXV Confer?ia Latinoamericana de Inform?ca - CLEI’2009, Pelotas, RS, Brazil, Sep 2009. Portuguese - A4 size - 14 p. File: GJP09.pdf

GTA-09-26 Jabour, F. C., Giancoli, E., and Pedroza, A. C. P. - “3M: Um protocolo de enlace multicanal para redes de sensores”, In XXXV Confer?ia Latinoamericana de Inform?ca - CLEI’2009, Pelotas, RS, Brazil, Sep 2009. Portuguese - A4 size - 14 p. File: JabourCLEI.pdf

GTA-09-27 Giancoli, E., Jabour, F. C., and Pedroza, A. C. P. - “CTCP: Analyzing Delivery, Latency and Power Consumption”, to appear in IEEE Latin America, Agosto 2009. Portuguese - A4 size - 8 p. File: GJP09a.pdf

GTA-09-28 Jabour, F. C., Giancoli, E., and Pedroza, A. C. P. - “A multithreaded discreteevent driven mobile sensor network simulator”, In 8th International Information and Telecommunication Technologies Symposium, Florian??is, SC, Brazil, Dec. 2009. English - A4 size - 4 p. File: JabourI2TS09.pdf

GTA-09-29 Penteado, D., Pedroza, A. C. P., and Costa, L. H. M. K. - “Underwater Sensor Networks: a Feasibility Study”, In 8th International Information and Telecommunication Technologies Symposium, Florian??is, SC, Brazil, Dec. 2009. English - A4 size - 3 p. File: PPC09.pdf

GTA-09-30 Fernandes, N. C., Moreira, M. D. D., Duarte, O. C. M. B. - “A Self-Organized Mechanism for Thwarting Malicious Access in Ad Hoc Networks”, to be published in 2010 Conference on Computer Communications (INFOCOM miniconference), San Diego, USA, March, 2010. English, A4 size, 5 p., file: FeMoDu10.pdf

GTA-09-31 Doctor of Science Thesis Author - Igor Monteiro Moraes Title - “Distribui? de V?eo sobre Redes Par-a-Par” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - December 2009 Portuguese - A4 size - 115 p. - Directory: Igor09

GTA-09-32 Moraes, I. M. and Duarte, O. C. M. B. - “A Lifetime-Based Peer Selection Mechanism for Peer-to-Peer Video-on-Demand Systems”, IEEE International Conference on Communications 2010 - ICC’2010, pp. 1-5, Cape Town, South Africa, May 2010. English, A4 size, 5 p., file: MoDu09a.pdf

GTA-09-33 Master Science Thesis Author - Raphael Melo Guedes Title - “Escassez de Recursos em Redes Tolerantes a Atrasos e Interrup?s” Advisor - Jos?erreira de Rezende COPPE/PEE/UFRJ - December 2009 Portuguese - A4 size - 112 p. - Directory: Guedes09

Technical Reports — 2008

GTA-08-01 Master Science Thesis Author - Natalia Castro Fernandes Title - “Controle de Acesso Distribu?o para Redes Ad Hoc” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - February 2008 Portuguese - A4 size - 132 p. - File: natalia08.pdf

GTA-08-02 Master Science Thesis Author - Carina Teixeira de Oliveira Title - “Uma Proposta de Roteamento Probabilístico para Redes Tolerantes a Atrasos e Desconexões” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - March 2008 Portuguese - A4 size - 100 p. - File: carina08.pdf

GTA-08-03 Oliveira, C. T., Braga, R. B., Taveira, D. M., Fernandes, N. C., and Duarte, O. C. M. B. - “A Predicted-contact Routing Scheme for Brazilian Rural Networks”, Technical Report, Electrical Engineering Program, COPPE/UFRJ, March 2008. English, A4 size, 12 p., file: OBT08.pdf

GTA-08-04 Master Science Thesis Author - Reinaldo Bezerra Braga Title - “MAPA: Mecanismo de Avaliação e Punição de nós egoístas em redes Ad hoc” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - March 2008 Portuguese - A4 size - 89 p. - File: Reinaldo08.pdf

GTA-08-05 Master Science Thesis Author - Danilo Michalczuk Taveira Title - “Mecanismo Anti-Spam Baseado em Autentica? e Reputa?” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - March 2008 Portuguese - A4 size - 98 p. - File: Danilo08.pdf

GTA-08-06 Taveira, D. M., Mattos, D. M. F., and Duarte, O. C. M. B. - “Ferramenta para An?se de Caracter?ticas de Spams e Mecanismos Anti-Spam”, in Sal?de Ferramentas do Simp?? Brasileiro de Redes de Computadores - SBRC’2008, pp. 51-58, Rio de Janeiro, RJ, Brazil, May 2008. Portuguese, A4 size, 8 p., file: TMD08.pdf

GTA-08-07 Taveira, D. M., and Duarte, O. C. M. B. - “Ferramenta para An?se de Caracter?ticas de Spams e Mecanismos Anti-Spam”, in 6th IEEE International Workshop on End-to-End Monitoring Techniques and Services - E2EMON 2008, Salvador, BA, Brazil, April 2008. English, A4 size, 8 p., file: TaDu08.pdf

GTA-08-08 Master Science Thesis Author - Natanael Delgado de Freitas Title - “Uma Análise da Utilização de Antenas Direcionais e Múltiplos Caminhos em Redes Ad Hoc Sem Fio” Advisor - Luís Henrique Maciel Kosmalski Costa COPPE/PEE/UFRJ - March 2008 Portuguese - A4 size - 94 p. - File: natanael08.pdf

GTA-08-09 Giancoli, E., Jabour, F. C., and Pedroza, A. C. P. - “Protocolo de Transporte Colaborativo para Redes de Sensores sem Fio”, in IX Workshop de Testes e Toler?ia a Falhas do Simp?? Brasileiro de Redes de Computadores - SBRC’2008, Rio de Janeiro, RJ, Brazil, May 2008. Portuguese, A4 size, 14 p., file: GEJFPA08.pdf

GTA-08-10 Cunha, D. O., Duarte, O. C. M. B., and Pujolle, G. - “Uma An?se do Roteamento em Redes com Diversidade Cooperativa e Canais com Desvanecimento”, in XXVI Simp?? Brasileiro de Redes de Computadores - SBRC’2008, pp. 301-314, Rio de Janeiro, RJ, Brazil, May 2008. Portuguese, A4 size, 14 p., file: CDP08a.pdf

GTA-08-11 Velloso, P. B., Laufer, R. P., Duarte, O. C. M. B., and Pujolle, G. - “An?se de um Modelo de Confian?para Redes Ad Hoc”, in XXVI Simp?? Brasileiro de Redes de Computadores - SBRC’2008, pp. 693-706, Rio de Janeiro, RJ, Brazil, May 2008. Portuguese, A4 size, 14 p., file: VLD08.pdf

GTA-08-12 Moraes, I. M., Campista, M. E. M., Moreira, M. D. D., Rubinstein, M. G., Costa, L. H. M. K., and Duarte, O. C. M. B. - "Distribui? de V?eo sobre Redes Par-a-Par: Arquiteturas, Mecanismos e Desafios”, in Minicursos do Simp?? Brasileiro de Redes de Computadores - SBRC’2008, pp. 115-171, Rio de Janeiro, RJ, Brazil, May 2008. Portuguese, A4 size, 57 p., file: MCM08.pdf

GTA-08-13 Guedes, R. M., Silva, M. W. R., and de Rezende, J. F. - “Escassez de Recursos em Redes Tolerantes a Atrasos e Interrup?s”, in XXVI Simp?? Brasileiro de Redes de Computadores - SBRC’2008, pp. 749-762, Rio de Janeiro, RJ, Brazil, May 2008. Portuguese, A4 size, 14 p., file: GSR08.pdf

GTA-08-14 Cardoso, K. V. and de Rezende, J. F. - “Adapta? Autom?ca de Taxa em Redes 802.11 Densas”, in XXVI Simp?? Brasileiro de Redes de Computadores - SBRC’2008, pp. 467-480, Rio de Janeiro, RJ, Brazil, May 2008. Portuguese, A4 size, 14 p., file: CaRe08.pdf

GTA-08-15 Silva, F. M. and de Rezende, J. F. - “Influ?ia do Ataque do Testemunho Mentiroso nos Modelos de Reputa?”, in XXVI Simp?? Brasileiro de Redes de Computadores - SBRC’2008, pp. 679-692, Rio de Janeiro, RJ, Brazil, May 2008. Portuguese, A4 size, 14 p., file: SiRe08.pdf

GTA-08-16 Augusto, C. H. P., Silva, M. W. R., Cardoso, K. V., Mendes, A. C., Guedes, R. M. and de Rezende, J. F. - “Segmenta? de Conex??TCP para Transfer?ia Fim-a-Fim em Alta Velocidade”, in VII Workshop em Desempenho de Sistemas Computacionais e de Comunica? - WPerformance’2008 (XXVIII Congresso da Sociedade Brasileira de Computa? - CSBC 2008), pp. 141-160, Bel? PA, Brazil, July 2008. Portuguese, A4 size, 20 p., file: ASC08.pdf

GTA-08-17 Velloso, P. B., Laufer, R. P., Duarte, O. C. M. B., and Pujolle, G. - “Analyzing a Human-based Trust Model for Mobile Ad Hoc Networks”, in IEEE Symposium on Computers and Communications ISCC’2008, Marrakech, Marocco, July 2008. English, Letter size, 6 p., file: VLD08a.pdf GTA-08-18 Velloso, P. B., Laufer, R. P., Duarte, O. C. M. B., and Pujolle, G. - “A Trust Model

Robust to Slander Attacks in Ad Hoc Networks”, in Worshop in Advanced Networking and Communications (ANC) jointly with ICCCN’2008, St. Thomas, Virgin Island, August 2008. English, Letter size, 6 p., file: VLD08b.pdf

GTA-08-19 Doctor of Science Thesis Author - Daniel de Oliveira Cunha Title - “Redes sem Fio de M??plos Saltos: Protocolos Espec?icos para Aplica?s e Roteamento com Suporte ?Diversidade Cooperativa” Advisors - Otto Carlos M. B. Duarte and Guy Pujolle COPPE/PEE/UFRJ and LIP6/UPMC - February 2008 Portuguese - A4 size - 173 p. - Directory: Cunha08

GTA-08-20 Master Science Thesis Author - S?o Rodrigues Cavalcanti Title - “VEER: Um Algoritmo de Sele? de Pares em Redes Ad Hoc Veiculares” Advisor - Luís Henrique Maciel Kosmalski Costa COPPE/PEE/UFRJ - March 2008 Portuguese - A4 size - 82 p. - File: savio08.pdf

GTA-08-21 Alves, R. S., Abdesslem, F. B., Cavalcanti, S. R., Campista, M. E. M., Costa, L. H. M. K., Rubinstein, M. G., Amorim, M. D., and Duarte, O. C M. B. - “Uma An?se Experimental da Capacidade de Redes Ad Hoc Veiculares”, in XXVI Simp?? Brasileiro de Telecomunica?s (SBrT’08), Rio de Janeiro, RJ, Brazil, September 2008. Portuguese, A4 size, 8 p., file: AAC08.pdf

GTA-08-22 Moraes, I. M. and Duarte, O. C. M. B. - “Um Mecanismo de Controle de Admiss?Baseado no Grau de Ocupa? para Redes ?ticas”, Revista Brasileira de Redes de Computadores e Sistemas Distribu?os, ISSN 1983-4217, vol. 01, no. 01, pp. 37-50, June 2008. Portuguese, A4 size, 14 p., file: MoDu08.pdf

GTA-08-23 Fernandes, N. C. and Duarte, O. C. M. B. - “Controle de Acesso Auto-Organiz?l e Robusto Baseado em N??elegados para Redes Ad Hoc”, in 8th Brazilian Symposium on Information and Computer System Security (SBSEG’08), Gramado, RS, Brazil, September, 2008. Portuguese, A4 size, 14 p., file:FeDu08a.pdf

GTA-08-24 Fernandes, N. C. and Duarte, O. C. M. B. - "An Efficient Group Key Management for Secure Routing in Ad Hoc Networks”, in IEEE Globecom 2008 Computer and Communications Network Security Symposium (GC’08 CCNS), New Orleans, LA, USA, December, 2008. English, A4 size, 6 p., file:FeDu08b.pdf

GTA-08-25 Campista, M. E. M., Costa, L. H. M. K., and Duarte, O. C. M. B. - “WPR: A Proactive Routing Protocol Tailored to Wireless Mesh Networks”, in IEEE Globecom 2008 Ad Hoc, Sensor and Mesh Networking Symposium (GC’08 AHSN), New Orleans, LA, USA, December, 2008. English, A4 size, 5 p., file: CCD08a.pdf

GTA-08-26 Carrano, R. C., Saade, D. C. M., Campista, M. E. M., Moraes, I. M., de Albuquerque, C. V. N., Magalh?, L. C. S., Rubinstein, M. G., Costa, L. H. M. K., Duarte, O. C. M. B. - “Multihop MAC: IEEE 802.11s Wireless Mesh Networks”, in Encyclopedia on Ad Hoc and Ubiquitous Computing, editors Dharma P Agrawal and Bin Xie, chapter 19, ISBN 9789812833488, Word Scientific Publishing, pp. 501-532, 2008. English, A4 size, 33 p., file: CSC08.pdf

GTA-08-27 Doctor of Science Thesis Author - Pedro Braconnot Velloso Title - “A humanbased trust model for ad hoc networks” Advisors - Otto Carlos M. B. Duarte and Guy Pujolle LIP6/UPMC - July 2008 English - A4 size - 119 p.

GTA-08-28 Moraes, I. M., Laufer, R. P., Cunha, D. O., and Duarte, O. C. M. B. - “An Efficient Admission Control Mechanism for Optical Burst-Switched Networks”, Photonic Network Communications, ISSN 1387-974X, vol. 18, no. 1, pp. 65-76, August 2009. English, A4 size, 12 p., file: MoDu08a.pdf

GTA-08-29 Cunha, D. O., Duarte, O. C. M. B., and Pujolle G. - “A Cooperation-Aware Routing Scheme for Fast Varying Fading Wireless Channels”, to appear in IEEE Communications Letters, ISSN 1089-7798, vol. 12, no. 10, pp. 794-796, October 2008. English, A4 size, 3 p., file:CDP08b.pdf

GTA-08-30 Moraes, I. M., Campista, M. E. M., Duarte, J. L., Passos, D. G., Costa, L. H. M. K., Rubinstein, M. G., de Albuquerque, C. V. N., and Duarte, O. C. M. B. - “On the Impact of User Mobility on Peer-to-Peer Video Streaming”, IEEE Wireless Communications, ISSN 1536-1284, vol. 15, no. 6, pp. 54-62, December 2008. English, A4 size, 17 p, file: MCD08.pdf

GTA-08-31 Esposito, P. M., Campista, M. E. M., Moraes, I. M., Rubinstein, M. G., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Implementing the Expected Transmission Time Metric for OLSR Wireless Mesh Networks”, in IFIP Wireless Days Conference, Dubai, United Arab Emirates, November, 2008. English, A4 size, 5 p., file: ECM08.pdf

GTA-08-32 Giancoli, E., Jabour, F. C., Pedroza, A., - “Collaborative Transport Control Protocol for Sensor Networks”, In: IEEE International Conference on Computer and Electrical Engineering, Phuket, Tail?ia, December, 2008 English, A4 size, 5p., file: GJP08A.pdf

GTA-08-33 Jabour, F. C., Giancoli, E., Pedroza, A., - “Mobility Support for Wireless Sensor Networks”, In: IEEE International Conference on Computer and Electrical Engineering, Phuket, Tail?ia, December, 200. English, A4 size, 5p., file: GJP08B.pdf

GTA-08-34 Giancoli, E., Jabour, F. C., Pedroza, A., - “Reliable Transport Control Protocol for Sensor Networks”, In: 2008 Fourth International Conference on Intelligent Sensors, Sensor Networks and Information Processing, Sydney, Australia, December, 2008. English, A4 size, 6p., file: GJP08C.pdf

GTA-08-35 Jabour, F. C., Giancoli, E., Pedroza, A., - “Mobility Support for Wireless Sensor Networks”, poster in: 3rd IEEE European Conference on Smart Sensing and Context (EuroSSC), Zurich, October, 2008. English, A4 size 2p., A0 size 1p., files: GJP08D.pdf GJP08E.pdf GTA-08-36 Jabour, F. C., Giancoli, E., Pedroza, A., - “Redes de sensores m??s: an?se da

velocidade, comunica? e esfor? computacional.” In: 7th International Information and Telecommunication Technologies Symposium (I2TS), Foz do Igua? Brazil, Deicember, 2008. Portuguese, A4 size 8p., files: GJP08F.pdf

GTA-08-37 Giancoli, E., Jabour, F. C., Pedroza, A., - “CTCP: Analisando a Entrega, Lat?ia e Consumo de Energia.” In: 7th International Information and Telecommunication Technologies Symposium (I2TS), Foz do Igua? Brazil, December, 2008. Portugu? A4 size 8p., files: GJP08G.pdf

GTA-08-38 Silva, M. W. R., Cardoso, K. V., Moura, D. F. C., Galdino, J. F., Salles, R. M., de Rezende, J. F., - “Avalia? do Compartilhamento Espectral em Redes Cognitivas Utilizando Teoria dos Jogos”, in XXVI Simp?? Brasileiro de Telecomunica?s (SBrT’08), Rio de Janeiro, RJ, Brazil, September 2008. Portuguese, A4 size, 6 p., file: SCM08.pdf

GTA-08-39 Doctor of Science Thesis Author - Miguel Elias Mitre Campista Title - “Um Novo Protocolo de Roteamento para Redes em Malha Sem Fio” Advisors - Lu? Henrique M. K. Costa and Otto Carlos M. B. Duarte PEE/COPPE/UFRJ - December 2008 Portuguese - A4 size - 107 p. - Directory: Campista08

GTA-08-40 Master Science Thesis Author - Fabiana Martins da Silva Title - “Avalia? de Desempenho de M?dos Matem?cos usados nos Modelos de Reputa? de Incentivo ?Coopera?" Advisor - Jos?erreira de Rezende PEE/COPPE/UFRJ - October 2008 <br> Portuguese - A4 size - 140 p. - File:

GTA-08-41 Santos Fo., J. V., Gomes, R. L., Pedroza, A. C. P. and Cortiat,J.-P. - “RECOL-LVE: REpresenting COLLaboration in Virtual Environments”, in Proceedings of World-Comp Congress in Computer Science, Computer Engineering, and Applied Computing - CGVR’2008, Las-Vegas, USA, July 2008. English, A4 size, 7 p., file: SGPC08.pdf

GTA-08-42 Santos Fo., J. V., Gomes, R. L., Pedroza, A. C. P. and Cortiat,J.-P. - “The Agent-based Architecture of RECOLLVE”, in Anais do V Simp?? Brasileiro de Sistemas Colaborativos , Vila Velha-ES, Brasil, Oct 2008 English, A4 size, 12 p., file: SGPC08b.pdf

GTA-08-43 Santos Fo., J. V., Gomes, R. L., Pedroza, A. C. P. and Cortiat,J.-P. - “Virtual Interface for Collaborative Work with RECOLLVE”, in Proceedings of the 7th International Information and Telecommunication Technologies Symposium, Foz do Igua?PR, Brazil, Dec. 2008 English, A4 size, 7 p., file: SGPC08c.pdf

Technical Reports — 2007

GTA-07-01 Undergraduate Final Project Author - Felipe Ortigão Sampaio Buarque Schiller Title - “Estudo, Implementação e Análise de Métricas Baseadas na Qualidade do Enlace para o Protocolo OLSR" Advisor - Lu? Henrique Maciel Kosmalski Costa DEL/POLI/UFRJ - January 2007 Portuguese - A4 size - 75 p. - Directory: Felipe07 GTA-07-02 Oliveira, C. T., Moreira, M. D. D., Rubinstein, M. G., Costa, L. H. M. K., Duarte,

O. C. M. B. - “Redes Tolerantes a Atrasos e Desconex??’, in Minicursos do XXV Simp?? Brasileiro de Redes de Computadores - SBRC’2007, Bel? Brazil, PA, May 2007. Portuguese, A4 size, 54 p., file: OlDu07.pdf

GTA-07-03 Esposito, P. M., Schiller, F. O. B., Campista, M. E. M., Moraes, I. M., Rubinstein, M. G., Costa, L. H. M. K., Duarte, O. C. M. B. - “Implementa? da M?ica de Roteamento Tempo Esperado de Transmiss?em Redes em Malha Sem Fio”, in XII Workshop de Ger?ia e Opera? de Redes e Servi? - WGRS’2007, pp. 25-36, Bel? PA, May 2007. Portuguese, A4 size, 12 p., file: ESC07.pdf

GTA-07-04 Cerveira, C. R. e Costa, L. H. M. K. - “Diferencia? de Tr?go e Controle de Admiss?para Redes Ad Hoc IEEE 802.11”, in XXV Simp?? Brasileiro de Redes de Computadores - SBRC’2007, Bel? PA, May 2007. Portuguese, A4 size, 14 p., file: CeCo07.pdf

GTA-07-05 Master Science Thesis Author - Carlos Rodrigo Cerveira Title - “Diferencia? de Tr?go e Controle de Admiss?em Redes Ad Hoc IEEE 802.11” Advisor - Lu? Henrique Maciel Kosmalski Costa COPPE/PEE/UFRJ - March 2007 Portuguese - A4 size - 81 p. - Directory: Cerveira07

GTA-07-06 Master Science Thesis Author - Gabriel Epsztejn Title - “Proposta de uma Arquitetura Par-a-Par Organizada por ?dices” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - March 2007 Portuguese - A4 size - 75 p. - File: Gabriel07

GTA-07-07 Cardoso, K. V. and de Rezende, J. F. - “An?se e Modelagem de Perda de Pacotes em Redes 802.11 em Ambientes Indoor” - in XXV Simp?? Brasileiro de Redes de Computadores - SBRC’2007, pp. 1117-1122, Bel? PA, May 2007. Portuguese, A4 size, 6 p., file: CaRe07.pdf

GTA-07-08 Silva, M. W. R. and de Rezende, J. F. - “SDCD: Um Novo Mecanismo para a Sele? Autom?a de Canal em Redes IEEE 802.11 Independentes” - in XXV Simp?? Brasileiro de Redes de Computadores - SBRC’2007, pp. 621-634, Bel? PA, May 2007. Portuguese, A4 size, 14 p., file: SiRe07.pdf

GTA-07-09 Silva, F. M. and de Rezende, J. F. - “Avalia? de M?dos Matem?cos usados nos Modelos de Reputa? de Incentivo ?Coopera?” - in XXV Simp?? Brasileiro de Redes de Computadores - SBRC’2007, pp. 999-1012, Bel? PA, May 2007. Portuguese, A4 size, 14 p., file: SiRe07a.pdf

GTA-07-10 Master Science Thesis Author - Carlos Henrique Pereira Augusto Title - “Arquitetura Robusta e Escal?l para Descoberta de Servi? em Redes Ad Hoc” Advisor - Jos?erreira de Rezende COPPE/PEE/UFRJ - May 2007 Portuguese - A4 size - 110 p. - Directory: Carlos07 GTA-07-11 Vilela, U. C., Cardoso, K. V., and de Rezende, J. F. - “Redes 802.11 em Cen-

tros Urbanos: Varredura, Estat?ticas e Aplica?s” - in VI Workshop em Desempenho de Sistemas Computacionais e de Comunica? - WPerformance’2007 (XXVII Congresso da Sociedade Brasileira de Computa? - CSBC 2007), pp. 703-718, Rio de Janeiro, RJ, June 2007. Portuguese, A4 size, 16 p., file: VCR07.pdf

GTA-07-12 Cardoso, K. V. and de Rezende, J. F. - “Um Modelo de Markov Oculto para Representa? de Perda de Pacotes em Redes 802.11 em Ambientes Indoor” - in VI Workshop em Desempenho de Sistemas Computacionais e de Comunica? - WPerformance’2007 (XXVII Congresso da Sociedade Brasileira de Computa? - CSBC 2007), pp. 653-672, Rio de Janeiro, RJ, June 2007. Portuguese, A4 size, 20 p., file: CaRe07a.pdf

GTA-07-13 Campista, M. E. M., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Privileging Long-life Packets in Multihop Wireless Networks”, to appear in 9th IFIP/IEEE International Conference on Mobile and Wireless Communications Networks - MWCN’2007, Cork, Ireland, September 2007. English, A4 size, 5 p., file: CCD07.pdf

GTA-07-14 Moreira, M. D. D., Laufer, R. P., Velloso, P. B., and Duarte, O. C. M. B. - “Uma Proposta de Marca? de Pacotes para Rastreamento Robusto a Ataques”, to appear in Anais do Simp?? Brasileiro de Seguran?da Informa? e de Sistemas Computacionais - SBSeg’2007, Rio de Janeiro, RJ, Brazil, August 2007. Portuguese, A4 size, 14 p., file: MLVD07.pdf

GTA-07-15 Moraes, I. M. and Duarte, O. C. M. B. - “Multilink Performance of the Load-Level-Based Admission Control Mechanism for OBS Networks”, in IEEE International Conference on Communications 2008 - ICC’2008, pp. 40-44, Beijing, China, May 2008. English, A4 size, 5 p., file: MoDu07.pdf

GTA-07-16 Cunha, D. O., Duarte, O. C. M. B., and Pujolle, G. - “An Efficient Data Transport Protocol for Event-Driven Field-Estimation on Sensor Networks”, in 9th IFIP/IEEE International Conference on Mobile and Wireless Communications Networks - MWCN’2007, pp. 31-35, Cork, Ireland, September 2007. English, A4 size, 5 p., file: CDP07.pdf

GTA-07-17 Fernandes, N. C. and Duarte, O. C. M. B. - “Autoconfiguracao de Enderecos Baseada em Filtros de Bloom para Redes Ad Hoc”, in XXVI Simp?? Brasileiro de Redes de Computadores - SBRC’2008, pp. 273-286, Rio de Janeiro, RJ, Brazil, May 2008. Portuguese, A4 size, 14 p., file: FeDu08.pdf

GTA-07-18 Amodei Jr., A., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Increasing the throughput of HomePNA”, International Journal of Communication Systems, ISSN 1074-5351, vol. 21, no.7, pp. 695-711, July 2008. English, A4 size, 17 p., file: ACD08.pdf

GTA-07-19 Campista, M. E. M., Passos, D. G., Esposito, P. M., Moraes, I. M., de Albuquerque, C. V. N., Saade, D. C. M., Rubinstein, M. G., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Routing Metrics and Protocols for Wireless Mesh Networks”, in IEEE Network, ISSN 0890-8044, vol. 22, no. 1, pp. 6-12, January 2008. English, A4 size, 30 p., file: CPE08.pdf GTA-07-20 Master Science Thesis

Author - Sergio Torres dos Santos Title - “Redes de Sensores Sem Fio em Monitoramento e Controle" Advisors - Aloysio de Castro Pinto Pedroza Lu? Henrique Maciel Kosmalski Costa COPPE/PEE/UFRJ - June 2007 Portuguese - A4 size - 87 p. - Directory: Sergio07

GTA-07-21 Cunha, D. O., Duarte, O. C. M. B., and Pujolle, G. - “An Enhanced Routing Metric for Fading Wireless Channels”, in IEEE Wireless Communications and Networking Conference - IEEE WCNC 2008 , pp. 2723-2728, Las Vegas, USA, March/April 2008. English, A4 size, 6 p., file: CDP08.pdf

GTA-07-22 Jabour Neto, F. C., Giancoli, E., and Pedroza, A. C. P. - “Um Esquema em Duas Camadas para Suporte ?Mobilidade em Redes de Sensores Sem Fio”, in XXV Simp?? Brasileiro de Telecomunica?s - SBrT’2007, Recife, PE, Setembro 2007. Portuguese - A4 size - 6 p., file: JGP07.pdf

GTA-07-23 Velloso, P. B., Rubinstein, M. G., Duarte, O. C. M. B. - “Evaluating Voice Traffic Requirements on IEEE 802.11 Ad Hoc Networks”, in Annals of Telecommunications, ISSN 0003-4347, vol. 63, no. 5-6, pp. 321-329, June 2008. English, A4 size, 9 p., file: VRD06.pdf

GTA-07-24 Campista, M. E. M., Costa, L. H. M. K., and Duarte, O. C. M. B. - “WPR: Um Protocolo de Roteamento Pr??ivo Adaptado ? Redes em Malha Sem Fio”, in XXVI Simp?? Brasileiro de Redes de Computadores - SBRC’2008, pp. 889-902, Rio de Janeiro, RJ, Brazil, May 2008. Portuguese, A4 size, 14 p., file: CCD08.pdf

GTA-07-25 de Freitas, N. D., Campista, M. E. M., and Costa, L. H. M. K. - “Uma An?se da Utiliza? de Antenas Direcionais e M??plos Caminhos em Redes Ad Hoc Sem Fio”, in XXVI Simp?? Brasileiro de Redes de Computadores - SBRC’2008, pp. 287-300, Rio de Janeiro, RJ, Brazil, May 2008. Portuguese, A4 size, 14 p., file: FCC08.pdf

GTA-07-26 Cavalcanti, S. R., Campista, M. E. M., Abdesslem, F. B., Costa, L. H. M. K., and Amorim, M. D. - “VEER: Um Algoritmo de Sele? de Pares em Redes Ad Hoc Veiculares”, in XXVI Simp?? Brasileiro de Redes de Computadores - SBRC’2008, pp. 315-328, Rio de Janeiro, RJ, Brazil, May 2008. Portuguese, A4 size, 14 p., file: CCA08.pdf

GTA-07-27 Taveira, D. M. and Duarte, O. C. M. B. - “Mecanismo Anti-Spam Baseado em Autentica? e Reputa?”, in XXVI Simpósio Brasileiro de Redes de Computadores - SBRC’2008, pp. 861-874, Rio de Janeiro, RJ, Brazil, May 2008. Portuguese, A4 size, 14 p., file: TaDu07.pdf

GTA-07-28 Oliveira, C. T., Taveira, D. M., Braga, R. B., and Duarte, O. C. M. B. - “Uma Proposta de Roteamento Probabilístico para Redes Tolerantes a Atrasos e Desconexões”, in XXVI Simpósio Brasileiro de Redes de Computadores - SBRC’2008, pp. 735-748, Rio de Janeiro, RJ, Brazil, May 2008. Portuguese, A4 size, 14 p., file: OTB07.pdf GTA-07-29 Braga, R. B., Taveira, D. M., and Duarte, O. C. M. B. - “MAPA: Mecanismo

de Avaliação e Punição de nós egoístas em redes Ad hoc”, in XXVI Simpósio Brasileiro de Redes de Computadores - SBRC’2008, pp. 665-678, Rio de Janeiro, RJ, Brazil, May 2008. Portuguese, A4 size, 14 p., file: BTD07.pdf

Technical Reports — 2006

GTA-06-01 Master Science Thesis Author - Igor Monteiro Moraes Title - “Diferencia? de Servi? em Redes de Comuta? de Rajadas ?ticas" Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - March 2006 Portuguese - A4 size - 81 p. - Directory: Igor06

GTA-06-02 Campista, M. E. M., Moraes, I. M., Esposito, P. M., Amodei Jr., A., Cunha, D. O., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Wireless Ad Hoc Networks on Underserved Communities: An Efficient Solution for Interactive Digital TV”, 8th IFIP IEEE International Conference on Mobile and Wireless Communication Networks - MWCN’2006, pp. 187-198, Santiago, Chile, August 2006. English, Letter size, 12 p., file: CME06.pdf

GTA-06-03 Campista, M. E. M., Moraes, I. M., Esposito, P. M., Amodei Jr., A., Cunha, D. O., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Capacidade do Canal de Interatividade Ad Hoc para a TV Digital”, in XXIV Simp?? Brasileiro de Redes de Computadores - SBRC’2006, pp. 1065-1068, Curitiba, PR, May 2006. Portuguese, A4 size, 4 p., file: CME06a.pdf

GTA-06-04 Cerveira, C.R. and Costa, L. H. M. K. - “A Time-Based Admission Control Mechanism for IEEE 802.11 Ad Hoc Networks”, 8th IFIP IEEE International Conference on Mobile and Wireless Communication Networks - MWCN’2006, pp. 217-228, Santiago, Chile, August 2006. English, Letter size, 12 p., file: CC06.pdf

GTA-06-05 Dos Santos Filho, J. V., Pedroza, A. C. P., and Courtiat, J.-P. “Ger?ia de Direitos de Acesso de Usu?o em Ambientes Virtuais Colaborativos”, VIII Symposium on Virtual Reality - SVR’2006, pp. 187-198, Bel? PA, May 2006. Portuguese, A4 size, 4 p., file: SPC06.pdf

GTA-06-06 Viana, A. C., Amorim, M. D., Viniotis, Y., Fdida S., and de Rezende, J. F. - "Twins: A Dual Addressing Space Representation for Self-organizing Networks”, in IEEE Transactions on Parallel and Distributed Systems, ISSN 1045-9219, vol. 17, n. 12, pp 1468-1481, December 2006. English, A4 size, 24 p.

GTA-06-07 Delicato, F. C., Protti, F., de Rezende, J. F., and Pirmez, L. - “An Efficient Heuristic for Selecting Active Nodes in Wireless Sensor Networks”, in Computer Networks Journal, Elsevier Science, ISSN 1389-1286, vol. 50, n. 18, pp. 3701-3720, December 2006. English, A4 size, 20 p.

GTA-06-08 Freitag, J., Fonseca, N. L. S., and de Rezende, J. F. - "A Tuning Mechanism for 802.11e Access Parameters for Coping with Delay Asymmetry”, in IEEE International Conference on Communications 2006, ICC’2006, Instanbul, Turkey, June 2006. English, Letter size, 5p., file: FFR06.pdf

GTA-06-09 Silva, M. W. R. and de Rezende, J. F. - " Um Mecanismo para a Sele? Din?ca de Canal em Redes IEEE 802.11 Infra-estruturadas”, XXIV Simp?? Brasileiro de Redes de Computadores - SBRC’2006, pp. 359-374, Curitiba, PR, May 2006. Portuguese, A4 size, 15 p., file: SiRe06.pdf

GTA-06-10 Amorim, G. F. and de Rezende, J. F. - "Avalia? de Desempenho da T?ica de Varredura em Roteamento Ad Hoc com Antenas Direcionais”, XXIV Simp?? Brasileiro de Redes de Computadores - SBRC’2006, pp. 36-47, Curitiba, PR, May 2006. Portuguese, A4 size, 12 p., file: AmRe06.pdf

GTA-06-11 Ramella, I. and de Rezende, J. F. - "Evitando Interfer?ias Prim?as em Redes Ad Hoc com Antenas Direcionais”, in XXIV Simp?? Brasileiro de Redes de Computadores - SBRC’2006, pp. 49-63, Curitiba, PR, May 2006. Portuguese, A4 size, 15 p., file: RaRe06.pdf

GTA-06-12 Augusto, C. H. P. and de Rezende, J. F. - "Uma Abordagem Adaptativa e Escal?l para Descoberta de Servi? em Redes Ad Hoc”, in XXIV Simp?? Brasileiro de Redes de Computadores - SBRC’2006, pp. 945-960, Curitiba, PR, May 2006. Portuguese, A4 size, 16 p., file: AuRe06.pdf

GTA-06-13 Delicato, F. C., Pirmez, L., Pires, P. F., and de Rezende, J. F. - "Exploiting Web Technologies to Build Autonomic Wireless Sensor Networks”, in 8th IFIP IEEE International Conference on Mobile and Wireless Communication Networks - MWCN’2006, pp. 99-114, Santiago, Chile, August 2006. English, Letter size, 16 p., file: DPPR06.pdf

GTA-06-14 Augusto, C. H. P. and de Rezende, J. F. - "An Adaptive Approach to Service Discovery in Ad Hoc Networks”, in 8th IFIP IEEE International Conference on Mobile and Wireless Communication Networks - MWCN’2006, pp. 61-75, Santiago, Chile, August 2006. English, Letter size, 15 p., file: AuRe06a.pdf

GTA-06-15 Master Science Thesis Author - Rafael Rosa da Rocha Title - “Um Mecanismo Para Evitar a Surdez Intr?seca em Redes Ad Hoc de M??plos Saltos utilizando Antenas Direcionais" Advisors - Jos?erreira de Rezende and Lu? Henrique Maciel Kosmalski Costa COPPE/PEE/UFRJ - March 2006 Portuguese - A4 size - 119 p. - Directory: Rafael06

GTA-06-16 Master Science Thesis Author - Italo Ramella Title - “A Interfer?ia em Redes Ad Hoc com Antenas Direcionais" Advisor - Jos?erreira de Rezende COPPE/PEE/UFRJ - April 2006 Portuguese - A4 size - 81 p. - Directory: Italo06 GTA-06-17 Epsztejn, G. and Duarte, O. C. M. B. - "Proposta de Rede P2P Organizada por

?dices”, II Workshop de Peer-to-Peer WP2P 2006, Curitiba, PR, May 2006. Portuguese, A4 size, 12 p., file: EpDu06.pdf

GTA-06-18 Velloso, P. B., Laufer, R. P., Duarte, O. C. M. B., and Pujolle, G. - “Um Novo Modelo para Confian?em Rede ad hoc”, XXIV Simp?? Brasileiro de Redes de Computadores - SBRC’2006, pp. 1027-1042, Curitiba, PR, May 2006. Portuguese, A4 size, 16 p., file: VLDP06.pdf

GTA-06-19 Antas, R. B. Z., Fernandes, N. C., Taveira D. M., Campista, M. E. M., Costa, L. H. M. K., and Duarte, O. C. M. B. - "An?se da Qualidade de Voz em uma Rede Ad Hoc Comunit?a”, XI Workshop de Ger?ia e Opera? de Redes e Servi? - WGRS 2006, Curitiba, PR, May 2006. Portuguese, A4 size, 12 p., file: AFT06.pdf

GTA-06-20 Moreira, M. D. D., Coutinho, G. L., Moraes, I. M., Laufer, R. P., and Duarte, O. C. M. B. - "RAT: Implementa? de um Servi?de Rastreamento de Pacotes”, XI Workshop de Ger?ia e Opera? de Redes e Servi? - WGRS 2006, Curitiba, PR, May 2006. Portuguese, A4 size, 12 p., file: MCM06.pdf

GTA-06-21 Freitag, J., Fonseca, N. L. S., and de Rezende, J. F. - "Tuning of 802.11e Network Parameters”, in IEEE Communication Letters, IEEE Press, vol. 10, n. 8, pp. 611-613, ISSN: 1089-7798, August 2006. English, A4 size, 3p.

GTA-06-22 Cunha, D. O., Duarte, O. C. M. B., and Pujolle, G. - "Event-Driven Field Estimation for Wireless Sensor Networks”, 8th IFIP IEEE International Conference on Mobile and Wireless Communication Networks - MWCN’2006, pp. 89-98, Santiago, Chile, August 2006. English, Letter size, 10 p., file: CDP06.pdf

GTA-06-23 Velloso, P. B., Laufer, R. P., Duarte, O. C. M. B., and Pujolle, G. - "HIT: A Human-Inspired Trust Model”, 8th IFIP IEEE International Conference on Mobile and Wireless Communication Networks - MWCN’2006, pp. 35-46, Santiago, Chile, August 2006. English, Letter size, 12 p., file: VLDP06a.pdf

GTA-06-24 Silva, M. W. R. and de Rezende, J. F. - "A Dynamic Channel Allocation Mechanism for IEEE 802.11 Networks”, in VI International Telecommunications Symposium - ITS’2006, pp. 403-408, Fortaleza, CE, September 2006. English, A4 size, 6 p., file: SiRe06a.pdf

GTA-06-25 Undergraduate Final Project Author - Ricardo Gomes Clemente Title - “Uma Solu? de Streaming de V?eo para Celulares: Conceitos, Protocolos e Aplicativo" Advisor - Jos?erreira de Rezende DEL/POLI/UFRJ - July 2006 Portuguese - A4 size - 64 p. Directory: Clemente06

GTA-06-26 Taveira, D. M., Moraes, I. M., Rubinstein, M. G., and Duarte, O. C. M. B. - “T?icas de Defesa Contra Spam”, in Minicursos do Simp?? Brasileiro de Seguran?da Informa? e de Sistemas Computacionais - SBSeg’2006, Santos, SP, Brazil, pp. 202-250, August 2006. Portuguese, A4 size, 49 p., file: TMRD06.pdf

GTA-06-27 Fernandes, N. C., Moreira, M. D. D., Velloso, P. B., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Ataques e Mecanismos de Seguran?em Redes Ad Hoc”, in Minicursos do Simp?? Brasileiro de Seguran?da Informa? e de Sistemas Computacionais - SBSeg’2006, Santos, SP, Brazil, pp. 49-102, August 2006. Portuguese, A4 size, 54 p., file: FMVC06.pdf

GTA-06-28 Rubinstein, M. G., Moraes, I. M., Campista, M. E. M., Costa, L. H. M. K., and Duarte, O. C. M. B. - “A Survey on Wireless Ad Hoc Networks”, in Mobile and Wireless Communications Networks. ISBN: 0-387-34634-1, Springer-Verlag. pp. 1-34, August 2006. English, A4 size, 34 p., file: RMCC06.pdf

GTA-06-29 Delicato, F. C., Pirmez, L., de Rezende, J. F. - "Middleware Baseado em Servi? para Redes de Sensores sem Fio”, in XIX Concurso de Teses e Disserta?s, Anais do XXVI Congresso da SBC, , Campo Grande, MS, ISBN 85-7669-074-8, pp. 41-48, July 2006. English, A4 size, 8 p., file: DPR06.pdf

GTA-06-30 Master Science Thesis Author - Luiz Antonio Ferreira da Silva Title - “An?se de Desempenho de Protocolos de Transporte para Redes de Alta Velocidade" Advisor - Jos?erreira de Rezende COPPE/PEE/UFRJ - August 2006 Portuguese - A4 size - 115 p.- Directory: Luiz06

GTA-06-31 Campista, M. E. M., Moraes, I. M., Esposito, P. M., Amodei Jr., A., Cunha, D. O., Costa, L. H. M. K., and Duarte, O. C. M. B. - “The Ad Hoc Return Channel: a Low-Cost Solution for Brazilian Interactive Digital TV”, in IEEE Communications Magazine, ISSN 0163-6804, vol. 45, no. 1, pp. 136 - 143, January 2007. English, A4 size, 28 p., file: CME06b.pdf

GTA-06-32 Undergraduate Final Project Author - Danilo Michalczuk Taveira Title - “An?se de T?icas de Defesa Contra Spam Advisor - Otto Carlos Muniz Bandeira Duarte DEL/POLI/UFRJ - December 2006 Portuguese - A4 size - 85 p. - File: Danilo06/Danilo06.pdf

GTA-06-33 Campista, M. E. M., Costa, L. H. M. K., and Duarte, O. C. M. B. - ”Um Mecanismo para Privilegiar Pacotes de Vida Longa em Redes Sem Fio de M??plos Saltos”, XXV Simp?? Brasileiro de Redes de Computadores - SBRC’2007, pp. 3-16, Bel? PA, Brazil, May 2007. Portuguese, A4 size, 14 p., file: CCD06a.pdf

GTA-06-34 Campista, M. E. M., Costa, L. H. M. K., and Duarte, O. C. M. B. - ”Um Algoritmo Eficiente de Dissemina? dos Estados de Enlace para Redes em Malha Sem Fio”, XXV Simp?? Brasileiro de Redes de Computadores - SBRC’2007, pp. 751-764, Bel? PA, Brazil, May 2007. Portuguese, A4 size, 14 p., file: CCD06b.pdf GTA-06-35 Moraes, I. M. and Duarte, O. C. M. B. - ”Uma Avalia? Multienlace do Mecanismo

de Controle de Admiss?Baseado no Grau de Ocupa? em Redes ?ticas”, XXV Simp?? Brasileiro de Redes de Computadores - SBRC’2007, pp. 147-160, Bel? PA, Brazil, May 2007. Portuguese, A4 size, 14 p., file: MoDu06.pdf

GTA-06-36 Undergraduate Final Project Author - Natalia Castro Fernandes Title - “An?se de Ataques e Mecanismos de Seguran?em Redes Ad Hoc" Advisor - Otto Carlos Muniz Bandeira Duarte DEL/POLI/UFRJ - December 2006 Portuguese - A4 size - 117 p. - File: Natalia06/pf.pdf

GTA-06-37 Fernandes, N. C. and Duarte, O. C. M. B. - ”CHARADAS: Uma proposta para uso de CHAve de grupo no Roteamento Atrav?de Distribui? Assim?ica Segura”, XXV Simp?? Brasileiro de Telecomunica?s - SBrT’07, Recife, PE, Brazil, September, 2007. Portuguese, A4 size, 6 p., file: FeDu07.pdf

GTA-06-38 Laufer, R. P., Velloso, P. B., Cunha, D. O., Moraes, I. M., Bicudo, M. D. D., Moreira, M. D. D., and Duarte, O. C. M. B. - “Towards Stateless Single-Packet IP Traceback”, 32nd IEEE Conference on Local Computer Networks - LCN’2007, pp. 458-555, Dublin, Ireland, October 2007. English, A4 size, 8 p., file:LVC07.pdf

GTA-06-39 Master Science Thesis Author - Marcel William Rocha da Silva Title - “Aloca? de Canal em Redes Sem Fio IEEE 802.11 Independentes" Advisor - Jos?erreira de Rezende COPPE/PEE/UFRJ - December 2006 Portuguese - A4 size - 83 p. - Directory: Marcel06

GTA-06-40 Master Science Thesis Author - Marcio Elkind Title - “Um Sistema Confi?l de Distribui? de M?ia" Advisor - Jorge Lopes de Souza Le? COPPE/PEE/UFRJ - October 2006 Portuguese - A4 size - 83 p. - Directory: Elkind06

GTA-06-41 Braga, R. B. and Duarte, O. C. M. B. - “T?icas de Avalia? e Puni? ao Mau Comportamento em Redes Ad Hoc”, Technical Report, Electrical Engineering Program, COPPE/UFRJ, December 2006. Portuguese, A4 size, 13 p., file:BrDu06.pdf

GTA-06-42 Oliveira, C. T. and Duarte, O. C. M. B. - “Uma An?se da Probabilidade de Entrega de Mensagens em Redes Tolerantes a Atrasos e Desconex??’, Technical Report, Electrical Engineering Program, COPPE/UFRJ, December 2006. Portuguese, A4 size, 13 p., file:OlDu06.pdf

Technical Reports — 2005

GTA-06-43 Taveira, D. M. and Duarte, O. C. M. B. - “Avali? de Desempenho de Mecanismos Anti-Spam e An?se de Caracter?ticas de Spams”, Technical Report, Electrical Engineering Program, COPPE/UFRJ, December 2006. Portuguese, A4 size, 14 p., file:TaDu06.pdf GTA-05-01 Delicato, F. C., Pires, P. F, Rust, L. F. C. C., Pirmez, L.i, and de Rezende,

J. F. - “Reflective Middleware for Wireless Sensor Networks”, in the 20th Annual ACM Symposium on Applied Computing - ACM SAC’2005, Santa Fe, New Mexico, USA, March 2005. English, A4 size, 5 p., file: DPRPR05.pdf

GTA-05-02 Delicato, F. C., Protti, F., de Rezende, J. F., Rust, L. F. C. C., and Pirmez, L. - “Application-Driven Node Management in Multihop Wireless Sensor Networks”, Lectures Notes in Computer Science, ISSN 0302-9743, Springer-Verlag, to appear in the 4th International Conference on Networking - IEEE ICN’2005, Reunion Island, April 2005. English, A4 size, 8 p., file: DPRRP05.pdf

GTA-05-03 Pires, A. A., de Rezende, J. F., and Cordeiro, C. - “ALCA: A New Scheme for Power Control on 802.11 Ad Hoc Networks”, in the IEEE International Symposium on a World of Wireless, Mobile and Multimedia Networks - IEEE WoWMoM’2005, Taormina, Italy, June 2005. English, A4 size, 3 p., file:

GTA-05-04 Bicudo, M. D. D., Moraes, I. M., Laufer, R. P., Cunha, D. O., Velloso, P. B., and Duarte, O. C. M. B. - “Protection and Minimal Interference in WDM Mesh Networks”, to appear in 12th International Conference on Telecommunications - ICT’2005, Capetown, South Africa, May 2005. English, Letter size, 5 p.

GTA-05-05 Laufer, R. P., Velloso, P. B., and Duarte, O. C. M. B. - “A Stateless Scheme for Single-Packet IP Traceback”, IEEE Infocom 2005 - Student Workshop, Miami, USA, March 2005. English, Letter size, 2 p., file: LVD05a.pdf

GTA-05-06 Laufer, R. P., Velloso, P. B., and Duarte, O. C. M. B. - “Defeating DoS Attacks with IP Traceback”, to appear in IFIP Open Conference on Metropolitan Area Networks - MAN’2005, Ho Chi Minh, Viet Nam, April 2005. English, A4 size, 18 p., file: LVD05b.pdf

GTA-05-07 Laufer, R. P., Velloso, P. B., Cunha, D. O., Moraes, I. M., Bicudo, M. D. D., and Duarte, O. C. M. B. - “A New IP Traceback System against Denial-of-Service Attacks”, 12th International Conference on Telecommunications - ICT’2005, Capetown, South Africa, May 2005. English, Letter size, 5 p., file: LVCM05a.pdf

GTA-05-08 Moraes, I. M. and Duarte, O. C. M. B. - “Differentiated Services in Optical Burst-Switched Networks”, IEEE Infocom 2005 - Student Workshop, Miami, USA, March 2005. English, Letter size, 2 p, file: MoDu05a.pdf.

GTA-05-09 Moraes, I. M., , Cunha, D. O., Bicudo, M. D. D., Laufer, R. P., and Duarte, O. C. M. B. - “An Admission Control Mechanism for Providing Service Differentiation in Optical Burst-Switching Networks”, 12th International Conference on Telecommunications - ICT’2005, Capetown, South Africa, May 2005. English, Letter size, 5 p., file: MCBL05a.pdf GTA-05-10 Cunha, D. O. and Duarte, O. C. M. B. - “Bio-Inspired Data Acquisition in Sensor

Networks”, IEEE Infocom 2005 - Student Workshop, Miami, USA, March 2005. English, Letter size, 2 p,file: CuDu05a.pdf

GTA-05-11 Cunha, D. de O., Laufer, R. P., Moraes, I. M., , Bicudo, M. D. D., Velloso, P. B., and Duarte, O. C. M. B. - “Bio-Inspired Field Estimation with Wireless Sensor Networks”, 12th International Conference on Telecommunications - ICT’2005, Capetown, South Africa, May 2005. English, Letter size, 5 p, file: CLMB05a.pdf

GTA-05-12 A. C. Viana and M. D. Amorim and Y. Viniotis and S. Fdida and J. F. de Rezende - “Easily-Managed Location in SONs by exploiting Space-Filling Curves”, IEEE Infocom 2005 - Student Workshop, Miami, USA, March, 2005. English, Letter size, 2 p., file: VAVFR05a.pdf

GTA-05-13 Carvalho, C. B. and de Rezende, J. F. - “Priority Control Mechanism for Performance Improvement in IEEE 802.11b Wireless LANs”, IEEE Infocom 2005 - Student Workshop, Miami, USA, March, 2005. English, Letter size, 2 p.

GTA-05-14 Amodei Jr., A., Costa, L. H. M. K., and Duarte, O. C. M. B. - “An Efficient Medium Access Control Mechanism for HomePNA”, in 12th International Conference on Telecommunications - ICT’2005, Capetown, South Africa, May 2005. English, Letter size, 5 p, file: ACD05.pdf

GTA-05-15 Viana, A. C., Amorim, M. D., Viniotis, Y., Fdida, S., and de Rezende, J. F. - “Easily-managed and Topological-Independent Location Service for Self-Organizing Networks”, in Sixth ACM International Symposium on Mobile Ad Hoc Networking and Computing - MobiHoc’2005, Urbana-Champaign, USA, May 2005. English, Letter size, 12 p., file: VAVFR05.pdf

GTA-05-16 Master Science Thesis Author - Andr?Martins Araujo Title - “ANDSR - Protocolo de Roteamento An?imo para Redes Ad Hoc" Advisor - Aloysio de Castro P. Pedroza COPPE/PEE/UFRJ - March 2005 Portuguese - A4 size - 133 p. - Directory: Andrea05

GTA-05-17 Dos Santos Filho, J.V., Courtiat J.-P., and Pedroza, A. C. P. - “Interfaces Virtuelles pour des Applications Coop?tives”,in 5?e Congres de Doctorants (EDSYS), LAAS/CNRS, May 2005. French, A4 size, 8p., file: ACD05.pdf

GTA-05-18 Taveira, D. M., Moraes, I. M., Cunha, D. O., Laufer, R. P., Bicudo, M. D. D., Campista, M. E. M., Amodei Jr. A., and Duarte, O. C. M. B. - “Comunica? Fim-a-Fim a Alta Velocidade em Redes Gigabit”, VI Workshop de Testes e Toler?ia a Falhas - WTF2005, Fortaleza, CE, May 2005. Portuguese, A4 size, 4 p., file: TMCL.pdf GTA-05-19 Master Science Thesis

Author - Aurelio Amodei Junior Title - “Um Novo Mecanismo de Acesso ao Meio para Redes Domiciliares Atrav?da Fia? Telef?ica" Advisors - Otto Carlos Muniz Bandeira Duarte and Lu? Henrique Maciel Kosmalski Costa COPPE/PEE/UFRJ - March 2005 Portuguese - A4 size - 83 p. - Directory: Aurelio05

GTA-05-20 Moraes, I. M. and Duarte, O. C. M. B. - “Um Mecanismo de Controle de Admiss?para Provis?de Servi? Diferenciados em Redes de Rajadas ?ticas”, XXIII Simp?? Brasileiro de Redes de Computadores - SBRC’2005, Fortaleza, CE, Brazil, May 2005 Portuguese, A4 size, 14 p., file: MoDu05b.pdf

GTA-05-21 Moura, G. M., de Rezende, J. F. and Jaime, G. D. - “Mecanismos de Fragmenta? Adaptativa em Redes 802.11”, XXIII Simp?? Brasileiro de Redes de Computadores - SBRC’2005, Fortaleza, CE, Brazil, May 2005 Portuguese, A4 size, 14 p., file: MRJ05.pdf

GTA-05-22 Campista, M. E. M., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Um Mecanismo Eficiente de Redu? de Colis?? para a Transmiss?de Dados Atrav?da Fia? El?ica Domiciliar”, XXIII Simp?? Brasileiro de Redes de Computadores - SBRC’2005, Fortaleza, CE, Brazil, May 2005 Portuguese, A4 size, 14 p., file: CCD05.pdf

GTA-05-23 Laufer, R. P., Velloso, P. B., and Duarte, O. C. M. B. - “Um Novo Sistema de Rastreamento de Pacotes contra Ataques de Nega? de Servi?’, XXIII Simp?? Brasileiro de Redes de Computadores - SBRC’2005, Fortaleza, CE, Brazil, May 2005 Portuguese, A4 size, 14 p., file: LVD05c.pdf

GTA-05-24 Cunha, D. O. and Duarte, O. C. M. B. - “Um Esquema Bio-inspirado para Estima? de Campo com Redes de Sensores sem Fio”, in XXIII Simp?? Brasileiro de Redes de Computadores - SBRC’2005, Fortaleza, CE, Brazil, May 2005. Portuguese, A4 size, 13 p., file CuDu05b.pdf

GTA-05-25 Bicudo, M. D. D. and Duarte, O. C. M. B. - “Um Mecanismo de Prote? em Redes WDM em Malha" , X Workshop de Ger?ia e Opera? de Redes e Servi? - WGRS2005, Fortaleza, CE, May 2005. Portuguese, A4 size, 12 p. , file: BiDu05a.pdf

GTA-05-26 Vasques, A. T.x, Abel? A. J. G. e Duarte, O. C. M. B. - “Seguran?em Comunica? Multicast”, Tutorial in XXIII Simp?? Brasileiro de Redes de Computadores SBRC’2005, Fortaleza, CE, Brazil, May 2005. Portuguese, A4 size, 49 p., file: VAD05.zip

GTA-05-27 Ziviani, A. e Duarte, O. C. M. B. - “Metrologia na Internet”, Tutorial in XXIII Simp?? Brasileiro de Redes de Computadores SBRC’2005, Fortaleza, CE, Brazil, May 2005. Portuguese, A4 size, 44 p. , file: ZiDu05.pdf GTA-05-28 Amodei Jr., A., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Um Novo

Mecanismo de Acesso ao Meio para o Padr? HomePNA”, in XXIII Simp?? Brasileiro de Redes de Computadores - SBRC’2005, Fortaleza, CE, Brazil, May 2005 Portuguese, A4 size, 14 p., file: ACD05a.pdf

GTA-05-29 Ziviani, A., Fdida, S., de Rezende, J. F., and Duarte, O. C. M. B., - “Uma Abordagem Adequada ao TCP para o Gerenciamento de Localiza? em Redes Ad Hoc M??s”, XXIII Simp?? Brasileiro de Redes de Computadores - SBRC’2005, Fortaleza, CE, Brazil, May 2005. Portuguese, A4 size, 14 p.

GTA-05-30 Master Science Thesis Author - Ingrid Teixeira Title - “Roteamento com Balanceamento de Consumo de Energia para Redes de Sensores Sem Fio" Advisor - Aloysio de Castro P. Pedroza e Jos?erreira de Rezende COPPE/PEE/UFRJ - April 2005 Portuguese - A4 size - 94 p. - Directory: Ingrid05

GTA-05-31 Master Science Thesis Author - Miguel Elias Mitre Campista Title - “Aumentando a Efici?ia do M?do de Acesso M??plo em Redes Domiciliares" Advisors - Otto Carlos Muniz Bandeira Duarte and Lu? Henrique Maciel Kosmalski Costa COPPE/PEE/UFRJ - May 2005 Portuguese - A4 size - 113 p. - Directory: Miguel05

GTA-05-32 Moraes, I. M. and Duarte, O. C. M. B. - “ Multiclass Admission Control Mechanism for Optical Burst-Switched Networks”, 19th International Teletraffic Congress - ITC19, pp. 1937-1946, Beijing, China, Aug 2005. English, Letter size, 10 p., file: MoDu05c.pdf

GTA-05-33 Delicato, F. C., Protti, F., de Rezende, J. F., and Pirmez, L. - “Uma Abordagem Baseada em QoS para Selecao de Nos Ativos em Redes de Sensores sem Fio - XXIII Simposio Brasileiro de Redes de Computadores - SBRC’2005, Fortaleza, CE, Brazil, May 2005. Portuguese, A4 size, 14 p., file:

GTA-05-34 Carvalho, C. B., and de Rezende, J. F. - “Definindo Criterios no Compartilhamento de Banda em Redes 802.11b sob Degradacao de Performance" - XXIII Simposio Brasileiro de Redes de Computadores - SBRC’2005, Fortaleza, CE, Brazil, May 2005. Portuguese, A4 size, 14 p.

GTA-05-35 Pires, A. A., de Rezende, J. F., and Cordeiro, C. - “FN-ALCA: Esquema de Controle de Potencia para Economia de Energia e Aumento de Capacidade de Redes Ad Hoc" - XXIII Simposio Brasileiro de Redes de Computadores - SBRC’2005, Fortaleza, CE, Brazil, May 2005. Portuguese, A4 size, 14 p.

GTA-05-36 Freitag, J., Fonseca, N. L. S., and de Rezende, J. F. - “Mecanismos para Aliviar o Problema de Assimetria de Retardos em Redes IEEE 802.11e" - XXIII Simposio Brasileiro de Redes de Computadores - SBRC’2005, Fortaleza, CE, Brazil, May 2005. Portuguese, A4 size, 14 p.

GTA-05-37 Master Science Thesis Author - Celso Barbosa Carvalho Title - “Mecanismo de Controle de Prioridade para Melhoria de Desempenho de Redes Infra-Estruturadas IEEE 802.11" Advisor - Jos?erreira de Rezende COPPE/PEE/UFRJ - April 2005 Portuguese - A4 size - 94 p. - Directory: Celso05

GTA-05-38 Master Science Thesis Author - Alexandre Andrade Pires Title - “Controle de Potencia em Redes Ad Hoc 802.11" Advisor - Jos?erreira de Rezende COPPE/PEE/UFRJ - June 2005 Portuguese - A4 size - 102 p. - Directory: Alexandre05

GTA-05-39 Campista, M. E. M., de Rezende, J. F., Costa, L. H. M. K., and Duarte, O. C. M. B.- “Um Mecanismo Inteligente para Controle de Taxa de Transmiss?em Redes Locais Sem Fio”, in XXI Simp?? Brasileiro de Telecomunica?s - SBrT’2005, pp. 888-893, Campinas, Brazil, September 2005. Portuguese, A4 size, 6 p.

GTA-05-40 Amodei Jr., A., Moraes, I. M., Cunha, D. O., Campista, M. E. M., Esposito, P., Costa, L. H. M. K., and Duarte, O. C. M. B.- “Uma An?se da Conectividade do Canal de Interatividade Ad Hoc para a TV Digital”, XXI Simp?? Brasileiro de Telecomunica?s - SBrT’2005, pp. 1147-1152, Campinas, Brazil, September 2005. Portuguese, A4 size, 6 p., file: AMC05a.pdf

GTA-05-41 Ziviani, A., Fdida, S., de Rezende, J. F., and Duarte, O. C. M. B., - “A TCP-Tailored Approach to the Location Management in Mobile Ad Hoc Networks”, IEEE Communication Letters, to appear. English, A4 size, 3 p.

GTA-05-42 Laufer, R. P., Velloso, P. B., Cunha, D. O., and Duarte, O. C. M. B.- “Um Procedimento Alternativo de Reconstru? de Rota para o Rastreamento de Pacotes IP”, XXI Simp?? Brasileiro de Telecomunica?s - SBrT’2005, Campinas, Brazil, September 2005. Portuguese, A4 size, 6 p.

GTA-05-43 Laufer, R. P., Velloso, P. B., and Duarte, O. C. M. B.- “Generalized Bloom Filters”, Technical Report, Electrical Engineering Program, COPPE/UFRJ, 2005. English, Letter size, 9 p.

GTA-05-44 Costa, L. H. M. K., Fdida, S., and Duarte, O. C. M. B. - “Incremental Service Deployment Using the Hop By Hop Multicast Routing Protocol”, in IEEE/ACM Transactions on Networking, ISSN 1063-6692, vol. 14, no. 3, pp. 543-556, June 2006. English, Letter size, 14 p.

GTA-05-45 Campista, M. E. M., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Improving the Data Transmission Throughput over the Home Electrical Wiring”, To appear in 30th Annual IEEE Conference on Local Computer Networks - LCN’2005, Sydney, Australia, November 2005. English, Letter size, 8 p.

GTA-05-46 de Rezende, J. F., Costa, L. H. M. K., and Rubinstein, M. G. - “Avalia? Experimental e Simula? do Protocolo TCP em Redes de Alta Velocidade”, To appear in XXI Simp?? Brasileiro de Telecomunica?s - SBrT’2005, Campinas, Brazil, September 2005. Portuguese, A4 size, 6 p.

GTA-05-47 Cunha, D. O., Laufer, R. P., Moraes, I. M., Bicudo, M. D. D., Velloso, P. B., and Duarte, O. C. M. B. - “A Bio-Inspired Field Estimation Scheme for Wireless Sensor Networks”, Annals of Telecommunications - Special Issue in Wireless Sensor Networks, vol. 60, n. 7/8, pp. 806-818, ISSN 0003-4347, July/August 2005. English, A4 size, 13 p.

GTA-05-48 Laufer, R. P., Moraes, I. M., Velloso, P. B., Bicudo, M. D. D., Campista, M. E. M., Cunha, D. O., Costa, L. H. M. K., Duarte, O. C. M. B. - “Nega? de Servi? Ataques e Contramedidas”, in Minicursos do Simp?? Brasileiro de Seguran?da Informa? e de Sistemas Computacionais - SBSeg’2005 , Florian??is, Brazil, pp. 1-63, September 2005. Portuguese, A4 size, 63 p., file: LMVB05a.pdf

GTA-05-49 Cunha, D. O., Costa, L. H. M. K., and Duarte, O. C. M. B. - “On the Limits of Power Saving Techniques for Ad Hoc Networks”, Technical Report, Electrical Engineering Program, COPPE/UFRJ, 2005. English, Letter size, 24 p., file: CuCoDu05a.pdf

GTA-05-50 Moraes, I. M. and Duarte, O. C. M. B. - “An Efficient Multiclass Mechanism for Optical Burst-Switching Networks”, 5th IEEE/CreateNet International Workshop on Optical Burst/Packet Switching - WOBS’05, Boston, USA, October 2005. English, Letter size, 10 p., file: MoDu05d.pdf

GTA-05-51 Master Science Thesis Author - Rafael Pinaud Laufer Title - “Rastreamento de Pacotes IP contra Ataques de Nega? de Servi? Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - September 2005 Portuguese - A4 size - 93 p. - Directory: Laufer05

GTA-05-52 Pires, A. A. and de Rezende, J. F. - “Protecting Transmissions when using Power Control on 802.11 Ad Hoc Networks”, in Fourth Annual Mediterranean Workshop on Ad Hoc Networks - MedHocNet’2005, ?e de Porquerolles, France, June 2005. English, A4 size, 10 p., file: PiRe05.pdf

GTA-05-53 Augusto, C. H. P., Cerveira, C. R., and de Rezende, J. F. - "Controle de Admiss?Adaptativo para Redes Ad Hoc 802.11”, in XXI Simp?? Brasileiro de Telecomunica?s - SBrT’2005, Campinas, Brazil, September 2005. Portuguese, A4 size, 6 p., file: ACR05.pdf

GTA-05-54 Ramella, I. and de Rezende, J. F. - “Minimizando a Interfer?ia em Redes Ad Hoc com Antenas Direcionais”, in XXI Simp?? Brasileiro de Telecomunica?s - SBrT’2005, Campinas, Brazil, September 2005. Portuguese, A4 size, 6 p., file: RaRe05.pdf

GTA-05-55 Doctor of Science Thesis Author - Fl?a Coimbra Delicato Title - “Middleware Baseado em Servi? para Redes de Sensores Sem Fio" Advisors - Jos?erreira de Rezende and Luci Pirmez COPPE/PEE/UFRJ - June 2005 Portuguese - A4 size - 180 p.

GTA-05-56 Doctor of Science Thesis Author - Aline Carneiro Viana Title - “Localisation et Routage dans les R?aux Auto Organisables ?arge-?helle: des Tables de Hachage Distribu? aux Structures d’Adressage Adaptatives" Advisors - Serge Fdida and Jos?erreira de Rezende Univ. Paris VI - July 2005 English - A4 size - 100 p.

GTA-05-57 Campista, M. E. M., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Improving the Multiple Access Method of CSMA/CA Home Networks”, IEEE Consumer Communication and Networking Conference - CCNC’2006, Las Vegas, Nevada, USA, January 2006. English, A4 size, 5 p., file: CCD06.pdf

GTA-05-58 Delicato, F. C., Pires, P. F, de Rezende, J. F., and Pirmez, L. - “Uma Framework baseada em Tecnologias Web para a Constru? de Redes de Sensores sem Fio Flex?eis" - to appear in XI Simp?? Brasileiro de Sistemas Multim?ia e Web - WebMedia’2005, Po? de Caldas, MG, Brazil, December 2005. Portuguese, A4 size, 10 p., file:

GTA-05-59 Rubinstein, M. G., Costa, L. H. M. K., Campista, M. E. M., Cunha, D. O., Amodei Jr., A., Velloso, P. B., and Duarte, O. C. M. B. - “Analysis of Medium Access Control Protocols for Home Networks”, Journal of Communication and Informatio Systems, ISSN 1980-6604, v. 22, p. 10-23, 2007. English, A4 size, 35 p., file: RCCC07.pdf

GTA-05-60 Master Science Thesis Author - Marco Dias Dutra Bicudo Title - “Sobreviv?ia em Redes ?ticas Transparentes" Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - December 2005 Portuguese - A4 size - 75 p. - Directory: Bicudo05

Technical Reports — 2004

GTA-04-01 Viana, A. C., Amorim, M. D., Fdida, S., and de Rezende, J. F. - “An Underlay Strategy for Indirect Routing”, to appear in ACM Wireless Networks and Applications, vol. 10, no. 6, November 2004, ISSN 1022-0038. English, Letter size, 8 p.

GTA-04-02 Ziviani, A., Fdida, S., de Rezende, J. F., and Duarte, O. C. M. B. - “Toward a Measurement-based Geographic Location Service”, Lectures Notes in Computer Science 3015, Chadi Barakat and Ian Pratt (Eds.), pp. 43-52, ISSN 0302-9743, ISBN 3-540-21492-5, Springer-Verlag, 5th Passive & Active Measurement Workshop - PAM’2004, Antibes Juan-les-Pins, France, April 2004. English, A4 size, 10 p., file: ZFRD04.ps.gz GTA-04-03 Delicato, F. C., Pires, P. F., Lages, A., de Rezende, J. F., and Pirmez, L. - “Mid-

dleware Orientado a Servi? para Redes de Sensores sem Fio”, in XXII Simp?? Brasileiro de Redes de Computadores - SBRC’2004, pp. 15-28, Gramado, RS, May 2004. Portuguese, A4 size, 14 p., file: DPLRP04.pdf

GTA-04-04 Pires, A. A., Fontes, M. F., and de Rezende, J. F. - “Proposta e Avalia? de um Esquema de Controle de Pot?ia com Mem?? em Redes Ad Hoc 802.11”, in XXII Simp?? Brasileiro de Redes de Computadores - SBRC’2004, pp. 351-364, Gramado, RS, May 2004. Portuguese, A4 size, 14 p., file: PFR04.pdf

GTA-04-05 Pinto, A. J. G., Stochero, J., and de Rezende, J. F. - “Descoberta de Rotas e Mecanismo de Agrega? em Redes de Sensores”, in XXII Simp?? Brasileiro de Redes de Computadores - SBRC’2004, pp. 567-570, Gramado, RS, May 2004. Portuguese, A4 size, 4 p., file: PSR04.pdf

GTA-04-06 Filho, J. C. M., de Rezende, J. F., and Pirmez, L. - “Dynamic EDCF Parameters for QoS Control in 802.11 Hot Spots”, in Journal of the Brazilian Computer Society, vol. 9, no. 2, pp. 5-15, ISSN 0104-6500, Brazilian Computer Society, November 2003. English, A4 size, 11 p.

GTA-04-07 Undergraduate Final Project Author - Igor Monteiro Moraes Title - “Um Prot??o de Canal de TV Usando um Servi? de Difus?IP com Fonte Espec?ica” Advisor - Otto Carlos Muniz Bandeira Duarte DEL/POLI/UFRJ - March 2004 Portuguese - A4 size - 109 p. - Directory: Igor04

GTA-04-08 Undergraduate Final Project Author - Rafael Pinaud Laufer Title - “Projeto e Implementa? de uma Zona Desmilitarizada” Advisor - Otto Carlos Muniz Bandeira Duarte DEL/POLI/UFRJ - March 2004 Portuguese - A4 size - 103 p. - Directory: Laufer04

GTA-04-09 Undergraduate Final Project Author - Marco Dias Dutra Bicudo Title - “Implementa? e An?se de Desempenho de um Ambiente de Testes de Qualidade de Servi?’ Advisor - Otto Carlos Muniz Bandeira Duarte DEL/POLI/UFRJ - March 2004 Portuguese - A4 size - 68 p. - Directory: Bicudo04

GTA-04-10 Gueye, B., Ziviani, A., Fdida, S., de Rezende, J. F., and Duarte, O. C. M. B. - “Two-Tier Geographic Location of Internet Hosts”, in Lecture Notes in Computer Science 3079, pp. 730-739, ISSN 0302-9743, ISBN 5-540-22262-6, 7th IEEE International Conference on High Speed Networks and Multimedia Communications - HSNMC’04, Toulouse, France, June 2004. English, A4 size, 10 p., file: GZFRD04.pdf

GTA-04-11 Villela, B. A. M. and Duarte, O. C. M. B. - “Maximum Throughput Analysis in Ad Hoc Networks”, Lecture Notes in Computer Science - Networking’2004, pp. 223-234, Vol. 3042, ISSN 0302-9743, ISBN 3-540-21959-5, Springer-Verlag, May 2004.

GTA-04-12 Fernandez, M. P., Pedroza, A. C. P., and de Rezende, J. F. de. - “Dynamic QoS Provisioning in DiffServ Domains using Fuzzy Logic Controllers”, in Kluwer Telecommunication Systems, ISSN 1018-4864, Vol. 26, No. 1, pp. 9-32, May 2004. English, Letter size, 24 p.

GTA-04-13 Master Science Thesis Author - Daniel de Oliveira Cunha Title - “Conserva? de Energia em Redes Ad Hoc" Advisors - Otto Carlos Muniz Bandeira Duarte and Lu? Henrique Maciel Kosmalski Costa COPPE/PEE/UFRJ - April 2004 Portuguese - A4 size - 98 p. - Directory: Cunha04

GTA-04-14 Master Science Thesis Author - Fabr?io Jorge Lopes Ribeiro Title - “An?se de Associa?s Seguras em Sistemas M??s de Terceira Gera?" Advisor - Aloysio de Castro P. Pedroza COPPE/PEE/UFRJ - March 2004 Portuguese - A4 size - 114 p. - Directory: Fabricio04

GTA-04-15 Stochero, J. M., Pinto, A. J. G., and de Rezende, J. F. - “Proposta e Avalia? de um Mecanismo de Agrega? de Dados em Redes de Sensores”, to appear in III Workshop em Desempenho de Sistemas Computacionais e de Comunica? - WPerformance’2004 (XXIV Congresso da Sociedade Brasileira de Computa? - SBC 2004), Salvador, BA, August 2004. Portuguese, A4 size, 20 p, file: SPR04.pdf.

GTA-04-16 Priess, W., de Rezende, J. F., and Pirmez, L. - “Adaptive Interpiconet Scheduling for Multipurpose Scatternet Scenarios”, to appear in Elsevier Ad Hoc Networks Journal, ISSN 1570-8705, December 2004. English, Letter size, 20 p.

GTA-04-17 Velloso, P. B., Cunha, D. O., Amodei Jr., A., Rubinstein, M. G., and Duarte, O. C. M. B. - “Redes Domiciliares: Princ?ios e Desafios das Tecnologias sem Novos Fios”, in Minicursos Simp?? Brasileiro de Redes de Computadores, pp. 221, 269, May 2004. Portuguese, A4 size, 48 p., file: VCAR04.pdf Presentation, file: redesdom.pdf

GTA-04-18 Araujo, A.M. and Pedroza, A.C.P. - “Especifica? Formal de um Protocolo de Roteamento An?imo para Redes Ad-Hoc”, IV Workshop em Seguran?de Sistemas Computacionas, WSEG2004, p.151-162, Gramado, RS, Brazil, May 2004 Portuguese, A4 size, 12p., file: AP04b.ps.gz

GTA-04-19 Araujo, A.M. and Pedroza, A.C.P. - “Um Protocolo de Roteamento An?imo para Redes M??s Ad Hoc”, I Workshop de Computa? da Regi?Sul, I WORKCOMP-SUL, Florian??is,SC, Brazil, May 2004 Portuguese, A4 size, 10p., file: AP04.ps.gz

GTA-04-20 Pereira, I.C.M. and Pedroza, A.C.P. - “An?se de Redes M??s Ad Hoc para Cen?os de Opera?s Militares”, I Workshop de Computa? da Regi?Sul, I WORKCOMP-SUL, Florian??is,SC, Brazil, May 2004 Portuguese, A4 size, 10p., file: PP04.pdf.gz

GTA-04-21 Master Science Thesis Author - Antonio Jos?on?ves Pinto Title - “Mecanismo de Agrega? de Dados Empregando T?icas Param?icas em Redes de Sensores" Advisor - Jos?erreira de Rezende COPPE/PEE/UFRJ - June 2004 Portuguese - A4 size - 115 p. - Directory: Antonio04

GTA-04-22 Villela, B. A. M. and Duarte, O. C. M. B. - “Calculating the Maximum Throughput in Multihop Ad Hoc Networks”, Lecture Notes in Computer Science, vol. 3124, pp. 443-449, ISSN: 0302-9743, ISBN: 3-540-22571-4, Springer-Verlag, Proc. of the International Conference on Telecommunications (ICT’04), Fortaleza, Brazil, August 2004. English, A4 size, 7 p.

GTA-04-23 Master Science Thesis Author - Bernardo Antunes Maciel Villela Title - “An?se da Vaz?M?ma em Redes Ad Hoc" Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - June 2004 Portuguese - A4 size - 79 p. - Directory: Bernardo04

GTA-04-24 Master Science Thesis Author - Jorgito Matiuzzi Stochero Title - “Roteamento Favorecendo Agrega? de Dados em Redes de Sensores" Advisor - Jos?erreira de Rezende COPPE/PEE/UFRJ - June 2004 Portuguese - A4 size - 95 p. - Directory: Jorgito04

GTA-04-25 Pinto, A. J. G., Stochero, J., and de Rezende, J. F. - "Aggregation-aware Routing on Wireless Sensor Networks”, to appear in the Lecture Notes in Computer Science - Proc. of 9th Intl. Conference on Personal Wireless Comunications (PWC’04), Delft, The Netherlands, September 2004. English, A4 size, 10 p.

GTA-04-26 Master Science Thesis Author - Jaime Cesar Ribeiro Lopes Title - “An?se de uma Arquitetura de Seguran?em Sistemas M??s de Terceira Gera?" Advisor - Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - March 2004 Portuguese - A4 size - 120 p. - Directory: Jaime04

GTA-04-27 Doctor of Science Thesis Author - Ricardo Neimo Bel?Lima Title - “Uma Metodologia para Implementa? de Protocolos de Comunica? utilizando Hardware/Software Codesign" Advisor - Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - March 2004 Portuguese - A4 size - 141 p. - Directory: Belem04

GTA-04-28 Campista, M. E. M., Velloso, P. B., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Uma An?se da Capacidade de Transmiss?na Rede de Energia El?ica Domiciliar”, in XXI Simp?? Brasileiro de Telecomunica?s - SBT’2004, Bel? Brazil, September 2004. Portuguese, A4 size, 6 p., file: CVCD04.pdf

GTA-04-29 Amodei Jr., A., Cunha, D. O., Costa, L. H. M. K., Rubinstein, M. G., and Duarte, O. C. M. B. - “Uma An?se de Desempenho do Padr?HomePNA”, in XXI Simp?? Brasileiro de Telecomunica?s - SBT’2004, Bel? Brazil, September 2004. Portuguese, A4 size, 6 p, file: ACC04.pdf.

GTA-04-30 Rocha, R. R., Costa, L. H. M. K., and de Rezende, J. F. - “Uma An?se de Desempenho de Redes Ad Hoc de M??plos Saltos Utilizando Antenas Direcionais”, to appear in XXI Simp?? Brasileiro de Telecomunica?s - SBT’2004, Bel? Brazil, September 2004. Portuguese, A4 size, 6 p, file: RCR04.pdf.

GTA-04-31 Carvalho, C.B. and de Rezende, J. F. - “Sele? Din?ca de Par?tros de QoS na Resolu? da Anomalia de Performance do IEEE 802.11b”, to appear in XXI Simp?? Brasileiro de Telecomunica?s - SBT’2004 , Bel? Brazil, September 2004. Portuguese, A4 size, 6 p.

GTA-04-32 Teixeira, I., de Rezende, J. F., and Pedroza, A.C.P. - "Wireless Sensor Network: Improving the Network Energy Consumption”, XXI Simp?? Brasileiro de Telecomunica?s - SBT’2004, Bel? Brazil, September 2004. English, A4 size, 6 p, file: TRP.pdf.

GTA-04-33 Freitag, J., Fonseca, N.L.S., and de Rezende, J. F. - "Mecanismos de Controle de Admiss?Baseado em Medi?s em Redes IEEE 802.11e”, to appear in XXI Simp?? Brasileiro de Telecomunica?s - SBT’2004, Bel? Brazil, September 2004. Portuguese, A4 size, 6 p.

GTA-04-34 Araujo, A.M. and Pedroza, A.C.P. - “Um Protocolo de Roteamento An?imo para Redes Ad Hoc”, to apper in XV Congresso Brasileiro de Autom?ca - CBA’2004, Gramado, RS, Brazil, September 2004. Portuguese, A4 size, 6p.

GTA-04-35 Master Science Thesis Author - Ivana Cardial de Miranda Pereirar Title - “An?se de Roteamento em Redes M??s Ad Hoc em Cen?os de Opera?s Militares" Advisor - Aloysio de Castro P. Pedroza COPPE/PEE/UFRJ - June 2004 Portuguese - A4 size - 92 p. - Directory: Ivana04

GTA-04-36 Pereira, I.C.M and Pedroza, A.C.P. - “Redes M??s Ad Hoc Aplicadas a Cen?os Militares”, to appear in IV Congresso Brasileiro de Computa? - CBComp’2004, Itaja? SC, Brazil, October 2004. Portuguese, A4 size, 6 p.

GTA-04-37 Moraes, I. M., and Duarte, O. C. M. B. - “Uma Compara? entre Protocolos de Roteamento /it Multicast para Distribui? de TV na Internet”, XXI Simp?? Brasileiro de Telecomunica?s - SBT’2004, Bel? Brazil, September 2004. Portuguese, A4 size, 6 p.

GTA-04-38 Amodei Jr., A., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Increasing the Throughput of the HomePNA MAC Protocol”, 29th Annual IEEE Conference on Local Computer Networks - LCN2004, Tampa, USA, November 2004. English, A4 size, pp. 293-301, file: ACD04.pdf.

GTA-04-39 Ara?? S. G, Mesquita A. C., and Pedroza, A.C.P. - “A Scenario Based Approach to Protocol Design Using Evolutionary Techniques”, in Proceedings of the 1st Euro-pean Workshop on Evolutionary Computation in Communications, Networks and Connected Systems (EVOCOMNET’04), LNCS, Vol. 3005, pp. 178-187, Coimbra, Portugal, April 2004.

GTA-04-40 Ara?? S. G, Mesquita A. C., and Pedroza, A.C.P. - “Improvements in FSM Evolu-tions from Partial Input/Output Sequences”, in Proceedings of the 1st International Workshop on Evolvable Hardware, (E-HARD’2004), LNCS, pp. 1311-1319, Kra-kow, Poland, June 2004.

GTA-04-41 Doctor of Science Thesis Author - S?io Granato de Ara?? Title - “S?tese de Sistemas Digitais Utilizando T?icas Evolutivas" Advisor - Aloysio de Castro P. Pedroza and Ant?io C. de Mesquita F. COPPE/PEE/UFRJ - July 2004 Portuguese - A4 size - Directory: Granato04

GTA-04-42 Master Science Thesis Author - Fagundes Pereira da Silva Title - “Um Benchmark para Sistemas de Gerenciamento de Workflow" Advisor - Jorge Lopes de Souza Leão COPPE/PEE/UFRJ - March 2004 Portuguese - A4 size - 91 p. - Directory: Fagundes04

GTA-04-43 Doctor of Science Thesis Author - Paulo André da Silva Gonçalves Title - “Communications de Groupes : Analyse et Performances des Mécanismes d’agrégation de l’information de Contrôle (Group Communications: Analysis and Performance of Feedback Aggregation Mechanisms)” Supervisors - Guy Pujolle, Otto Carlos Muniz Bandeira Duarte, and José Ferreira de Rezende Université Pierre et Marie Curie (Paris 6), France - June 2004 English - A4 size - 123 p. - Directory: Goncalves04

GTA-04-44 Gon?ves, P. A. da S. - “Les Réseaux de Capteurs”, L’Internet Ambiant, chapter 8, pp. 193-214, Hermes Science Publications, ISBN 2-7462-0988-8, October 2004. // French, A4 size, 23 p.

GTA-04-45 Cunha, D. O., Costa, L. H. M. K., and Duarte, O. C. M. B. - "Uma An?se das T?icas de Conserva? de Energia em Redes Ad Hoc”, to appear in XXI Simp?? Brasileiro de Telecomunica?s - SBT’2004, Bel? Brazil, September 2004.

Portuguese, A4 size, 6 p.

GTA-04-46 Cunha, D. O., Costa, L. H. M. K., and Duarte, O. C. M. B. - "Analyzing the Energy Consumption of IEEE 802.11 Ad Hoc Networks”, to appear in Sixth IFIP IEEE International Conference on Mobile and Wireless Communication Networks - MWCN’2004, Paris, France, October 2004. English, Letter size, 12 p.

GTA-04-47 Viana, A. C., Amorim, M. D., Fdida, S., and de Rezende, J. F. - “Self-organization in spontaneous networks: the approach of DHT-based routing protocols”, to appear in Ad Hoc Networks Journal, Elsevier Academic Publishers, ISSN 1570-8705, 2004. English, Letter size, 20 p.

GTA-04-48 Freitag, J., Fonseca, N.L.S., and de Rezende, J. F. - "Admission Control in IEEE 802.11 Networks”, to appear in 10th IEEE Workshop on Computer-Aided Modeling, Analysis and Design of Communication Links and Networks (CAMAD’2004), Dallas, USA, November 2004 English, Letter size, 8p.

GTA-04-49 Master Science Thesis Author - Luiz Gustavo Silva Rocha Title - “Uma An?se dos Impactos de A?s Maliciosas de N??o Roteamento em Redes Ad Hoc" Advisors - Otto Carlos Muniz Bandeira Duarte and Lu? Henrique Maciel Kosmalski Costa COPPE/PEE/UFRJ - July 2004 Portuguese - A4 size - 61 p. - file:Rocha04/Rocha04.pdf

GTA-04-50 Ziviani, A., Fdida, S., de Rezende, J. F., and Duarte, O. C. M. B. - “Improving the Accuracy of Measurement-Based Geographic Location of Internet Hosts”, Computer Networks Journal, Elsevier Science, ISSN 1389-1286, to appear. English, A4 size, 33 p., file: ZFRD04a.ps.gz

GTA-04-51 Pires, A. A. and de Rezende, J. F. - “Controle de Pot?ia em Redes Ad Hoc: Evitando Colis??em Enlaces Assim?icos”, to appear in VI Workshop de Comunica? sem Fio e Computa? M?? - WCSF’2004, Fortaleza, CE, RS, October 2004. Portuguese, A4 size, 14 p.

GTA-04-52 Vida, R., Costa, L. H. M. K., and Fdida, S. - “Mobile Hop-By-Hop Multicast Routing”, Computer Networks Journal, vol. 44, no. 6, April 2004, ISSN 1389-1286. English, A4 size, pp. 789-81.

GTA-04-53 Costa, L. H. M. K., Amorim, M. D., and Fdida, S. - "Reducing Latency and Overhead of Route Repair with Controlled Flooding”, ACM/Kluwer Wireless Networks, vol. 10, no. 4, pp. 347-358, July 2004, ISSN 1022-0038. English, A4 size, 12 p.

GTA-04-54 Cunha, D. O., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Uma An?se do Consumo de Energia em Redes Ad Hoc (vers?estendida)”, Telecomunica?s, vol. 07, no. 01, pp. 39-47, July 2004, ISSN 1516-2338. Portuguese, A4 size, 9 p.,

GTA-04-55 Undergraduate Final Project Author - Rog?o de Castro Melo Title - “Ferramenta de Aux?io ?Instala? de Redes 802.11 Infra-estruturadas" Advisor - Jos?erreira de Rezende DEL/POLI/UFRJ - October 2004 Portuguese - A4 size - 109 p. - Directory: Rogerio04

GTA-04-56 Velloso, P. B., Cunha, D. O., Amodei Jr., A., Rubinstein, M. G., and Duarte, O. C. M. B. - “Redes Domiciliares: Aplica?s, Tecnologias, Desafios e Tend?ias”, in XXI Simp?? Brasileiro de Telecomunica?s - SBT’2004, Bel? Brazil, September 2004. Portuguese, A4 size, 93 p., file: VCAR04a.pdf

GTA-04-57 Velloso, P. B., Campista, M. E. M., Cunha, D. de O., Costa, L. H. M. K., and Duarte, O. C. M. B. - "Analyzing the Performance of Wireless Local Area Networks with an Improved Collision Avoidance Mechanism”, Revista da Sociedade Brasileira de Telecomunica?s, vol. 19, no. 3, pp. 53-60, December 2004. English, A4 size, 8 p., file: VCC04.pdf

Technical Reports — 2003

GTA-03-01 Anton, E. R. and Duarte, O. C. M. B. - “Uma An?se de Protocolos para o Estabelecimento de Chave de Grupo em Redes Ad Hoc”, March 2003. Portuguese, A4 size, 8 p., file: AnDu03.ps.gz

GTA-03-02 Velloso, P. B., Rubinstein, M. G., and Duarte, O. C. M. B. - “Analyzing Voice Transmission Capacity on Ad Hoc Networks”, International Conference on Communications Technology - ICCT 2003, vol.2, pp. 1254-1257, Beijing, China, April 2003. English, A4 size, 4 p., file: VRD03.ps.gz

GTA-03-03 Ziviani, A., Fdida, S., de Rezende, J. F., and Duarte, O. C. M. B., “Une Approche TCP pour la Gestion de la Localisation dans les R?aux Ad Hoc Mobiles”, Colloque Francophone sur la Gestion de R?aux et de Services - GRES’2003, pp. 443-454, Fortaleza, Brazil, February 2003. French, A4 size, 12p., file: ZFRD03.ps.gz

GTA-03-04 Viana, A. C., Amorim, M. D., Fdida, S., and de Rezende, J. F., “Indirect Routing using Distributed Location Information”, IEEE International Conference on Pervasive Computing and Communications - PERCOM’2003, Dallas-Fort Worth, Texas, March 2003. English, Letter size, 8 p., file: VAFR03.ps.gz

GTA-03-05 Costa, L. H. M. K., Amorim, M. D., and Fdida, S., “Utilisant l’inondation contr??pour la d?uverte de routes dans les r?aux ad hoc”, Colloque Francophone sur la Gestion de R?aux et de Services - GRES’2003, pp. 387-398, Fortaleza, Brazil, February 2003. French, A4 size, 12 p., file: CoAmFd03.ps.gz GTA-03-06 Anton, E. R. and Duarte, O. C. M. B. - “Performance Analysis of Group Key

Establishment Protocols in Ad Hoc Networks”, March 2003. English, A4 size, 6 p., file: AnDu03a.ps.gz

GTA-03-07 Velloso, P. B., Rubinstein, M. G., and Duarte, O. C. M. B. - “Uma An?se da Capacidade da Transmiss?de Voz em Redes Ad Hoc”, XXI Simp?? Brasileiro de Redes de Computadores - SBRC’2003, Natal, RN, Brazil, May 2003. Portuguese, A4 size, 12 p., file: VRD03a.ps.gz

GTA-03-08 Cunha, D. O., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Um Mecanismo de Roteamento para o Consumo Balanceado de Energia em Redes M??s Ad Hoc”, XXI Simp?? Brasileiro de Redes de Computadores - SBRC’2003, Natal, RN, Brazil, May 2003. Portuguese, A4 size, 14 p., file: CCD03.ps.gz

GTA-03-09 Gon?ves, P. A. da S., de Rezende, J. F., Duarte, O. C. M. B., and Pujolle G. - “Improving Feedback Merging for Source-Adaptive Layered Multicast Schemes”, To appear in Kluwer Cluster Computer Journal, ISSN 1386-7857, 2003. English, A4 size, 24 p., file: GRDP03a.ps.gz

GTA-03-10 Lima, M. M. A. E., Fonseca, N. L. S., and de Rezende, J. F. - “On the Performance of TCP Loss Recovery Mechanisms”, International Conference on Communications - ICC’2003 2003, Anchorage, AL, USA, May 2003. English, A4 size, 5 p., file: LFR03.pdf

GTA-03-11 Priess, W., de Rezende, J. F., Carmo, L. F. R. C., and Pirmez, L. - “Um Mecanismo de Escalonamento Parametriz?l para Scatternets Bluetooth”, XXI Simp?? Brasileiro de Redes de Computadores - SBRC’2003, Natal, RN, Brazil, May 2003. Portuguese, A4 size, 14 p. file: PRCP03.ps.gz

GTA-03-12 Filho, J. C. M., de Rezende, J. F., and Pirmez, L. - “Sele? Din?ca de Par?tros de Controle de Qualidade de Servi?em Redes IEEE 802.11 Infra-Estruturadas”, XXI Simp?? Brasileiro de Redes de Computadores - SBRC’2003, Natal, RN, Brazil, May 2003. Portuguese, A4 size, 14 p. file: FRP03.ps.gz

GTA-03-13 Fernandez, M. P., Pedroza, A. C. P., and de Rezende, J. F. de. - “Implementa? de Pol?icas de Gerenciamento atrav?de l??a Fuzzy visando melhoria da Qualidade de Servi?(QoS)”, XXI Simp?? Brasileiro de Redes de Computadores - SBRC’2003, Natal, RN, Brazil, May 2003. Portuguese, A4 size, 16 p.

GTA-03-14 Master Science Thesis Author - Werner Priess Title - ‘Mecanismos de Escalonamento com Qualidade de Servi?em Redes com Tecnologia Bluetooth’ Advisor - Jos?erreira de Rezende and Luci Pirmez NCE/UFRJ - January 2003 Portuguese - A4 size - 111 p. - File: Werner03.pdf

GTA-03-15 Master Science Thesis Author -Jos?oelho de Melo Filho Title - ‘Mecanismos de Controle de Qualidade de Servi? em Redes 802.11’ Advisor - Jos?erreira de Rezende and Luci Pirmez NCE/UFRJ - January 2003 Portuguese - A4 size - 120 p. - File: Coelho03.pdf GTA-03-16 Master Science Thesis

Author - Eric Ricardo Anton Title - “An?se de Desempenho de Protocolos para Estabelecimento de Chave de Grupo em Redes Ad Hoc” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - March 2003 Portuguese - A4 size - 53 p. - File: Eric03.ps.gz

GTA-03-17 Viana, A. C., Amorim, M. D., Fdida, S., and de Rezende, J. F., “Routage pair- ?air dans les r?aux spontan??arge ?elle”, 5? Rencontres Francophones sur les Aspects Algorithmiques des T?communications - ALGOTEL’2003, pp. 55-60, France, May 2003. French, Letter size, 7 p., file: VAFR03a.ps.gz

GTA-03-18 Vidal, P. C. S. and Duarte, O. C. M. B. - “Reducing the Establishment Time in IP Tunnels by Using Staged Refresh Timers”, 2nd IEEE International Symposium on Network Computing and Applications - NCA 2003, pp. 389-396, Cambridge, MA, USA, April 2003. English, A4 size, 8 p., file: ViDu03.ps.gz

GTA-03-19 Rocha, L. G. S., Costa, L. H. M. K., and Duarte, O. C. M. B. - "Avali? do Impacto da A? Maliciosa de N??o Roteamento em Redes Ad Hoc”, Workshop em Seguran?de Sistemas Computacionais - WSeg 2003, pp. 69-76, Natal, RN, Brazil, May 2003. Portuguese, A4 size, 8 p., file: RCD03.ps

GTA-03-20 Moraes, I. M., Bicudo, M. D. D., Cardoso, K. V., Vasconcelos, S. V., de Rezende, J. F., and Duarte, O. C. M. B. - "Desenvolvimento de um Ambiente de Testes com Suporte ?Qualidade de Servi?para Transmiss?de V?eo Digital”, IV Workshop RNP2 - WRNP2’2003, Natal, RN, Brazil, May 2003. Portuguese, A4 size, 4 p., file: MBCVRD03.pdf

GTA-03-21 Velloso, P. B., Rubinstein, M. G., and Duarte, O. C. M. B. - “Uma Avalia? da Capacidade da Transmiss?de Voz em Redes 802.11 em Modo Ad Hoc”, Revista da Sociedade Brasileira de Telecomunica?s, vol. 18, no. 2, pp. 186-197, ISSN 0102-986X, October 2003. Portuguese, A4 size, 12 p., file: VRD03b.ps.gz

GTA-03-22 Costa, L. H. M. K. and Duarte, O. C. M. B. - "Roteamento Mulicast na Internet”, Mini-curso do XXI Simp?? Brasileiro de Redes de Computadores - SBRC’2003, Natal, RN, Brazil, May 2003. Portuguese, A4 size, 47 p., file: mc-sbrc03.pdf

GTA-03-23 Ziviani, A., Fdida, S., de Rezende, J. F., and Duarte, O. C. M. B., “Similarity Models for Internet Host Location”, IEEE International Conference on Networks - ICON’2003, pp. 81–86, Sydney, Australia, September 2003. English, A4 size, 6 p., file: ZFRD03b.ps.gz

GTA-03-24 Ziviani, A., Fdida, S., de Rezende, J. F., and Duarte, O. C. M. B., “Demographic Placement for Internet Host Location”, IEEE GLOBECOM’2003, vol. 7, pp. 3813-3817, San Francisco, CA, USA, December 2003. English, Letter size, 5 p., file: ZFRD03d.ps.gz GTA-03-25 Villela, B. A. M. and Duarte, O. C. M. B., “Uma An?se de Protocolos de Rotea-

mento Sob Demanda de Redes Ad Hoc”, SBT’2003, Rio de Janeiro, RJ, Brasil, Outubro 2003. Portuguese, A4 size, 6 p., file: VD03.ps.gz

GTA-03-26 Priess, W., de Rezende, J. F., Carmo, L. F. R. C., and Pirmez, L. - “Enhancing Performance in Scatternets via Scheduling Algorithm Parametrization, Lectures Noets in Computer Science 2775, Marco Conti, Silvia Giordano, Enrico Gregori Stephan Olariu (Eds.), pp. 741-775, ISSN 0302-9743, ISBN 3-540-20123, Springer-Verlag, September 2003, IFIP-TC6 8th International Conference on in Personal Wireless Communications - PWC’2003, Venice, Italy, September 2003. English, A4 size, 12 p. File: PRP03.pdf

GTA-03-27 Fernandez, M.P., Pedroza, A.C.P., and de Rezende, J. F. - “Converting QoS Policy Specification into Fuzzy Logic Parameters”, in the 18th International Teletraffic Congress, Berlin, Germany, September 2003. English, A4 size, 12 p. File: FPR03a.pdf

GTA-03-28 Miranda, M. N., Lima, R. N. B., Pedroza, A. C. P., and Mesquita F, A. C. M.- “Projeto de Protocolos Utilizando HW/SW Codesign Baseado na Otimiza? de Desempenho por Algoritmos Gen?cos”, SBT’2003, Rio de Janeiro, RJ, Brazil, October 2003. Portuguese, A4 size, 20 p. - file: MRAA03.ps.gz

GTA-03-29 Undergraduate Final Project Author - Andrea Martins Araujo Title - “Modelagem, Verificacao e Prototipagem do protocolo de comunicacao L2CAP” Advisor - Aloysio de Castro P. Pedroza DEL/EE/UFRJ - January 2003 Portuguese - A4 size - 89 p. - file: Andrea03.ps.gz

GTA-03-30 Master Science Thesis Author - Renato Bagatelli Title - “Projeto Formal de Protocolos de Acesso ao Meio e de Roteamento para Redes Ad Hoc” Advisor - Aloysio de Castro P. Pedroza COPPE/PEE/UFRJ - April 2003 Portuguese - A4 size - 149 p. - File: Bagatelli03.ps.gz

GTA-03-31 Ziviani, A., Fdida, S., de Rezende, J. F., and Duarte, O. C. M. B., “Placement Issues in Measurement-Based Host Location”, 9th European Summer School and IFIP Workshop on Next Generation Networks - EUNICE 2003, pp. 173-180, Budapest - Balatonfured, Hungary, September 200. English, A4 size, 8 p., file: ZFRD03c.ps.gz

GTA-03-32 Velloso, P. B., Rubinstein, M. G., and Duarte, O. C. M. B., “The effect of mobility on voice transmission capacity in mobile ad hoc networks”, The 3rd IEEE Latin American Network Operations and Management Symposium - LANOMS’2003, pp. 63-70, Foz do Igua? Brazil, September 2003. English, A4 size, 8 p., file: VRD03c.ps.gz

GTA-03-33 Cunha, D. O., Costa, L. H. M. K., and Duarte, O. C. M. B., “An Energy-Aware Routing Mechanism for Mobile Ad Hoc Networks”, The 3rd IEEE Latin American Network Operations and Management Symposium - LANOMS’2003, Foz do Igua? Brazil, September 2003. English, A4 size, 8 p., file: CCD03a.ps.gz

GTA-03-34 Rocha, L. G. S., Costa, L. H. M. K., and Duarte, O. C. M. B., “Analyzing the Impact of Misbehaving Nodes in Ad Hoc Routing”, The 3rd IEEE Latin American Network Operations and Management Symposium - LANOMS’2003, pp. 5-12, Foz do Igua? Brasil, September 2003. English, A4 size, 8 p., file: RCD03a.ps.gz

GTA-03-35 Master Science Thesis Author - Pedro Braconnot Velloso Title - “Transmiss?de Voz em Redes Ad Hoc" Advisors - Otto Carlos Muniz Bandeira Duarte and Marcelo Gon?ves Rubinstein COPPE/PEE/UFRJ - August 2003 Portuguese - A4 size - 64 p. - Directory:

GTA-03-36 Undergraduate Final Project Author - Anelise Godolphim Sacks Title - “Sistema de Gerenciamento de Redes e Processos atrav?de Computadores Port?is via Bluetooth” Advisor - Jos?erreira de Rezende DEL/EE/UFRJ - July 2003 Portuguese - A4 size - 98 p. - Directory: Anelise03

GTA-03-37 Stochero, J. M, Pinto, A. J. G., and de Rezende, J. F. - “Roteamento com Agrega? de Dados em Redes de Sensores”, I Escola Regional de Redes de Computadores (ERRC’03), pp. 103-108, Porto Alegre, September 2003. Portuguese, A4 size, 6 p., file: SPR03.pdf

GTA-03-38 Fernandez, M.P., Pedroza, A.C.P., and de Rezende, J. F. - “Implementa? de Pol?icas de Gerenciamento com L??a Fuzzy e Algoritmo Gen?co visando Melhoria da Qualidade de Servi?(QoS)”, to appear in Revista Brasileira de Telecomunica?, vol. 18, n. 2, October 2003. Portuguese, A4 size, 15 p. file: FPR03c.pdf

GTA-03-39 Cunha, D. O., Costa, L. H. M. K., and Duarte, O. C. M. B. - Energia em Redes Ad Hoc”, V Workshop de Comunica?s sem Fio e Computa? M?? - WCSF’2003, pp. 115-124, S?Louren? Brazil, October 2003. Portuguese, Letter size, 10 p., file: CCD03b.ps.gz

GTA-03-40 Viana, A. C., Amorim, M. D. de, Fdida, S., and de Rezende, J. F. de - “Routage bas?ur ancre dans les r?aux ?large ?elle auto-organisables”, Colloque Francophone sur l’Ing?erie des Protocoles - CFIP’2003, pp. 393-408, Paris, France, October 2003. French, A4 size, 16 p., file: VAFR03b.ps.gz

GTA-03-41 Ribeiro, F. L., Lopes, J. C. R., and Pedroza, A. C. - “Analise dos Processos em Sistemas Moveis de 3a Geracao”, I Escola Regional de Redes de Computadores - ERRC 2003, Porto Alegre, RS, Brazil, September 2003. Portuguese, A4 size, 6 p., file: RiLoPe03a.pdf GTA-03-42 Doctor of Science Thesis

Author - Paulo Cesar Salgado Vidal Title - “Agrega? de Estados do Protocolo RSVP para Fluxos Multicast” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - September 2003 Portuguese - A4 size - 121 p. - file: Vidal03.tar.gz

GTA-03-43 Rubinstein, M. G and de Rezende, J. F. - “Quality of Service in Ad Hoc 802.11 Networks”, Journal of the Brazilian Computer Society, vol. 9, no. 1, pp. 41-52, ISSN 0104-6500, Brazilian Computer Society, July 2003. English, A4 size, 12 p., file: RuRe02a.ps.gz

GTA-03-44 Ribeiro, F. J. L., Lopes, J. C. R., and Pedroza, A. C. - “An?se do Estabelecimento de Associa?s Seguras em Sistemas M??s de 3a Gera?”, 2nd International Information and Telecommunication Technologies Symposium - I2TS’2003, Florian??is, SC, Brazil, November 2003. Portuguese, Letter size, 7 p., file: RiLoPe03b.pdf

GTA-03-45 Doctor of Science Thesis Author - Artur Ziviani Title - “Qualité de Service et Localisation d’Hôtes (Quality of Service and Location-Awarenessi)” Supervisors - Serge Fdida, Otto Carlos Muniz Bandeira Duarte, and José Ferreira de Rezende Université Pierre et Marie Curie (Paris 6), France - December 2001 English - A4 size - 148 p. - Directory: Ziviani03

GTA-03-46 Pereira, I.C.M., Pedroza, A.C.P. - “Aplica?s Militares Empregando Redes M??s Ad Hoc”, VI Simp?? de Pesquisa Operacional e VII Simp?? de Log?tica da Marinha, SPOLM 2003, Rio de Janeiro, RJ, Brazil, 2003 Portuguese, A4 size, 14p., file: PP04b.pdf.gz

GTA-03-47 Ara?? S.G., Mesquita A. and Pedroza, A.C.P. - “Evolutionary Synthesis of Communication Protocols”, in Proceedings of the 10th International Conference on Telecommunications (ICT’03), pp. 986-993, Tahiti, French Polynesia, February 2003, file: AMP03a.pdf.gz.

GTA-03-48 Ara?? S. G., Mesquita A. C., and Pedroza, A.C.P. - “Using Genetic Programming and High Level Synthesis to Design Optimized Datapath”, in Proceedings of the 5th International Conference of Evolvable Systems (ICES’03), LNCS, Vol. 2606, pp. 434-445, Trondheim, Norway, March 2003, file: AMP03b.pdf.gz.

GTA-03-49 Ara?? S.G., Mesquita A. C., and Pedroza, A.C.P. - “Optimized Datapath Design by Evolutionary Computation”, in Proceedings of the 3rd IEEE International Workshop on System-on-Chip for Real-Time Applications (IWSOC’03), pp. 06-10, Calgary, Canada, June 2003, file: AMP03c.pdf.gz.

GTA-03-50 Ara?? S.G., Mesquita A. C., and Pedroza, A.C.P. - “S?tese de Circuitos Digitais Otimizados via Programa? Gen?ca”, 30o Semin?o Integrado de Software e Hardware (SEMISH’03), Vol. III pp. 273-285, Campinas, Brasil, Agosto de 2003, file: AMP03d.pdf.gz.

GTA-03-51 Ara?? S.G., Mesquita A. C., and Pedroza, A.C.P. - “Uma Metodologia de Projeto de Protocolos de Comunica? Baseada em T?icas Evolutivas”, 20o Simp?? Brasileiro de Telecomunica?s (SBT’03), Rio de Janeiro, Brasil, Outubro de 2003, AMP03e.pdf.gz.

GTA-02-01 Cardoso, K. V., de Rezende, J. F. and Fonseca, N. L. S. da - “On the Effectiveness of Push-out Mechanisms for the Discard of TCP Packets”, International Conference on Communications 2002, New York City, NY, USA, April 2002. English, Letter size, 5 p., file: CRF02a.ps.gz

GTA-02-02 Costa, L. H. M. K., Fdida, S., and Duarte, O. C. M. B. - “Developing Scalable Protocols for Three-metric QoS-routing”, Computer Networks Journal, vol. 39, n. 6, pp. 713-727, ISSN 1389-1286, Elsevier Science, August 2002. English, Letter size, 15 p., file: CFD02.ps.gz

GTA-02-03 Ziviani, A., de Rezende, J. F., Duarte, O. C. M. B., and Fdida, S. - “Services Diff?nci?pour les Flux MPEG”, Journ? Doctorales Informatique et R?aux - JDIR’2002, pp. 51-60, Toulouse, France, March 2002. French, A4 size, 10 p., file: ZRDF02a.ps.gz

GTA-02-04 Master Science Thesis Author - Flavio Duboc de Mendonca Title - “Implementa? de Sistemas com Restri?s Temporais a partir de uma Especifica? Formal” Advisor - Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - March 2002 Portuguese - A4 size - 191 p. - File: Duboc01.ps.gz

GTA-02-05 Rubinstein, M. G. and de Rezende, J. F. - “Qualidade de Servi?em Redes 802.11”, XX Simp?? Brasileiro de Redes de Computadores (SBRC2002), B??s, RJ, Brazil, May 2002. Portuguese, A4 size, 16 p., file: RuRe02.pdf

GTA-02-06 Anton, E. R. and Duarte, O. C. M. B. - “Estabelecimento de Chave de Grupo em Redes Ad Hoc”, Workshop em Seguran?de Sistemas Computacionais (WSeg2002), B??s, RJ, Brazil, May 2002. Portuguese, A4 size, 8 p., file: AnDu02.ps.gz

GTA-02-07 Velloso, P. B., Rubinstein, M. G., and Duarte, O. C. M. B. - “Uma Ferramenta de Gerenciamento de Redes Baseada em Agentes M??s”, XIV Congresso Brasileiro de Autom?ca, pp. 679-684, Natal, RN, Brazil, September 2002. Portuguese, A4 size, 8 p., file: VRD02.ps.gz

GTA-02-08 Gonçalves, P. A. da S., de Rezende, J. F., Duarte, O. C. M. B., and Pujolle, G. - “Optimal Feedback for Quality Source-Adaptive Schemes in Multicast Multi-layered Video Environments”, Lecture Notes in Computer Science - Networking’2002, pp. 563-574, Vol. 2345, ISSN 0302-9743, Springer, May 2002. English, A4 size, 12 p., file: GRDP02.ps.gz

GTA-02-09 Rubinstein, M. G., Duarte, O. C. M. B., and Pujolle, G. - “Scalability of a Mobile Agents Based Network Management Application”, IEEE/KICS Journal of Communications and Networks, vol. 5, no. 3, pp. 240-248, ISSN 1229-2370, September 2003. English, A4 size, 9 p.

GTA-02-10 Anton, E. R. and Duarte, O. C. M. B. - “Seguran?em Redes Sem Fio Ad Hoc: Gerenciamento de Chave de Grupo”, XIV Congresso Brasileiro de Autom?ca, Natal, RN, Brazil, September 2002. Portuguese, A4 size, 6 p., file: AnDu02a.ps.gz

GTA-02-11 Anton, E. R. and Duarte, O. C. M. B. - “Group Key Establishment in Wireless Ad Hoc Networks”, Workshop em Qualidade de Servio e Mobilidade - WQoSM 2002, Angra dos Reis, RJ, Brazil, November 2002. English, A4 size, 10 p., file: AnDu02b.pdf

GTA-02-12 Ziviani, A., Wolfinger, B. E., de Rezende, J. F., Duarte, O. C. M. B., and Fdida, S. “On the Combined Adoption of QoS Schemes to Improve the Delivery Quality of MPEG Video Streams”, Intern. Symp. on Performance Evaluation of Computer and Telecommunications Systems - SPECTS’2002, San Diego, USA, July 2002. English, Letter, 8 p., file: ZWRDF02.ps.gz

GTA-02-13 Cardoso, K. V. and de Rezende, J. F. - “HTTP Traffic Modeling: Development and Application”, IEEE International Telecommunications Symposium (ITS2002), Natal, RN, Brasil, September 2002. English, A4 size, 6 p., file: CaRe02a.ps.gz

GTA-02-14 Fernandez, M. P., Pedroza, A. C. P., and de Rezende, J. F. - “Otimiza? de Controlador Fuzzy para Provisionamento de Recursos em ambiente DiffServ atrav?de Algoritmo Gen?co”, Simposio Brasileiro de Redes de Computadores (SBRC2002), B??s, RJ, Brazil, May 2002. Portuguese, A4 size, 16 p., file: FPR02.ps.gz

GTA-02-15 Fernandez, M. P., Pedroza, A. C. P., and de Rezende, J. F. - “Optimizing Fuzzy Controllers with Genetic Algorithms for QoS Improvement”, IEEE International Telecommunications Symposium (ITS2002), Natal, RN, Brasil, September 2002. English, A4 size, 6 p., file: FPR02a.ps.gz

GTA-02-16 Miranda, M. N., Lima, R. N. B., Pedroza, A. C. P., and Mesquita F, A. C. M. - “Design of a Reliable Multicast Protocol Using HW/SW Codesign Based on Performance Optimization with Genetic Algorithms”, IEEE International Telecommunications Symposium (ITS2002), Natal, RN, Brasil, September 2002. English, A4 size, 6 p., file: MBAM02a.ps.gz

GTA-02-17 Lima, R. N. B., Miranda, M. N., Pedroza, A. C. P., and Mesquita F, A. C. M. - “HW/SW Codesign of Handoff Protocol for Wireless ATM Networks based on Performance Optimization using Genetic Algorithms”, SBCCI 2002 - 15th Symposium on Integrated Circuits and System Design, Porto Alegre, RS, Brasil, September 2002. English, A4 size, 6 p., file: MBAM02b.ps.gz

GTA-02-18 Amorim, M. D., Duarte, O. C. M. B., and Pujolle, G., “Improving User Satisfaction in Adaptive Multicast Video”, IEEE/KICS Journal of Communications and Networks, ISSN 1229-2370, vol 4, no 3, pp. 221-229, September 2002. English, A4 size, 9 p., file: ADP02a.ps.gz

GTA-02-19 Master Science Thesis Author - Kleber Vieira Cardoso Title - “Modelagem de Tr?go HTTP: Desenvolvimento e Aplica?”

Supervisor - Jos?erreira de Rezende COPPE/PEE/UFRJ - July 2002 Portuguese - A4 size - 89 p. - Directory: Cardoso02

GTA-02-20 Ziviani, A., Fdida, S., de Rezende, J. F., and Duarte, O. C. M. B., “Enhancing Location Management in Mobile Ad Hoc Networks”, IFIP Mediterranean Ad Hoc Networking Workshop – Med-Hoc-Net 2002, Sardegna, Italy, September 2002. English, A4 size, 7 p., file: ZFRD02.ps.gz

GTA-02-21 Amorim, M. D., Duarte, O. C. M. B., and Pujolle, G. - “Distinguishing Video Quality Through Differential Matrices”, ACM Multimedia System Journal, Special Issue on Multimedia Distribution, vol. 9, no. 1, pp. 94-103, ISSN 0942-4962, Springer, July 2002. English, A4 size, 9 p., file: ADP02b.ps.gz

GTA-02-22 Costa, L. H. M. K., Fdida, S., and Duarte, O. C. M. B. - “Analyzing the Effects of Asymmetric Unicast Routes on Multicast Routing Protocols”, Journal of the Brazilian Computer Society, vol. 17, n. 2, pp. 5-15, ISSN 0104-6500, Brazilian Computer Society, 2001. English, A4 size, 11 p., file: CFD02a.ps.gz

GTA-02-23 Amorim, M. D., Duarte, O. C. M. B., and Pujolle, G. - “Virtual Layering and Efficient Merging in Non-cooperative Multicast Trees”, Journal of the Brazilian Computer Society, vol. 17, n. 2, pp. 28-38, ISSN 0104-6500, Brazilian Computer Society, 2001. English, A4 size, 10 p., file: ADP02c.ps.gz

GTA-02-24 Bagatelli R., Moura, D. F. C., and Pedroza, A. C. P. - “Especifica? Formal de uma Arquitetura de Suporte ?Descoberta de Servi? em Redes M??s Ad Hoc”, V Workshop de M?dos Formais (WMF’2002), Gramado, RS, October 2002. Portuguese, A4 size, 16 p., file: BMP02.ps.gz

GTA-02-25 Doctor of Science Thesis Author - Alexandre Sztajnberg Title - “Flexibilidade e Separa? de Interesses para a Concep? e Evolu? de Aplica?s Distribu?as” Advisors - Jorge Lopes de Souza Leao and Orlando Loques (CAA/UFF) COPPE/PEE/UFRJ - May 2002 Portuguese - A4 size - 230 p. - file: Sztajnberg02.tar.gz

GTA-02-26 Master Science Thesis Author - Myrna Cec?ia Martins dos Santos Title - “An?se Formal de Protocolos de Autentica? para Redes Celulares” Advisors - Jos?erreira de Rezende and Jos?nt?io Moreira Xex? IME/UFRJ - June 2002 Portuguese - A4 size - 128 p. - File: Myrna02.pdf

GTA-02-27 Master Science Thesis Author - Glauco Fiorott Amorim Title - “An?se de Desempenho de Protocolos de Roteamento com Diferencia? de Servi? em Redes de Comunica? M?? Ad Hoc” Advisor - Jos?erreira de Rezende IME/UFRJ - July 2002 Portuguese - A4 size - 127 p. - File: Glauco02.pdf GTA-02-28 Santos, M. C. M, Xex? J. A. M., and de Rezende, J. F. - “An?se Formal de

Protocolos de Seguran?para Redes Celulares”, IV Workshop de Comunica? Sem Fio (WCSF’02), pp. 78-89, S?Paulo, SP, October 2002. Portuguese, A4 size, 12 p., file: SXR02.pdf

GTA-02-29 Priess, W., de Rezende, J. F., Carmo, L. F. R. C., and Pirmez, L. - “BlueNetS: Uma Ferramenta para o Estudo de Desempenho do Bluetooth”, IV Workshop de Comunica? Sem Fio (WCSF’02), pp. 161-171, S?Paulo, SP, October 2002. Portuguese, A4 size, 11 p., file: PRCP02.ps.gz

GTA-02-30 Filho, J. C. M., de Rezende, J. F., Pirmez, L., and Carmo, L. F. R. C. - “Controle de Admiss?e Controle de Carga para Redes IEEE 802.11 Infra-Estruturadas”, IV Workshop de Comunica? Sem Fio (WCSF’02), pp. 100-111, S?Paulo, SP, October 2002. Portuguese, A4 size, 12 p., file: FRPC02.ps.gz

GTA-02-31 Master Science Thesis Author - Saulo Vaz de Vasconcellos Title - “Provisionamento de Recursos e QoS em redes de N??o IP para Sistemas Celulares de 3^{a} Gera?” Supervisor - Jos?erreira de Rezende COPPE/PEE/UFRJ - September 2002 Portuguese - A4 size - 119 p. - Directory: Vasc02

GTA-02-32 Velloso, P. B., Rubinstein, M. G., and Duarte, O. C. M. B. - "Transmiss?de Voz em Redes Ad Hoc”, Workshop em Qualidade de Servi?e Mobilidade - WQoSM 2002, Angra dos Reis, RJ, Brazil, November 2002. Portuguese, A4 size, 12 p., file: VRD02a.pdf

GTA-02-33 Gon?ves, P. A. S., Amorim, M. D., de Rezende, J. F., and Duarte, O. C. M. B. - “QoS Mechanisms for Multilayered Multicast Applications”, Tenth International Conference on Telecommunications Systems, pp. 131-143, Monterey, USA, October 2002. English, A4 size, 13 p., file: GARD02a.ps.gz

GTA-02-34 Ziviani, A., Fdida, S., de Rezende, J. F., and Duarte, O. C. M. B. - “Placing Landmarks to Locate Internet Hosts”, Workshop em Qualidade de Servi?e Mobilidade - WQoSM 2002, Angra dos Reis, RJ, Brazil, English, A4 size, 16 p., file: ZFRD02b.ps.gz

GTA-02-35 Ziviani, A., Wolfinger, B. E., de Rezende, J. F., Duarte, O. C. M. B., and Fdida, S. - “Joint Adoption of QoS Schemes for MPEG Streams”, to appear in Multimedia Tools and Applications, Kluwer, ISSN: 1380-7501, 2003. English, A4 size, 35 p., file: ZWRDF02b.ps.gz

GTA-02-36 Gon?ves, P. A. S., Amorim, M. D., de Rezende, J. F., Duarte, O. C. M. B., Pujolle, G. - “Agrega? e Supress?de Feedback de Grupos Multiponto Heterog?os”, Workshop em Qualidade de Servi?e Mobilidade - WQoSM 2002, Angra dos Reis, RJ, Brazil, Portuguese, A4 size, 19 p., file: GARDP02b.ps.gz

GTA-02-37 Vasconcellos, S. V. and de Rezende, J. F. - “QoS and mobility in 3G Core Networks”, Workshop em Qualidade de Servi?e Mobilidade - WQoSM 2002, Angra dos Reis, RJ, Brazil, November 2002. English, A4 size, 14 p., file: VaRe02.ps.gz GTA-02-38 Cardoso, K. V. and de Rezende, J. F. - “Design and Use of an aggregated HTTP

Traffic Model”, Workshop em Qualidade de Servi?e Mobilidade - WQoSM 2002, Angra dos Reis, RJ, Brazil, November 2002. English, A4 size, 14 p., file: CaRe02b.pdf

GTA-02-39 Master Science Thesis Author - Jos?alentim dos Santos Filho Title - “Um Simulador para Estelle Tempo-Real” Advisors - Aloysio de Castro Pinto Pedroza e Jorge Lopes de Souza Le? COPPE/PEE/UFRJ - October 2002 Portuguese - A4 size - 88 p. - Directory: Santos02

GTA-02-40 Doctor of Science Thesis Author - Marcial Porto Fernandez Title - “Provisionamento de Recursos em Arquitetura Diffserv para Melhoria da Qualidade de Servi?(QoS)” Advisors - Aloysio Castro P. Pedroza and Jos?erreira de Rezende COPPE/PEE/UFRJ - October 2002 Portuguese - A4 size - 168 p. - file: Fernandez02.ps.gz

GTA-02-41 Rubinstein, M. G. and de Rezende, J. F. - “Qualidade de Servi?no Controle de Acesso ao Meio de Redes 802.11”, Workshop em Qualidade de Servi?e Mobilidade - WQoSM 2002, Angra dos Reis, RJ, Brazil, November 2002. Portuguese, A4 size, 11 p., file: RuRe02b.pdf

GTA-02-42 de Rezende, J. F. de, Lima, M. M. de A. E. and Fonseca, N. L. S. da. - Mobility over Transport Control Protocol/Internet Protocol (TCP/IP). In: ILYAS, Mohammad. (Org.). The Handbook of Ad Hoc Wireless Networks. Boca Raton, Florida, 2002. English - Letter - 15 p.

GTA-02-43 Undergraduate Final Project Author - Bernardo Antunes Maciel Villela Title - “Uma an?se do Roteamento em Redes Ad Hoc” Advisor - Otto Carlos Muniz Bandeira Duarte DEL/EE/UFRJ - December 2002 Portuguese - A4 size - 57 p. - Directory: Bernardo02

GTA-02-44 Rocha, L. G. S. and Duarte, O. C. M. B. - “Aspectos e Mecanismos de Seguran?em Redes Ad Hoc”, Workshop em Qualidade de Servi?e Mobilidade - WQoSM 2002, Angra dos Reis, RJ, Brazil, November 2002. Portuguese, A4 size, 8 p., file: RoDu2.pdf

GTA-02-45 Vidal, P. C. S. and Duarte, O. C. M. B. - “Multicast Flow Aggregation in IP Tunnels with Resource Reservation”, International Information Technology Symposium - I2TS 2002, Florian??is, SC, Brazil, October 2002. English, A4 size, 8 p., file: ViDu02a.pdf.gz

GTA-02-46 Vidal, P. C. S. and Duarte, O. C. M. B. - “Estabelecimento de Reserva de Recursos para Fluxos Multicast em T??s IP-sobre-IP”, V Simp?? Brasileiro de Sistemas Multim?ia e Hiperm?ia - SBMIDIA 2002, pp. 252-259, Fortaleza, CE, Brazil, October 2002. Portuguese, A4 size, 8 p., file: ViDu02b.ps.gz

GTA-02-47 Doctor of Science Thesis Author - Marcio Nunes de Miranda Title - “Uma Metodologia de Hardware/Software Codesign de Protocolos de Comunica? Baseada na Otimiza? de Desempenho por Algoritmos Gen?os” Advisors - Aloysio Castro P. Pedroza and A. C. Mesquita Filho COPPE/PEE/UFRJ - August 2002 Portuguese - A4 size - 157 p. - file: Mira02.ps.gz

GTA-02-48 Undergraduate Final Project Author - Ingrid Teixeira Title - “Sistema de Informa?s M??s para Departamento de Combust?eis” Advisors - Aloysio Castro P. Pedroza and Antonio. C. G. de Souza DEL/EE/UFRJ - September 2002 Portuguese - A4 size - 157 p. - file: Ingrid02.ps.gz GTA-01-01 Rubinstein, M. G. and Duarte, O. C. M. B. - “Implementa? e Simula? de uma

Aplica? Gerenciamento Utilizando Agentes M??s”, 19^{o} Simp?? Brasileiro de Telecomunica?s, Fortaleza, CE, Brazil, September 2001. Portuguese, A4 size, 6 p., file: RuDu01.ps.gz

GTA-01-02 Doctor of Science Thesis Author - Marcelo Gon?ves Rubinstein Title - “Avalia? do Desempenho de Agentes M??s no Gerenciamento de Redes” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - March 2001 Portuguese - A4 size - 101 p. - File: Rubi01/Rubi01.ps.gz

GTA-01-03 Viana, A. C., Guedes, L. G. R., de Rezende, J. F., and Duarte, O. C. M. B. - “Avalia? de um Ambiente de Diferencia? de Servi? Proporcional Tr?go TCP”, XIX Simp?? Brasileiro de Redes de Computadores SBRC’2001, pp. 226 – 241, Florian??is, SC, May 2001. Portuguese, A4 size, 16 p., file: AGRD01.ps.gz

GTA-01-04 Vasconcellos, S. V. and de Rezende, J. F. - “Mobilidade num ambiente de servi? diferenciados”, XIX Simp?? Brasileiro de Redes de Computadores SBRC’2001, pp. 242 – 257, Florian??is, SC, May 2001. Portuguese, A4 size, 16 p., file: VaRe01a.ps.gz

GTA-01-05 Cardoso, K. V., Vasconcellos, S. V., de Rezende, J. F., and Fonseca, N. L. S. da - “Avaliando Estrat?as de Push-out no Descarte Seletivo de Pacotes de Tr?go TCP de Curta e Longa Dura?”, XIX Simp?? Brasileiro de Redes de Computadores SBRC’2001, pp. 638 – 652, Florian??is, SC, May 2001. Portuguese, A4 size, 15 p., file: CVRF01a.ps.gz

GTA-01-06 Undergraduate Final Project Author - Eric Ricardo Anton Title - “Projeto e Implementa? de um Protocolo para Transmiss?Multidestinat?a Segura” Advisor - Otto Carlos Muniz Bandeira Duarte DEL/EE/UFRJ - March 2001 Portuguese - A4 size - 125 p. - Directory: Eric01

GTA-01-07 Fernandez, M.P., Pedroza, A.C.P. and de Rezende, J. F. - “Qualidade de Servico em um Dominio DiffServ atraves de Gerenciamento Baseado em Politicas”, XIX Simp?? Brasileiro de Redes de Computadores SBRC’2001, pp. 590 – 605, Florian??is, SC, May 2001. Portuguese, A4 size, 16 p., file: FPR01.ps.gz

GTA-01-08 Costa, L. H. M. K., Fdida, S., and Duarte, O. C. M. B. - “HBH: Um Protocolo de Roteamento para a Implanta? Progressiva do Servi?Multicast”, XIX Simp?? Brasileiro de Redes de Computadores SBRC’2001, pp. 448 – 463, Florian??is, SC, May 2001. Portuguese, A4 size, 16 p., file: CFD01a.ps.gz

GTA-01-09 Amorim, M. D., Duarte, O. C. M. B., and Pujolle, G. - “Otimiza? da Qualidade de V?eo em Redes Multicast n?Cooperativas”, XIX Simp?? Brasileiro de Redes de Computadores SBRC’2001, pp. 434 – 447, Florian??is, SC, May 2001. Portuguese, A4 size, 15 p., file: ADP01b.ps.gz GTA-01-10 Costa, L. H. M. K., Fdida, S., and Duarte, O. C. M. B. - “Hop-by-hop Multicast

Routing Protocol”, ACM SIGCOMM, August 2001. English, A4 size, 14 p., file: CFD01b.ps.gz

GTA-01-11 Miranda, M. N., Lima, R. N. B., Pedroza, A. C. P., and Mesquita F^{o}, A. C. - “HwSw Codesign de Protocolos Baseado na Otimização de Desempenho por Algoritmos Genéticos”, XIX Simpósio Brasileiro de Redes de Computadores SBRC’2001, pp. 727 – 742, Florianópolis, SC, May 2001. Portuguese, A4 size, 16 p., file: MBAM01a.ps.gz

GTA-01-12 Master Science Thesis Author - Rodrigo de Carvalho Roman Title - “Um Editor Visual para o Aux?io a Especifica? de Protocolos de Comunica? com Restri? Temporal” Advisors - Aloysio Castro Pinto Pedroza and Jorge Lopes de Souza Le? COPPE/PEE/UFRJ - June 2001 Portuguese - A4 size - 136 p. - File: Roman01.ps.gz

GTA-01-13 Costa, L. H. M. K., Fdida, S., and Duarte, O. C. M. B. - “Enabling the Progressive Multicast Service Deployment”, IEEE Symposium on Computers and Communications ISCC’01, Tunisia, July 2001. English, A4 size, 6 p., file: CFD01c.ps.gz

GTA-01-14 Master Science Thesis Author - Aline Carneiro Viana Title - “Avalia? de um Ambiente de Diferencia?o de Servi?Proporcional com Tr?go TCP” Advisors - Jos?erreira de Rezende, Leonardo Guerra de Rezende Guedes, and Otto Carlos Muniz Bandeira Duarte EEE/UFG e COPPE/PEE/UFRJ - April 2001 Portuguese - A4 size - 95 p. - Directory: Aline01

GTA-01-15 Miranda, M. N., Lima, R. N. B., Pedroza, A. C. P., and Mesquita F^{o}, A.C. - “HW/SW Codesign of Protocols Based on Performance Optimization Using Genetic Algorithms”, 17^{th} International Teletraffic Congress, Salvador, BA, September 2001. English, A4 size, 16 p., file: MBAM01b.ps.gz

GTA-01-16 Amorim, M. D., Achir, N., Duarte, O. C. M. B., and Pujolle, G. - “Active Bandwidth Sharing for Multi-layered Video in Multi-Source Environments”, Networking and Information Systems Journal, Hermes Science Publications, vol. 3, no. 1, pp. 123-133, ISSN 1290-2926, 2000. English, A4 size, 11 p., file: AADP01a.ps.gz

GTA-01-17 Amorim, M. D., Duarte, O. C. M. B., and Pujolle, G. - “Improving the Quality Specification of Multi-Layered Applications”, Networking and Information Systems Journal, Hermes Science Publications, vol. 2, no. 6, ISSN 1290-2926, 2000. English, A4 size, 14 p., file: ADP01a.ps.gz

GTA-01-18 Doctor of Science Thesis Author - Marcelo Dias de Amorim Title - ‘Distribution de Flux en Couches dans les R?aux Multicast H?rog?es’ Advisors - Guy Pujolle and Otto Carlos Muniz Bandeira Duarte Universit?e Versailles, France - June 2001 French - A4 size - 140 p. - File: Amorim01/Amorim.ps.gz GTA-01-19 Vasconcellos, S. V. and de Rezende, J. F. - “Using Differentiated Services in 3G

Cellular Networks”, The 3rd. IEEE International Conference on Mobile and Wireless Communication Networks - MWCN’2001, Recife, PE, Brazil, August 2001. English, A4 size, 8 p., file: VaRe01b.ps.gz

GTA-01-20 Ziviani, A., de Rezende, J. F., Duarte, O. C. M. B., and Fdida, S. - “V?eo MPEG sobre Servi? Diferenciados”, XIX Simp?? Brasileiro de Telecomunica?s - SBrT’01, Fortaleza, CE, September 2001. Portuguese, A4 size, 6 p., file: ZRDF01b.ps.gz

GTA-01-21 Amorim, M. D., Duarte, O. C. M. B., and Pujolle, G. - “Application-aware Multicast”, IEEE Global Communications Conference - Globecom’01, November 2001. English, A4 size, 5 p., file: ADP01c.ps.gz

GTA-01-22 Fernandez, M. P., Pedroza, A. C. P., and de Rezende, J. F. - “Policy-Based Management providing QoS in a DifServ Domain”, IEEE Latin American Network Operations and Management Symposium - LANOMS’01, pp. 14-25, Belo Horizonte, MG, Brazil, August 2001 English, A4 size, 11 p., file: FPR01b.ps.gz

GTA-01-23 Undergraduate Final Project Author - Pedro Braconnot Velloso Title - “Projeto e implementa? de uma aplica? de gerenciamento de redes baseada em agentes m??s” Advisor - Otto Carlos Muniz Bandeira Duarte and Marcelo Gon?ves Rubinstein DEL/EE/UFRJ - October 2001 Portuguese - A4 size - 60p. - Directory: Pedro01

GTA-01-24 Rubinstein, M. G., Duarte, O. C. M. B., and Pujolle, G. - “Evaluating the Performance of a Network Management Application Based on Mobile Agents”, Lectures Notes in Computer Science 2345, Enrico Gregori, Marco Conti, Andrew Campbell, Guy Omidyar, and Moshe Zukerman (Eds.), pp. 515-526, ISSN 0302-9743, ISBN 3-540-43709-6, Springer-Verlag, May 2002. English, A4 size, 12 p., file: RDP01.ps.gz

GTA-01-25 Doctor of Science Thesis Author - Lu? Henrique Maciel Kosmalski Costa Title - “Routing in the Internet: Quality of Service and Group Communication” Supervisors - Serge Fdida and Otto Carlos Muniz Bandeira Duarte Universit?ierre et Marie Curie (Paris 6), France - December 2001 English - A4 size - 140 p. - Directory: Costa01

GTA-01-26 Fernandez, M.P., Pedroza, A. C. P., and de Rezende, J. F. - “QoS Provisioning across a DiffServ Domain using Policy-Based Management”, IEEE Global Telecommunications Conference - Globecom’2001, pp. 378-382, San Antonio, Texas, USA, November 2001. English, Letter size, 6 p., file: FPR01c.ps.gz

GTA-01-27 Fernandez, M. P., Pedroza, A. C. P., and de Rezende, J. F. - “Quality of Service in a DiffServ Domain using Policy-Based Management”, 17th International Teletraffic Congress, Salvador, BA, Brazil, December 2001. English, A4 size, 12 p., pp. 919-930 file: FPR01d.ps.gz GTA-01-28 Master Science Thesis

Author - Igor Briglia Habib de Almeida Alves Title - “Marca? de Tr?go para Justi?em Fluxos Agregados no Servi?Assegurado" Advisors - Jos?erreira de Rezende and Luis Felipe Magalh? de Moraes COPPE/PEE/UFRJ - October 2001 > Portuguese - A4 size - 184 p. - File: Igor01.pdf GTA-00-01 Rubinstein, M. G., Duarte, O. C. M. B., and Pujolle, G. - “Using Mobile Agent

Strategies for Reducing the Response Time in Network Management”, 16th IFIP World Computer Congress, ICCT2000, pp. 278-281, Beijing, China, August 2000. English, A4 size, 4 p., file: RDP00.ps.gz

GTA-00-02 Rubinstein, M. G., Duarte, O. C. M. B., and Pujolle, G. - “Reducing the Response Time in Network Management by Using Multiple Mobile Agents”, in Managing QoS in Multimedia Networks and Services, Jos?euman de Souza and Raouf Boutaba (Ed.), pp. 253-265, ISBN 0-7923-7962-4, Kluwer Academic Publishers, September 2000. English, A4 size, 13 p., file: RDP00a.ps.gz

GTA-00-03 Delfino, G.M., Santos F^{∘}, J.V., Roman, R.C., Le? J. L.S., and Pedroza, A.C.P. - “Sistema Integrado de Aux?io ao Projeto de Protocolos com Restri?s Temporais (Prot-CAD/RT)” - February 2000. Portuguese, A4 size, 15 p., file: DSR00a.ps.gz

GTA-00-04 Amorim, M. D., Duarte, O. C. M. B., and Pujolle, G. - “Multi-criteria arguments for improving the fairness of layered multicast applications”, Broadband Communications, High Performance Networking, and Performance of Communication Networks - Networking 2000, Lectures Notes in Computer Sciences n^{∘} 1815, Edited by G. Pujolle, H. Perros, S. Fdida, U. Körner,and I. Stavrakakis, pp. 1-10, ISSN 0302-9743, Springer, 2000. English, A4 size, 5 p., file: ADP00b.ps.gz

GTA-00-05 Amorim, M. D., Duarte, O. C. M. B., and Pujolle, G. - “An improved MPEG behavioral analysis with autonomous parameter-extracting algorithm for strict video applications”, IEEE International Conference on Communications ICC’2000, pp. 831-835, New Orleans, USA, June 18-22 2000. English, A4 size, 5 p., file: ADP00b.ps.gz

GTA-00-06 Master Science Thesis Author - Paulo Andr?a Silva Gon?ves Title - “Um Servi?Ativo de Distribui? de V?eo Multiponto” Advisors - Otto Carlos Muniz Bandeira Duarte and Jos?erreira de Rezende COPPE/PEE/UFRJ - March 2000 Portuguese - A4 size - 82 p. - Directory: Goncalves00

GTA-00-07 Gon?ves, P. A. da S., de Rezende, J. F., and Duarte, O. C. M. B. - “Um Servi?Ativo de Distribui? de V?eo Multiponto”, XVIII Simp?? Brasileiro de Redes de Computadores SBRC’2000, pp. 211-226, Belo Horizonte, MG, May 2000. Portuguese, A4 size, 16 p., file: GRD00.ps.gz

GTA-00-08 Ziviani, A., de Rezende, J. F., and Duarte, O. C. M. B. - “Avaliando o Suporte ao Tr?go de Voz em um Ambiente de Diferencia? de Servi?”, VI Simp?? Brasileiro de Multim?ia e Sistemas Hiperm?ia - SBMIDIA’2000, pp. 341-356, Natal, RN, June 2000. Portuguese, A4 size, 16 p., file: ZRD00.ps.gz

GTA-00-09 Alves, I. B. H. A., de Rezende, J. F., and Moraes, L. F. M. - “Avaliando a Justi?na Marca? de Tr?gos Agregados”, XVIII Simp?? Brasileiro de Redes de Computadores SBRC’2000, pp. 119-134, Belo Horizonte, MG, May 2000. Portuguese, A4 size, 16 p., file: ARM00a.ps.gz GTA-00-10 Miranda, M.N., Pedroza, A.C.P., and Mesquita F^{o}, A.C. - “Uma Metodologia para

Projeto de Protocolos Utilizando Algoritmos Gen?cos e Modelos Markovianos”, February 2000. Portuguese, A4 size, 14 p., file: MPM00.ps.gz

GTA-00-11 Rubinstein, M. G., Duarte, O. C. M. B., and Pujolle, G. - “Evaluating the Network Performance Management based on Mobile Agents”, Second International Workshop on Mobile Agents for Telecommunication Applications MATA’00, Lectures Notes in Computer Science 1931, Eric Horlait (Ed.), pp. 95-102, ISSN 0302-9743, Springer-Verlag, September 2000. English, A4 size, 8 p., file: RDP00b.ps.gz

GTA-00-12 Sztajnberg, A. and Loques, O. G. - ”Reflection in the R-RIO Configuration Programming Environment”, IEEE Middleware’2000, Workshop on Reflective Middleware, Palisades, NY, EUA, April, 2000. English, A4 size, 2 p., pp.21-22, file: SzLo00a.ps.gz

GTA-00-13 Loques, O., Sztajnberg, A., Leite, J., and Lobosco, M. - “On the Integration of Configuration and Meta-Level Programming Approaches”, Walter Cazzola, Robert J. Stroud, and Francesco Tisato Editors, Object Oriented Reflection in Software Engineering, Lecture Notes on Computer Science, Springer Verlag, to be launched in ECOOP’2000, France, June, 2000. English, A4 size, 20 p., file: LSLL00.ps.gz

GTA-00-14 Sztajnberg, A. and Loques, O. G. - “Bringing QoS Specifications to the Architectural Level”, ECOOP’2000 Workshop on QoS for Distributed Object Systems, June, France, 2000. (accepted for publication) English, A4 size, 1 p., file: SzLo00b.ps.gz

GTA-00-15 Ziviani, A., de Rezende, J. F., and Duarte, O. C. M. B. - “Evaluating the Expedited Forwarding of Voice Traffic in a Differentiated Services Network”, International Journal of Communication Systems, John Wiley & Sons, vol. 15, no. 9, pp. 799-813, ISSN: 1074-5351, November 2002. English, A4 size, 16 p., file: ZRD00b.ps.gz

GTA-00-16 Delfino, G. M., Santos F^{∘}, J. V., Roman, R. C., Pedroza, A. C. P., and Le? J. L. S. - “Especifica? em Estelle Tempo-Real de um Sistema de Transmiss?CBR”, June 2000. Portuguese, A4 size, 17 p., file: DSR00b.ps.gz

GTA-00-17 Costa, L. H. M. K., Fdida, S., and Duarte, O. C. M. B. - “Un Algorithme R?stant au Facteur d’?helle pour le Routage avec Trois M?iques de QoS”, In Colloque Francophone sur l’Ing?erie des Protocoles - CFIP’2000, pp. 233-247, Toulouse, France, October 2000. French, A4 size, 15 p., file: KFD00a.ps.gz

GTA-00-18 Vidal, P. C. S. and Duarte, O. C. M. B. - “Agrega? de Estados do Protocolo RSVP para Fluxos Multicast”, XVIII Simp?? Brasileiro de Telecomunica?s SBT’2000, Gramado, RGS, September 2000. Portuguese, A4 size, 6 p., file: ViDu00.ps.gz

GTA-00-19 Graduation Final Project Authors - Guilherme Paulo Teixeira dos Santos and Saulo Vaz de Vasconcellos Title - “Avalia? por Simula? dos Mecanismos de Controle de Congestionamento do TCP” Advisor - Jos?erreira de Rezende DEL/EE/UFRJ - August 2000 Portuguese - A4 size - 110 p. - Directory: Vasc00

GTA-00-20 Gon?ves, P. A. da S., de Rezende, J. F., and Duarte, O. C. M. B. - “An Active Service for Multicast Video Distribution”, Journal of the Brazilian Computer Society, Vol. 7, n^{∘} 1, pp. 43-51, ISSN 0104-6500, Brazilian Computer Society. English, A4 size, 9 p., file: GRD00a.ps.gz

GTA-00-21 Gon?ves, P. A. da S., de Rezende, J. F., and Duarte, O. C. M. B. - “Improving Layered Video Multicast using Active Networks”, October 2000. English, A4 size, 6 p., file: GRD00b.ps.gz

GTA-00-22 Delfino, G. M. D., Santos F^{∘}, J. V., Roman, R. C., Le? J. L. S., and Pedroza, A. C. P. - “An Integrated Toolset for the Design of Protocols with QoS Requirements (ProtCAD/RT)”, Protocols for Multimedia Systems - PROMS’2000 Conference, Krakow, Poland, October 2000. English, A4 size, 7 p., file: DSR00c.ps.gz

GTA-00-23 Costa, L. H. M. K., Fdida, S., and Duarte, O. C. M. B. - “Hop-by-hop Multicast”, Journées Doctorales Informatique et Réseaux - JDIR’2000, pp. 85-96, Paris, France - November 2000. French, A4 size, 12p., file: KFD00b.ps.gz

GTA-00-24 Sztajnberg, A. and Loques, O. - “R-RIO: reflective-Reconfigurable Interconnectable Objects”, ACM OOPSLA’00 Companion, pp.85-86, (extended abstract), Minneapolis, USA, October, 2000. English, Letter, 2p. file: SzLo00c.ps.gz

GTA-00-25 Doctor Science Thesis Proposal Author - Marcio Nunes de Miranda Title - “Uma Metodologia de Hardware/Software Codesign de Protocolos de Comunicação Baseada na Otimização de Desempenho por Algoritmos Genéticos" Advisors - Aloysio de C. P. Pedroza and A. C. Mesquita Filho COPPE/PEE/UFRJ - November 2000 Portuguese - A4 size - 67 p. - File: Mira00.ps.gz

GTA-00-26 Ziviani, A., de Rezende, J. F., Duarte, O. C. M. B., and Fdida, S. - “Improving the Delivery Quality of MPEG Video Streams by Using Differentiated Services”, 2nd European Conference on Universal Multiservice Networks – ECUMN’02, pp. 107-115, Colmar, France, April 2002. English, A4 size, 9 p., file: ZRDF00.ps.gz

GTA-00-27 Costa, L. H. M. K., Fdida, S., and Duarte, O. C. M. B. - “A Scalable Algorithm for Link-state QoS-based Routing with Three Metrics”, IEEE International Conference on Communications - ICC’2001, Helsink, Filand, June 2001. English - A4 size - 5 p., file: KFD00c.ps.gz

GTA-00-28 Master Science Thesis Author - Mauros de Campello Queiroz Title - “Um M?do de Acesso ao Meio com Garantia de Qualidade de Servi?para Modem a Cabo” Advisors - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - December 2000 Portuguese - A4 size - 82 p. - Directory: Mauros00

GTA-00-29 Amorim, M. D., Achir, N., Duarte, O. C. M. B., and Pujolle, G. - “Gestion Active de la Vid?dans un Environnement Multisource”, Journ? Doctorales Informatique et R?aux - JDIR’2000, pp. 75-84, Paris, France, November 2000. French, A4 size, 10 p., file: AADP00a.ps.gz GTA-99-01 Lima, R. N. B., Carli, E., Pedroza, A. C. P., Pirmez, L., and Mesquita, A. C.

- “Design for Hardware Implementation of Communication Protocols from a Description Technique”. Portuguese, A4 size, 20 p., file: LCP99a.ps.gz

GTA-99-02 Delfino, G. M., Roman, R. C., Fernadez, M. P., Lima, R. N. B., Mesquita, A. C., and Pedroza, A. C. P. - “ProtCAD - Uma Ferramenta Integrada para Desenvolvimento de Protocolos”. Portuguese, A4 size, 16 p., file: DRF99.ps.gz

GTA-99-03 Teixeira, R. C. and Duarte, O. C. M. B. - “Uma Plataforma para Sistemas Distribu?os de Realidade Virtual”, XXVI Semin?o Integrado de Software e Hardware - SEM-ISH’99, pp. 57-71, Rio de Janeiro, RJ, July 1999. Portuguese, A4 size, 15 p., file: TeDu99a.ps.gz

GTA-99-04 Ziviani, A., de Rezende, J. F., and Duarte, O. C. M. B. - “Tr?go de Voz em um Ambiente de Diferencia? de Servi? na Internet”, XVII Simp?? Brasileiro de Redes de Computadores SBRC’99, pp. 354-367, Salvador, BA, May 1999. Portuguese, A4 size, 14 p., file: ZRD99a.ps.gz

GTA-99-05 Lima, R. N. B., Carli, E., Pedroza, A. C. P., Pirmez, L., and Mesquita, A. C. - “Logic and High Level Synthesis for Communication Protocols”, SBCCI’99, Natal, RN, Brazil. English, A4 size, 4 p., file: LCP99b.ps.gz

GTA-99-06 Rubinstein, M. G. and Duarte, O. C. M. B. - “An?se da Efici?ia de Agentes M??s no Gerenciamento de Redes”, XIX Congresso Nacional da Sociedade Brasileira de Computa?, XXVI Semin?o Integrado de Software e Hardware - SEMISH’99, vol. 1, pp. 167-178, Rio de Janeiro, RJ, Brazil, July 1999. Portuguese, A4 size, 12 p., file: RuDu99.ps.gz

GTA-99-07 de Lima, R. M., Junqueira, F. P., Gon?ves, P. A. da S., and Duarte, O. C. M. B. - “SAMM: An Integrated Environment to Support Multimedia Synchronization of Pre-orchestrated Data”, IEEE International Conference on Multimedia Computing and Systems - ICMCS’99, Vol. 2, pp. 929-933i, Vol. 2, pp. 929-933, Firenze, Italy, June 7-11 1999. English, letter size, 5 p., file: LJGD99.ps.gz

GTA-99-08 Teixeira, R. C. and Duarte, O. C. M. B. - “Comunica?s Multidestinat?as em Ambientes Virtuais Distribu?os Escal?is”. Portuguese, A4 size, 13p., file: TeDu99b.ps.gz

GTA-99-09 Bernardo Filho, O., Pedroza, A. C. P., and Le? J. L. S. - “Uma Ferramenta para Verifica? de Protocolos de Comunica?: Implementa? e Experi?ias”. Portuguese, A4 size, 16 p., file: BPL99.ps.gz

GTA-99-10 Villela, D. A. M. and Duarte, O. C. M. B. - “Improving Scalability on Reliable Multicast Communications”, Computer Communications, ISSN 0140-3664, vol. 24, no. 5-6, pp. 548-562, Elsevier Science, March 2001. English, letter size, 25 p., file: ViDu99.ps.gz GTA-99-11 Teixeira, R. C. and Duarte, O. C. M. B. - “Multicast Communication for Scalable

Distributed Virtual Environments”. English, A4 size, 17 p., file: TeDu99c.ps.gz

GTA-99-12 Ziviani, A., de Rezende, J. F., and Duarte, O. C. M. B. - “Towards a Differentiated Services Support for Voice Traffic”, IEEE Global Telecommunications Conference - Globecom’99, vol. 1A, pp. 59-63, Rio de Janeiro, RJ, Brazil, December 1999. English, Letter size, 5 p., file: ZRD99b.ps.gz

GTA-99-13 Rubinstein, M. G. and Duarte, O. C. M. B. - “Evaluating the Performance of Mobile Agents in Network Management”, IEEE Global Telecommunications Conference - Globecom’99, pp. 386-390, Rio de Janeiro, RJ, Brazil, December 1999. English, Letter size, 5 p., file: RuDu99a.ps.gz

GTA-99-14 Costa, L. H. M. K., Fdida, S., and Duarte, O. C. M. B. - “Distance-vector QoSbased Routing with Three Metrics”, Broadband Communications, High Performance Networking, and Performance of Communication Networks - Networking 2000, Lectures Notes in Computer Sciences n^{∘} 1815, Edited by G. Pujolle, H. Perros, S. Fdida, U. Körner, and I. Stavrakakis, pp. 847-858, ISSN 0302-9743, Springer, 2000.// English, A4 size, 12 p., file: KFD99a.ps.gz

GTA-99-15 Junqueira, F. P. and Duarte, O. C. M. B. - “Comunica? Multiponto para Esta?s M??s em Redes Ativas”, XVII Simp?? Brasileiro de Redes de Computadores SBRC’99, pp. 389-404, Salvador, BA, May 1999. Portuguese, A4 size, 16 p., file: JuDu99a.ps.gz

GTA-99-16 Junqueira, F. P. and Duarte, O. C. M. B. - “Active Networking Support for Mobile Multicast Communication”, February, 1999. English, A4 size, 19 p., file: JuDu99b.ps.gz

GTA-99-17 Junqueira, F. P. and Duarte, O. C. M. B. - “Comunica? Multiponto Confi?l em Redes Ativas Utilizando Cache Adaptativo”, XIX Congresso Nacional da Sociedade Brasileira de Computa? SBC’99, XXVI Semin?o Integrado de Software e Hardware - SEMISH’99, vol. 1, pp. 73-87, Rio de Janeiro, RJ, July 1999. Portuguese, A4 size, 15 p., file: JuDu99c.ps.gz

GTA-99-18 Ziviani, A., de Rezende, J. F., and Duarte, O. C. M. B. - “An?se de Desempenho de Mecanismos de Diferencia? de Servi? para Tr?gos de Voz na Internet”, XIX Congresso Nacional da Sociedade Brasileira de Computa? SBC’99, XXVI Semin?o Integrado de Software e Hardware SEMISH’99, vol. 1, pp. 23 - 37, July 1999. Portuguese, A5 size, 15 p., file: ZRD99c.ps.gz

GTA-99-19 Oliveira, F. S. G. and Pedroza, A. C. P. - “A Arquitetura AMoS - Uma Nova Abordagem para o Provimento de Seguran?em Ambientes com Suporte ?Mobilidade”, March, 1999. Portuguese, A4 size, 18 p., file: OP99a.ps.gz

GTA-99-20 Master Science Thesis Author - Jo?Carlos Mendes Lu? Title - “Implementa? e Medidas de Desempenho de Protocolos de Comunica? Usando STREAMS” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - March 1999 Portuguese - A4 size - 107 p. - Directory: Luis99

GTA-99-21 Amorim, M. D., Duarte, O. C. M. B., and Pujolle, G. - “Single-Loop Packet Merging for Receiver-Oriented Multicast Multi-Layered Video”, International Conference on Computer Communications ICCC’99, pp. 370-376, Tokyo, Japan, September 1999. English, A4 size, 9 p., file:ADP99a.ps.gz

GTA-99-22 Amorim, M. D. and Duarte, O. C. M. B. - “A Self-extracting Accurate Modeling for Bounded-delay Video Services”, Computer Communications, vol. 27, no. 3, pp. 197-207, ISSN 0140-3664, Elsevier Science, 2004. English, A4 size, 11 p., file:AmDu04a.pdf

GTA-99-23 Amorim, M. D., Duarte, O. C. M. B., and Pujolle, G. - “Improving User’s Satisfaction in Adaptive Multicast Video”, Journal of Telecommunication and Networks, ISSN 1229-2370, vol. 4, n. 3, pp. 221-229, Seul, Corea, 2002. English, A4 size, 18 p., file:ADP99b.ps.gz

GTA-99-24 Teixeira, R. C. and Duarte, O. C. M. B. - “Evaluating the Impact of the Communication System on Distributed Virtual Environments”, Multimedia Tools and Applications, vol. 19, no. 3, pp. 259-278, ISSN 1380-7501, Kluwer Academic Publishers, 2003. English, A4 size, 22 p., file:TeDu99d.ps.gz

GTA-99-25 Junqueira, F. P. and Duarte, O. C. M. B. - “Active Mobile Multicast”, April, 1999. English

GTA-99-26 Rubinstein, M. G. and Duarte, O. C. M. B. - “Analyzing Mobile Agent Scalability in Network Management”, IEEE Latin American Network Operations and Management Symposium - LANOMS’99, pp. 64-74, Rio de Janeiro, RJ, Brazil, December 1999. English, A4 size, 11p., file: RuDu99b.ps.gz

GTA-99-27 de Rezende, J. F. - “Avalia? do Servi?Assegurado para a Diferencia? de Servi? na Internet ”, XVII Simp?? Brasileiro de Redes de Computadores SBRC’99, pp. 339-353, Salvador, BA, May 1999. Portuguese, A4 size, 15 p., file: Reze99a.ps.gz

GTA-99-28 Master Science Thesis Author - Jorge Lu? Machado Amaral Title - “Programa? em L??a Difusa: Um Sistema Formal e um Interpretador” Advisor - Jorge Lopes de Souza Le? COPPE/PEE/UFRJ - March 1999 Portuguese - A4 size - 304 p. - File: Amaral99.tgz

GTA-99-29 de Rezende, J. F. - “Assured Service Evaluation ”, IEEE Global Telecommunications Conference - Globecom’99, pp. 100-104, Rio de Janeiro, RJ, Brazil, December 1999. English, Letter size, 5 p., file: Reze99b.ps.gz

GTA-99-30 Costa, L. H. M. K., Fdida, S., and Duarte, O. C. M. B. - “A QoS Routing Algorithm based on Three Different Metrics” - Technical Report, January 1999. English, A4 size, 24 p., file: KFD99b.ps.gz

GTA-99-31 Rubinstein, M. G. and Duarte, O. C. M. B. - “Evaluating Tradeoffs of Mobile Agents in Network Management”, Networking and Information Systems Journal, Hermes Science Publications, vol. 2, no. 2, pp. 237-252, ISSN 1290-2926, 1999. English, A4 size, 16 p., file: RuDu99c.ps.gz

GTA-99-32 Costa, L. H. M. K., Fdida, S., and Duarte, O. C. M. B. - “An Introduction to Virtual Private Networks: Towards D-VPNs”, Networking and Information Systems Journal, Hermes Science Publications, vol. 2, no. 6, ISSN 1290-2926, 2000. English, A4 size, 20 p., file: CFD99.ps.gz

GTA-99-33 Lima, R. N. B., Carli, E., Pedroza, A. C. P., Pirmez, L., and Mesquita, A. C. - “Communication Protocols in Integrated Circuits: Experiments and Results”, IEEE Global Telecommunications Conference - Globecom’99, pp. 178-182, Rio de Janeiro, RJ, Brazil, December 1999. English, Letter size, 6 p., file: LCP99c.ps.gz

GTA-99-34 Master Science Thesis Author - Renata Cruz Teixeira Title - “Uma Plataforma Escal?l para Sistemas Distribu?os de Realidade Virtual” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - July 1999 Portuguese - A4 size - 83 p. - File: Teix99.ps.gz

GTA-99-35 Master Science Thesis Author - Flavio Paiva Junqueira Title - “Comunica? Multiponto Confi?l para Esta?s M??s em Redes Ativas” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - August 1999 Portuguese - A4 size - 99 p. - File: Junqueira99.ps.gz

GTA-99-36 Master Science Thesis Author - Silvia Mara da Costa Campos Victer Title - “Contribui?s para um Sistema de Gerenciamento de Workflow Transacional para WWW” Advisor - Jorge Lopes de Souza Le? COPPE/PEE/UFRJ - June 1999 Portuguese - A4 size

GTA-99-37 Master Science Thesis Author - Frederico Sauer Guimar? Title - “Uma Arquitetura para Suporte a Mobilidade e Seguran?na Internet” Advisor - Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - March 1999 Portuguese - A4 size

GTA-99-38 Doctor Science Thesis Author - Orlando Bernardo Filho Title - “Verifica? de Protocolos de Comunica? com L??a Nebulosa” Advisor - Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - April 1999 Portuguese - A4 size

GTA-99-39 Loques, O., Leite, J., Sztajnberg, A., and Lobosco, M., “Towards Integrating Meta-Level Programming and Configuration Programming”, XIII Simp?? Brasileiro de Engenharia de Software 99, pp.341-354, Florian??is, October 1999. English, A4 size, 13 p., file: LLLS99a.ps.gz GTA-99-40 Sztajnberg, A., Lobosco, M., and Loques, O. G., “Configurando protocolos de

intera? na abordagem R-RIO”, XIII Simp?? Brasileiro de Engenharia de Software 99, pp. 19-45 Florian??is, October 1999. Portuguese, A4 size, 16 p., file: SLL99b.ps.gz

GTA-99-41 Lobosco, M. Loques, O., and Sztajnberg, A."R-RIO: Um Ambiente para Suporte ?Constru? e ?Evolu? de Sistemas”, XIII Simp?? Brasileiro de Engenharia de Software 99 (Caderno de Ferramentas), pp. 53-56, Florian??is, October 1999. Portuguese, A4 size, 4 p.

GTA-99-42 Sztajnberg, A., Lobosco, M., and Loques, O. G., “Separa? de Interesses em Sistemas Distribu?os”, (published as Poster). XVII Simp?? Brasileiro de Redes de Computadores SBRC’99, pp. 339-353, Salvador, BA, May 1999. Portuguese, A4 size, 18 p., file (full version): SLL99d.ps.gz

GTA-99-43 Doctor Science Thesis Proposal Author - Alexandre Sztajnberg Title - “Flexibilidade e Separa? de Interesses para a Concep? e Evolu? de Aplica?s Distribu?as" Advisor - Jorge Lopes de Souza Le? Co-advisor - Orlando Loques (CAA/UFF) COPPE/PEE/UFRJ - July 1999 Portuguese - A4 size - 194 p. - File: Sztaj99.ps.gz

GTA-99-44 Doctor Science Thesis Proposal Author - Marcelo Gon?ves Rubinstein Title - “Avalia? do Desempenho de Agentes M??s no Gerenciamento de Redes” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - August 1999 Portuguese - A4 size - 55 p. - File: Rubi99/Rubi99.ps.gz

GTA-99-45 Lima, R. N. B. Carli, E., Pedroza, A. C. P., Pirmez, L., and Mesquita, A. C. - “Protocolos de Comunica? em Circuitos Integrados: Experiências e Resultados”, XVII Simp?? Brasileiro de Telecomunicações, SBT’99, Vila Velha, ES. Portuguese, A4 size, 6 p., file: LCP99d.ps.gz

GTA-99-46 Master Science Thesis Author - Artur Ziviani Title - “Voz sobre Serviços Diferenciados na Internet” Advisors - Otto Carlos Muniz Bandeira Duarte and José Ferreira de Rezende COPPE/PEE/UFRJ - September 1999 Portuguese - A4 size - 91 p. - Directory: Ziviani99

GTA-99-47 Rubinstein, M. G. and Duarte, O. C. M. B. - “Reducing Response Time in Network Management Tasks by Using Mobile Agent Strategies”, October 1999. English, A4 size, 13 p., file: RuDu99d.ps.gz

GTA-99-48 Ziviani, A., de Rezende, J. F., Duarte, O. C. M. B., and Fdida, S. - “Evaluating Voice Traffic in a Differentiated Services Environment”, 17^{th} International Teletraffic Congress, Salvador, BA, September 2001. English, A4 size, 12 p., file: ZRDF01.ps.gz GTA-99-49 Rubinstein, M. G., Duarte, O. C. M. B., and Pujolle, G. - “Improving Management

Performance by Using Multiple Mobile Agents”, ACM Autonomous Agents 2000, ACM ISBN 1-58113-230-1, Edited by Carles Sierra, Maria Gini, and Jeffrei S. Rosenschein, pp. 165-166, ACM Press, 2000. English, A4 size, 2 p., file: RDP99.ps.gz

GTA-99-50 Loques, O., Leite, J., Sztajnberg, A., and Lobosco, M., “Integrating Meta-Level Programming and Configuration Programming”, OOPSLA’99 Workshop on Reflection and Software Engineering, pp.137-152, Denver, USA, November 1999. English, A4 size, 16 p., file: LLLS99e.ps.gz

GTA-99-51 Fernandez, M. P., Delfino, G. M., and Pedroza, A. C. P. - “Hierarchical Management Protocol for Large Networks”, LANOMS’99, Rio de Janeiro, RJ, Brazil, December 1999. English, A4 size, 10 p., file: FPD99.ps.gz

GTA-99-52 Graduation Final Project Author - Gardel Moreira Delfino Title - “Verifica?, Implementa? e Testes de um Protocolo de Gerenciamento Hier?uico (SGH)” Advisors - Aloysio de Castro Pinto Pedroza and Marcial Porto Fernandez DEL/EE/UFRJ - September 1999 Portuguese - A4 size - 177 p. - Directory: Delfino99 GTA-98-01 Rubinstein, M. G. and Duarte, O. C. M. B. - “Infra-estrutura GTA/AG: Um

Ambiente de Desenvolvimento de Sistemas Baseados em Agentes”, January 1998. Portuguese, A4 size, 11 p., file: RuDu98.ps.gz

GTA-98-02 Villela, D. A. M. and Duarte, O. C. M. B. - “Mecanismos com Prioridades para Escalabilidade e Alto Desempenho em Comunica?s Multiponto Confi?is”, XVI Simp?? Brasileiro de Redes de Computadores SBRC’98, pp. 301-318, Rio de Janeiro, RJ, May 1998. Portuguese, A4 size, 18 p., file: AnDu98.ps.gz

GTA-98-03 Junqueira, F. P. and Duarte, O. C. M. B. - “Comunicação Multiponto Móvel baseada em Redes Ativas”. Portuguese, A4 size, 15 p., file: JuDu98a.ps.gz

GTA-98-04 Villela, D. A. M. and Duarte, O. C. M. B. - “A Priority Scheme Applied to Scalable Reliable Multicast Communications”, IEEE Global Telecommunications Conference GLOBECOM’98, pp. 3368-3373, Sydney, Australia, November 1998. English, Letter size, 17 p., file: AnDu98b.ps.gz

GTA-98-05 Costa, L. H. M. K. and Duarte, O. C. M. B. - “Um Mecanismo Escal?l de Roteamento com Qualidade de Servi?’, XVI Simp?? Brasileiro de Redes de Computadores SBRC’98, pp. 355, Rio de Janeiro, RJ, May 1998. Portuguese, A4 size, Extended Abstract, 1 p., file: KoDu98.ps.gz

GTA-98-06 Oliveira, A. N., Bahiense, A. C. M., Nunes, A. C. F., and Le? J. L. S. - “W4MS Um Sistema Distribu?o de Gerenciamento de Workflow para a WWW”. Portuguese, A4 size, 12 p., file: OBNL98.ps.gz

GTA-98-07 Oliveira, F. S. G. and Pedroza, A. C. P. - “ Uma Arquitetura para Suporte a Mobilidade e Seguran?sobre IPv6”, XVI Simp?? Brasileiro de Redes de Computadores SBRC’98, pp. 596-612, Rio de Janeiro, RJ, May 1998. Portuguese, A4 size, 17 p., file: OP98a.ps.gz

GTA-98-08 Junqueira, F. and Duarte, O. M. B. - “Extens??do Modelo de Comunica? MultiMob-RA para Multim?ia”. Portuguese, A4 size, 12 p., file: JuDu98b.ps.gz

GTA-98-09 Rubinstein, M. G. and Duarte, O. C. M. B. - “Um Mecanismo de Sincroniza? Interm?ia com Taxa M?ima de Realimenta?”, February 1998. Portuguese, A4 size, 12 p., file: RuDu98a.ps.gz

GTA-98-10 de Lima, R. M., Junqueira, F. P., Gon?ves, P. A. da S., and Duarte, O. C. M. B. - “O Ambiente Integrado SAMM de Suporte ? Sincroniza? de Dados Multim?ias Pr?rquestrados”. Portuguese, A4 size, 18 p., file: MJGD98.ps.gz

GTA-98-11 de Lima, R. M. and Duarte, O. C. M. B. - “Um Sistema de Apresenta? Sincronizada de Objetos Multim?ias Pr?rquestrados”, IV Simp?? Brasileiro de Sistemas Multim?ia e Hiperm?ia, pp. 237-248, Rio de Janeiro, RJ, May 1998. Portuguese, A4 size, 12 p., file: MuDu98.ps.gz GTA-98-12 Albuquerque, C. V. N., Faerman, M., and Duarte, O. C. M. B. - “Implementa-

tions of Traffic Control Mechanisms for High Speed Networks”, SBT/IEEE International Telecommunications Symposium ITS’98, pp. 177-182, S?Paulo, Brasil, August 1998. English, Letter size, 6 p., file: AFD.ps.gz

GTA-98-13 Rubinstein, M. G. and Duarte, O. C. M. B. - “Service Location for Mobile Agent Systems”, SBT/IEEE International Telecommunications Symposium ITS’98, pp. 623-626, S?Paulo, SP, Brazil, August 1998. English, Letter size, 4 p., file: RuDu98b.ps.gz

GTA-98-14 de Lima, R. M. and Duarte, O. C. M. B. - “Multimedia Synchronization System for Pre-orchestrated Data Communications”, SBT/IEEE International Telecommunications Symposium ITS’98, pp. 267-272, S?Paulo, Brazil, August 1998. English, Letter size, 6 p., file: MuDu98a.ps.gz

GTA-98-15 Vidal, M. T. L. and Duarte, O. C. M. B. - “Envio de tokens em avan? otimizando a utiliza? de enlaces com alto retardo de propaga?”, XVI Simp?? Brasileiro de Redes de Computadores, pp. 339-352, Rio de Janeiro, RJ, May 1998. Portuguese, A4 size, 14 p., file: ViDu98.ps.gz

GTA-98-16 Amorim, M. D. and Duarte, O. C. M. B. - “XGOP-B: Um Modelo Estendido com Previs?de Ponto de Quebra para Tr?gos de V?eo MPEG”, XVI Simp?? Brasileiro de Redes de Computadores SBRC’98, pp. 723-741, Rio de Janeiro, RJ, May 1998. Portuguese, A4 size, 19 p., file: AmDu98a.ps.gz

GTA-98-17 Lima, R. N. B., Carli, E., Pedroza, A. C. P., Pirmez, L., and Mesquita, A. C. - “S?tese de Alto N?el de Protocolos para a Abordagem IP sobre ATM”. Portuguese, A4 size, 16 p., file: LCP98a.ps.gz

GTA-98-18 Amorim, M. D. and Duarte, O. C. M. B. - “A Novel Deterministic Traffic Model Based on a Two-level Analysis of MPEG Video Sources”, IEEE Global Telecommunications Conference GLOBECOM’98, pp. 244-249, Sydney, Australia, November 1998. English, A4 size, 15 p., file: AmDu98b.ps.gz

GTA-98-19 Carli, E., Lima, R. N. B., Pedroza, A. C. P., Pirmez, L., and Mesquita, A. C. - “Hardware Implementation of Protocols for High Speed Networks from a Formal Description Technique”. English, A4 size, 9 p., file: CLP98a.ps.gz

GTA-98-20 Amorim, M. D. and Duarte, O. C. M. B. - “Um Modelo Estendido para Caracteriza? Determin?tica de Tr?gos de V?eo MPEG”. Portuguese, A4 size, 20 p., file: AmDu98c.ps.gz

GTA-98-21 Lima, R. N. B., Carli, E., Pedroza, A. C. P., Pirmez, L., and Mesquita, A. C. - “High Level Synthesis of IPv6 and AAL5 Protocols”. English, A4 size, 4 p., file: LCP98b.ps.gz

GTA-98-22 Master Science Thesis Author - Marcelo Dias de Amorim Title - “Modelagem Determin?tica de Tr?gos de V?eo Comprimido no Padr?MPEG” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - March 1998 Portuguese - A4 size - 89 p. - Directory: Amorim98

GTA-98-23 Master Science Thesis Author - Marcial Porto Fernandez Title - “Protocolo para Gerenciamento Hier?uico de Redes de Telecomunicaç??’ Advisor - Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - January 1998 Portuguese - A4 size - 166 p. - Directory: Marcial98

GTA-98-24 Junqueira, F. P., Teixeira, R. C., and Duarte, O. C. M. B. - “Multicast Communication for Distributed Virtual Environments based on Active Networks”. Portuguese, A4 size, 6 p., file: JTD98.ps.gz

GTA-98-25 Villela, D. A. M. and Duarte, O. C. M. B. - “Comunica?s Multiponto Confi?is Escalav? atrav?de Transmiss??de M??plas C??s”, XVIII Congresso da Sociedade Brasileira de Computa?, XXV Semin?o Integrado de Software e Hardware - SEMISH’98, pp. 56–73, Belo Horizonte, MG, August 1998. Portuguese, A4 size, 15 p., file: AnDu98c.ps.gz

GTA-98-26 Carli, E., Lima, R. N. B., Pedroza, A. C. P., Pirmez, L., and Mesquita, A. C. - “S?tese L??a do Protocolo IPv6: Resultado de uma Metodologia visando o Projeto de Protocolos em Hardware”. Portuguese, A4 size, 15 p., file: CLP98b.ps.gz

GTA-98-27 Oliveira, F. S. G. and Pedroza, A. C. P. - “Um Suporte Escal?l a Mobilidade e Segurança sobre IPv6”. Portuguese, A4 size, 15 p., file: OP98b.ps.gz

GTA-98-28 Costa, L. H. M. K. and Duarte, O. C. M. B. - “Um Mecanismo de Roteamento para Aplica?s Multim?ias com Tr?Requisitos de Qualidade de Servi?’, XVIII Congresso da Sociedade Brasileira de Computaç? XXV Semin?o Integrado de Software e Hardware - SEMISH’98, pp. 107-124, Belo Horizonte, MG, August 1998. Portuguese, A4 size, 15 p., file: KoDu98a.ps.gz

GTA-98-29 Master Science Thesis Author - Marcelo Sp?dola Bacha Title - “Uma Plataforma de Teste de Fun?s de Gerenciamento para Redes ATM” Advisor - Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - March 1998 Portuguese - A4 size - 170 p. - Directory: Bacha98

GTA-98-30 Graduation Final Project Author - Artur Ziviani Title - “Implementa? de um Ambiente Virtual Distribu?o Baseado em Java e VRML" Advisor - Otto Carlos Muniz Bandeira Duarte DEL/EE/UFRJ - March 1998 Portuguese - A4 size - 128 p. - Directory: Ziviani98

GTA-98-31 Ziviani, A. and Duarte, O. C. M. B. - “Compartilhando Ambientes Virtuais pelo Interfaceamento das Linguagens VRML e Java". Portuguese, A4 size, 12 p., file: ZiDu98.ps.gz GTA-98-32 Amorim, M. D. and Duarte, O. C. M. B. - “Extracting Deterministic Parameters

from MPEG Video Sources with Quality Adjustment”. English, Letter size, 19 p., file: AmDu98e.ps.gz

GTA-98-33 Villela, D. A. M. and Duarte, O. C. M. B. - “Scalability Enhancement of Reliable Multicast Communications by Using Multicopy”. English, Letter size, 25 p.

GTA-98-34 Costa, L. H. M. K. and Duarte, O. C. M. B. - “A Scalable QoS-based Routing Machanism for Supporting Multimedia Applications”, IEEE International Conference on Multimedia Computing and Systems - ICMCS’99, Vol. 2, pp. 347-351, Firenze, Italy, June 7-11 1999. English, A4 size, 5 p., file: KoDu98b.ps.gz

GTA-98-35 Master Science Thesis Author - Daniel Antunes Maciel Villela Title - “Mecanismos para Escalabilidade em Comunica?s Multiponto Confi?is” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - July 1998 Portuguese - A4 size - 89 p. - Directory: Villela98

GTA-98-36 Master Science Thesis Author - Lu? Henrique Maciel Kosmalski Costa Title - “Uma Arquitetura Escal?l de Roteamento com Requisitos de Qualidade de Servi?’ Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - July 1998 Portuguese - A4 size - 90 p. - Directory: Kosmalski98

GTA-98-37 Graduation Final Project Author - Paulo Andr?a Silva Gon?ves Title - “Aplicativo em Tcl/Tk para Composi? de Objetos Multim?ias" Advisors - Otto Carlos Muniz Bandeira Duarte and Ronaldo Mussauer de Lima DEL/EE/UFRJ - March 1998 Portuguese - A4 size - 56 p. - Directory: Paulo98

GTA-98-38 Master Science Thesis Author - Ricardo Neimo Bel?Lima Title - “Projeto de Protocolos Visando a Implementa? em Hardware: Aplica? ao Caso do Protocolo AAL5 para Redes ATM” Advisor - Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - May 1998 Portuguese - A4 size - 130 p. - Directory: Belem98

GTA-98-39 Master Science Thesis Author - Ant?io Carlos Fernandes Nunes Title - “A Estrutura das Aplica?s no Sistema W4 de Gerenciamento de Workflow” Advisor - Jorge Lopes de Souza Le? COPPE/PEE/UFRJ - May 1998 Portuguese - A4 size - 61 p. - Directory: Antonio98

GTA-98-40 de Rezende, J. F. and Romano, S.-P. - “Arquitetura de Integra? de Servi? para a Comuta? IP”, XVI Simp?? Brasileiro de Redes de Computadores SBRC’98, pp. 670-681, Rio de Janeiro, RJ, May 1998. Portuguese, A4 size, 13 p., file: Reze98.ps.gz GTA-98-41 Master Science Thesis

Author - Emerson Carli Title - “Projeto de Protocolos visando a Implementa? em Hardware: aplica? ao Protocolo IPv6 no bito da Abordagem IP over ATM” Advisors - Aloysio de Castro Pinto Pedroza, Antonio Carneiro de Mesquita, and Luci Pirmez COPPE/PEE/UFRJ - April 1998 Portuguese - A4 size - 136 p. - Directory: Emerson98

GTA-98-42 Master Science Thesis Author - Jorge Roberto Mendes Filho Title - “Uma Extens?de um Modelo de Tr?go Autosimilar com Aplica?s em Controle de Admiss?e Tarifa? em Redes ATM” Advisor - Luis Felipe Magalh? de Moraes COPPE/PEE/UFRJ - September 1998

GTA-98-43 Master Science Thesis Author - Guilherme de Melo Baptista Domingues Title - “Modelagem e Desempenho de Arquiteturas de Computadores ATM com Armazenamento nas Sa?as e Diferentes Classes de Tr?go e Prioridades” Advisor - Luis Felipe Magalh? de Moraes COPPE/PEE/UFRJ - December 1998

GTA-98-44 Bernardo Filho, O., Fernandez, M. P., Pedroza, A. C. P., and Le? J. L. S. - “Uma T?ica com L??a Nebulosa para Verifica? de um Protocolo de Ger?ia Hier?uica de Redes”, XVI Simp?? Brasileiro de Redes de Computadores (SBRC’98), May 1998, Rio de Janeiro, RJ.

GTA-98-45 Fernandez, M. P., Delfino, G. M., and Pedroza, A. C. P. - “Implementa?, Verifica? e Teste do Protocolo para Gerenciamento Hier?uico (SGH)”, XVI SBRC, Rio de Janeiro, RJ, May 1998. Portuguese, A4 size, 17 p., file: FPD98a.ps.gz

GTA-98-46 Fernandez, M. P., Delfino, G. M., and Pedroza, A. C. P. - “Protocolo para Gerenciamento Hier?uico de Redes de Telecomunica?s (TMN)”, III Workshop em TMN, XVI SBRC, Rio de Janeiro, RJ, May 1998. Portuguese, A4 size, 11 p., file: FPD98b.ps.gz

GTA-98-47 Bernardo Filho, O., Pedroza, A. C. P., and Le? J. L. S. - “Uma T?ica com L??a Nebulosa para Verifica? de Sistemas Distribu?os”, XII CBA, Uberl?ia, MG, September 1998.

GTA-98-48 Master Science Thesis Author - Gilberto Fernandes Taboada Title - “Um Implementador Semi-Autom?co de Protocolos de Comunica? sobre Sistema Aberto” Advisor - Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - 1998.

GTA-98-49 Graduation Final Project Authors - Leandro Marques Rodrigues and Francisco Eduardo de Luna e Almeida Santos Title - “Especifica? de uma Arquitetura de Gerenciamento TMN Aplicada a Redes ATM”

Advisors - Aloysio de Castro Pinto Pedroza and Marcelo Sp?dola Bacha COPPE/EE/UFRJ - 1998.

GTA-98-50 Graduation Final Project Author - Rodrigo de Carvalho Roman Title - “Projeto Did?co: Sistema de Controle de Estoque Supermercado em Ambiente LINUX” Advisor - Aloysio de Castro Pinto Pedroza DEL/EE/UFRJ - 1998 Potuguese - A4 size - 97 p. - Directory: Roman98

GTA-98-51 Graduation Final Project Author - Denis Litwak Title - “Implementa? de um Protocolo de Negocia? em um Ambiente de Projeto Coletivo” Advisor - Aloysio de Castro Pinto Pedroza DEL/EE/UFRJ - 1998

GTA-98-52 Graduation Final Project Author - Fl?a Lacerda de Freitas Title - “Automa? de uma Mesa de Medida de Se?s de Choque Diferencial Angular para Espalhamento Gama” Advisor - Aloysio de Castro Pinto Pedroza DEL/EE/UFRJ - 1998

GTA-98-53 Lobosco, M., Loques, O. G., and Sztajnberg, A. “Configura? de Sistemas Distribu?os Baseada em Java”, I SICON, Rio de Janeiro, September 1998. Portuguese, A4 size, 6 p., file: LLS98a.ps.gz

GTA-98-54 Sztajnberg, A., Lobosco, M., and Loques, O. G., “Uma Plataforma para Suporte ?Evolu? de Sistemas Distribu?os.”, Workshop de Sistemas Distribu?os, Curitiba, May 1998. Portuguese, A4 size, 8 p., file: SLL98b.ps.gz

GTA-98-55 Sztajnberg, A., Lobosco, M., and Loques, O. G., “Separa? de Interesses em Sistemas Distribu?os”, I SICON, Rio de Janeiro, September 1998. Portuguese, A4 size, 6 p., file: SLL98c.ps.gz GTA-97-01 Rubinstein, M. G. and Duarte, O. C. M. B. - “Implementa? de um Servi?de

Sincroniza? Intra e Interm?ia Considerando Rela?s Temporais Impl?itas”, January 1997. Portuguese, A4 size, 18 p., file: RuDu97.ps.gz

GTA-97-02 de Lima, R. M. and Duarte, O. C. M. B. - “Implementa? de um Servi?de Sincroniza? Multim?ia para Comunica? de Dados Pr?rquestrados”, XV Simp?? Brasileiro de Redes de Computadores SBRC’97, pp. 19-35, S?Carlos, SP, May 1997. Portuguese, A4 size, 17 p., file: MuDu97.ps.gz

GTA-97-03 Graduation Final Project Author - Daniel Antunes Maciel Villela Title - “Implementa? e Medidas de Desempenho de Protocolos com Mecanismo STREAMS” Advisor - Otto Carlos Muniz Bandeira Duarte DEL/EE/UFRJ - February 1997 Portuguese - A4 size - 65 p. - Directory: Daniel97

GTA-97-04 Graduation Final Project Author - Flavio Paiva Junqueira Title - “Gera? Autom?ca de XOCPN para um Servi?de Sincroniza?” Advisors - Otto Carlos Muniz Bandeira Duarte and Ronaldo Mussauer de Lima DEL/EE/UFRJ - February 1997 Portuguese - A4 size - 83 p.

GTA-97-05 Graduation Final Project Author - Renata Cruz Teixeira Title - “Um Servi?de Sincroniza? Interfluxo para Comunica? Multim?ia” Advisors - Otto Carlos Muniz Bandeira Duarte and Marcelo Gon?ves Rubinstein DCC/IM/UFRJ - February 1997 Portuguese - A4 size - 68 p.

GTA-97-06 Amorim, M. D. and Duarte, O. C. M. B. - “Efici?ia de Modelos Determin?ticos de Tr?gos na Representa? de Fontes de V?eo MPEG”, XXIV Semin?o Integrado de Software e Hardware - SEMISH’97, pp. 253-264, August 1997. Portuguese, A4 size, 12 p., file: AmDu97a.ps.gz

GTA-97-07 Amorim, M. D. and Duarte, O. C. M. B. - “Um Novo Modelo Determin?tico para Tr?gos de V?eo MPEG Baseado em Grupo de Quadros”, XV Simp?? Brasileiro de Telecomunica?s - SBT’97, pp. 543-547, September 1997. Portuguese, A4 size, 5 p., file: AmDu97b.ps.gz

GTA-97-08 Graduation Final Project Author - Lu? Henrique Maciel Kosmalski Costa Title - “An?se do Mecanismo Multithread na Implementa? de Protocolos de Comunica?” Advisor - Otto Carlos Muniz Bandeira Duarte DEL/EE/UFRJ - February 1997 Portuguese - A4 size - 66 p. - File: KoGFP.ps.gz

GTA-97-09 Nunes, M. D. and Duarte, O. C. M. B. - “Incorporando Escalabilidade ao Protocolo MCF”, XV Simp?? Brasileiro de Telecomunica?s - SBT’97, pp. 239-243, Recife, PE, September 1997 Portuguese, A4 size, 5p., file: NuDu97.ps.gz GTA-97-10 Villela, D. A. M., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Implementando

Redes Locais em ATM”. Portuguese, A4 size, 34 p., file: VCD97.ps.gz

GTA-97-11 de Lima, R. M., Rubinstein, M. G., and Duarte, O. C. M. B. - “Sincroniza? Multim?ia”, April 1997. Portuguese, A4 size, 34 p., file: MRD97.tar.gz

GTA-97-12 Fernandez, M. P. and Pedroza, A. C. P. - “Especifica? de Protocolo para Gerenciamento Hier?uico em Ambiente TMN”, XV Simp?? Brasileiro de Redes de Computadores, Maio 1997. Portuguese, A4 size, 19 p., file: FePe97a.ps.gz

GTA-97-13 Amorim, M. D. and Duarte, O. C. M. B. - “Um Modelo Determin?tico para Caracteriza? de Tr?gos de V?eo MPEG”, II Semin?o Franco-Brasileiro de Sistemas Inform?cos Distribu?os, pp. 401-412, Fortaleza, CE, November 1997. Portuguese, A4 size, 12 p., file: AmDu97c.ps.gz

GTA-97-14 Costa, L. H. M. K. and Duarte, O. C. M. B. - “Roteamento Baseado em Requisitos de Qualidade de Servi?’. Portuguese, A4 size, 21 p., file: KoDu97.ps.gz

GTA-97-15 Villela, D. A. M. and Duarte, O. C. M. B. - “Novas Pol?icas de Escalonamento para Controle de Admiss?em Redes ATM”.// Portuguese, A4 size, 23 p., file: AnDu97.ps.gz

GTA-97-16 Rubinstein, M. G. and Duarte, O. C. M. B. - “Implementa? de um Servi?de Sincroniza? Baseado em Rela?s Temporais Impl?itas”, Semin?o Franco-Brasileiro de Sistemas Inform?cos Distribu?os, pp. 282-293, Fortaleza, CE, Brazil, November 1997. Portuguese, A4 size, 12 p., file: RuDu97a.ps.gz

GTA-97-17 Rubinstein, M. G. and Duarte, O. C. M. B. - “Utilizando o Paradigma de Programa? Remota em Aplica?s Multim?ias Distribu?as”, July 1997. Portuguese, A4 size, 15 p., file: RuDu97b.ps.gz

GTA-97-18 Abdalla, M. F. and Duarte, O. C. M. B. - “An?se de Mecanismos de CAC para Redes ATM”, XV Simp?? Brasileiro de Telecomunica?s - SBT’97, pp. 150-154, Recife, PE, September 1997. Portuguese, A4 size, 5 p., file: AbDu97.ps.gz

GTA-97-19 Nunes, M. D. and Duarte, O. C. M. B. - “MCF: incorporando escalabilidade, mobilidade e novas estrat?as de implementa? de alto desempenho”, Semin?o Franco-Brasileiro de Sistemas Inform?cos Distribu?os - SFBSID’97, pp. 221-232, Fortaleza, CE, November 1997. Portuguese, A4 size, 14 p., file: NuDu97b.ps.gz

GTA-97-20 Carli, E., Lima, R. N. B., Pedroza, A. C. P., Pirmez, L., and Mesquita, A. C. - “Implementa? em Hardware de Protocolos para Redes de Alta Velocidade a partir de uma Descri? Formal”, II Semin?o Franco-Brasileiro de Sistemas Inform?cos Distribu?os - SFBSID’97, pp. 211-220, Fortaleza, CE, November 1997. Portuguese, A4 size, 11 p., file: CLP97.ps.gz

GTA-97-21 Teixeira, R. C. and Duarte, O. C. M. B. - “Requisitos de Ambientes Virtuais Distribu?os de Larga Escala”, I Workshop de Realidade Virtual, pp. 26-35, S?Carlos, SP, November 1997. Portuguese, A4 size, 10 p., file: TeDu97.ps.gz

GTA-97-22 Junqueira, F. P. and Duarte, O. C. M. B. - “Uma Arquitetura de Implementa? para Redes Ativas”, November 1997. Portuguese, A4 size, 15 p., file: JuDu97.ps.gz

GTA-97-23 Oliveira, F. S. G., Pedroza, A. C. P., and Duarte, O. C. M. B. - “Suporte para Seguran?e Mobilidade no IPv6”, II Semin?o Franco-Brasileiro de Sistemas Inform?cos Distribu?os, pp. 233-244, Fortaleza, CE, November 1997. Portuguese, A4 size, 12 p., file: OPD97.ps.gz

GTA-97-24 Bacha, M. S., Pedroza, A. C. P., Rodrigues, L. M., and Santos, F. E. L. A. - “Uma Plataforma de Testes de Fun?s de Gerenciamento TMN para Redes ATM”, II Semin?o Franco-Brasileiro de Sistemas Inform?cos Distribu?os - SFBSID’97, pp. 335-346, Fortaleza, CE, November 1997. Portuguese, A4 size, 12 p., file: BPRS97.ps.gz

GTA-97-25 Fernandez, M. P. and Pedroza, A. C. P. - “Sistema de Gerenciamento Hier?uico para Grandes Redes de Telecomunica?s”, XV Simp?? Brasileiro de Telecomunica?s - SBT’97, pp. 557-561, Recife, PE, September 1997. Portuguese, A4 size, 5 p., file: FePe97b.ps.gz

GTA-97-26 Junqueira, F. P., Costa, L. H. M. K., and Duarte, O. C. M. B. - “Uma Arquitetura de Roteamento com Qualidade de Servi? Baseada em Redes Ativas”, II Semin?o Franco-Brasileiro em Sistemas Inform?cos Distribu?os - SFBSID’97, pp. 257-268, Fortaleza, CE, November 1997. Portuguese, A4 size, 12 p., file: JuKoDu97.ps.gz

GTA-97-27 Costa, L. H. M. K. and Duarte, O. C. M. B. - “Escalabilidade no Roteamento Baseado em Requisitos de Qualidade de Servi?’. Portuguese, A4 size, 19 p., file: KoDu97b.ps.gz

GTA-97-28 Master Science Thesis Author - Monica de Rezende Title - “Descri? e Implementa? de Apresenta?s Multim?ia Pr?rquestradas Interativas” Advisor - Jorge Lopes de Souza Le? COPPE/PEE/UFRJ - November 1997 Portuguese - A4 size - 136 p. - Directory: MonicaR97

GTA-97-29 Bernardo Filho, O., Pedroza, A. C. P., and Le? J. L. S. - “Uma T?ica com L??a Nebulosa para Verifica? de Protocolos de Comunica?”. Portuguese, A4 size, 20 p., file: BPL97.ps.gz

GTA-97-30 Vidal, M. T. L. and Duarte, O. C. M. B. - “Qualidade de Servi?at? N?el de Aplica?". Portuguese, A4 size, 18 p., file: ViDu97.ps.gz

GTA-97-31 Sobral, J. B. M., Le? J. L. S., Pedroza, A. C. P., and Westphal, C. B. - “Integrating CCS and Z languages for specifying distributed Systems”, First Workshop on Systems Design Languages, San Jose, California, USA, April 1997.

GTA-97-32 Doctor Science Thesis Author - Jorge Duarte Pires Val?o Title - “Aquisi? de Conhecimento Auxiliada por Computador ” Advisor - Jorge Lopes de Souza Le? COPPE/PEE/UFRJ - March 1997

GTA-97-33 Pirmez, L., Pedroza, A. C. P., and Mesquita, A. C. - “S?tese de Alto N?el de Protocolos de Comunica? Especificados na Linguagem Formal Estelle usando o Ambiente AMICAL”, Rev. da Soc. Bras. de Telecomunica?s, vol.12, no. 1, pp. 59 - 70, June 1997.

GTA-97-34 Pirmez, L., Pedroza, A. C. P., and Mesquita, A. C. - “High-Level Synthesis of Protocols Described by a Formal Technique”, VLSI’97, IX IFIP Int. Conf. on VLSI, Gramado, RS, August 1997.

GTA-97-35 Master Science Thesis Author - Marlos de Andrade Martins Title - “Um Ambiente Distribu?o para o Desenvolvimento de Protocolos” Advisor - Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - 1997.

GTA-97-36 Master Science Thesis Author - Ildemar Pinto Nunes Title - “Implementa? Semi-autom?ca de Protocolos em Plataformas UNIX” Advisor - Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - 1997.

GTA-97-37 Graduation Final Project Authors - Lu? Cl?io Spielmann and Tula Kaiser Title - “Projeto de um Sistema de Gerenciamento de Redes sobre uma Plataforma Distribu?a HORB” Advisor - Aloysio de Castro Pinto Pedroza DEL/EE/UFRJ - 1997

GTA-97-38 Graduation Final Project Authors - Eduardo Peixoto da Silva e Ricardo Peixoto da Silva Title - “Projeto Did?co de Protocolo da Camada de Rede” Advisor - Aloysio de Castro Pinto Pedroza DEL/EE/UFRJ - 1997

GTA-97-39 Graduation Final Project Author - Mauro Vallad?de Oliveira Title - “Plano de Recupera? de Desastres para Redes Locais de Computadores” Advisor - Aloysio de Castro Pinto Pedroza DEL/EE/UFRJ - 1997

GTA-97-40 Sztajnberg, A. and Loques, O. G., “Configura? e Reflex?Computacional em Aplica?s Multim?ia Distribu?as”, XV Simp?? Brasileiro de Redes de Computadores, S?Carlos, SP, January 1997. Portuguese, A4 size, 18 p., file: SzLo97a.ps.gz

GTA-97-41 Doctor Science Thesis Author - Nival Nunes de Almeida Title - “Modularidade e Composicionalidade de Sistemas Concorrentes” Advisor - Jorge Lopes de Souza Le? COPPE/PEE/UFRJ - March 1997 GTA-96-01 de Lima, R. M. and Duarte, O. C. M. B. - “Comunicaç? Sincronizada de Dados

Multim?ias Pr?rquestrados”, XVI Congresso da Sociedade Brasileira de Computaç? XXIII Semin?o Integrado de Software e Hardware SEMISH, pp. 285-296, Recife, PE, August 1996. Portuguese, A4 size, 12 p., file: MuDu96.ps.gz

GTA-96-02 Nunes, M. D. and Duarte, O. C. M. B. - “MCF: Um Protocolo Multidestinat?o Configur?l com Gerenciamento de Acesso por Fichas de Permiss?’, XIV Simp?? Brasileiro de Redes de Computadores SBRC’96, pp. 159-177, Fortaleza, CE, May 1996. Portuguese, A4 size, 19 p., file: NuDu96.ps.gz

GTA-96-03 Rubinstein, M. G. and Duarte, O. C. M. B. - “Definiç? de um Serviço de Sincronizaç?Cont?ua para Aplicaç??Multim?ias”, XIV Simp?? Brasileiro de Redes de Computadores SBRC’96, pp. 38-55, Fortaleza, CE, Brazil, May 1996. Portuguese, A4 size, 18 p., file: RuDu96.ps.gz

GTA-96-04 Taboada, G. F. and Pedroza, A. C. P. - “Um implementador semi-autom?co de protocolos descritos em Estelle para ambiente UNIX”, January 1996. Portuguese, A4 size, 19 p., File: TaPe96.ps.gz

GTA-96-05 Bacha, M. S. and Pedroza, A. C. P. - “Especificaç? Formal de uma Plataforma de Gerenciamento TMN para Redes ATM”, XIV Simp?? Brasileiro de Redes de Computadores SBRC’96, pp. 273-291, Fortaleza, CE, May 1996. Portuguese, A4 size, 19 p., File: BaPe96.ps.gz

GTA-96-06 Graduation Final Project Author - Hermes Alexandre Alcantara Silva Costa Title - “Implementaç?de Protocolo Multidestinat?o no MBONE” Advisor - Otto Carlos Muniz Bandeira Duarte DEL/EE/UFRJ - March 1996 Portuguese - A4 size - 55 p. - Directory: Costa96

GTA-96-07 Graduation Final Project Author - Marcelo Dias de Amorim Title - “Aplicaç?em Tcl/Tk para Protocolo Multidestinat?o” Advisor - Otto Carlos Muniz Bandeira Duarte DEL/EE/UFRJ - March 1996 Portuguese - A4 size - 65 p. - Directory: Amorim96

GTA-96-09 Graduation Final Project Author - Ricardo Neimo Belem Lima Title - “Projeto de Protocolo para Rede de Alta Velocidade” Advisor - Aloysio de Castro Pinto Pedroza DEL/EE/UFRJ - March 1996 Portuguese - A4 size - 73 p. - Directory: Belem96

GTA-96-10 de Lima, H. M. and Duarte, O. C. M. B. - “Um Novo Protocolo Ponto-a-Multiponto de Retransmiss?Seletiva com Previs? de Transbordo de Mem??’, XIV Simp?? Brasileiro de Telecomunicaç??TELEMO’96, pp. 323-328, Curitiba, PR, July 1996. Portuguese, A4 size, 6 p., file: LiDu96.ps.gz GTA-96-11 de Lima, H. M. and Duarte, O. C. M. B. - “Comments on Evaluation of the Mean

Error-Free Interval of a Noisy Channel”, IEEE Transactions on Communications, Vol. 44, n^{∘} 8, pp. 918-920, August 1996. English, Letter size, 2 p., file: LiDu96a.ps.gz

GTA-96-12 Master Science Thesis Author - Ronaldo Mussauer de Lima Title - “Um Serviço de Sincronizaç?Multim?ia de Dados Pr?rquestrados” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - June 1996 Portuguese - A4 size - 94 p. - Directory: Mussauer96

GTA-96-13 de Lima, H. M. and Duarte, O. C. M. B. - “A Point-to-Multipoint SR ARQ Scheme with Accumulative Acknowledgment for Satellite Communications”, Computer Networks and ISDN Systems, vol. 30, no. 14, pp. 1311-1315, August 1998. Portuguese, A4 size, 12 p., file: LiDu96b.ps.gz

GTA-96-14 Master Science Thesis Author - Michel Ferreira Abdalla Title - “An?se de Mecanismos de Controle de Admiss?de Conex?para Redes ATM” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - September 1996 Portuguese - A4 size - 131 p. - Directory: Abdalla96

GTA-96-15 Nunes, M. D. and Duarte, O. C. M. B. - “CTMP: A Configurable Token-based Multicast Protocol”, October 1996. English, A4 size, 19 p., file: NuDu96b.ps.gz

GTA-96-16 de Lima, R. M. and Duarte, O. C. M. B. - “A Multimedia Synchronization Service and Implementation Architecture for Pre-orchestrated Data Communications”, October 1996. English, A4 size, 4 p., file: MuDu96a.ps.gz

GTA-96-17 Master Science Thesis Author - Marcelo Gon?ves Rubinstein Title - “Um Servi?de Sincroniza? Cont?ua para Aplica?s Multim?ias” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - November 1996 Portuguese - A4 size - 119 p. - Directory: Rubi96

GTA-96-18 Doctor Science Thesis Author - Luci Pirmez Title - “Uma Metodologia para a S?tese de Alto N?el de Protocolos a Partir de uma Descriç?Formal” Advisors - Aloysio de Castro Pinto Pedroza and Ant?io Carneiro de Mesquita Filho COPPE/PEE/UFRJ - July 1996 Portuguese - A4 size - 246 p. - Directory: Luci96

GTA-96-19 Doctor Science Thesis Author - Jo?Bosco Mangueira Sobral Title - “Uma Liguagem para Especifica? de Sistemas Distribu?os em Alto N?el de Abstra?” Advisor - Jorge Lopes de Souza Le? COPPE/PEE/UFRJ - March 1996 GTA-96-20 Pirmez, L., Pedroza, A. C., Mesquita, A. C., and Jerraya, A. A. - “Mapeamento

de ESTELLE em VHDL e Estilos de Descri? de Protocolos de Comunica? para S?tese de Alto N?el”, XIV SBRC, Fortaleza, May 1996.

GTA-96-21 Mendes, M. C. and Pedroza, A. C. - “Uma Ferramenta para S?tese de Protocolos em ESTELLE”, XXIII SEMISH/SBC, Recife, August 1996.

GTA-96-22 Pirmez, L., Rahmouni, M., Kission, P., Pedroza, A. C., Mesquita, A. C., and Jerraya, A. A. - “Analysis of Different Protocol Description Styles in VHDL for High-Level Syntesis”, EURO-DAC’96, PALEXPO, Geneva, Switzerland, September 1996.

GTA-96-23 Master Science Thesis Author - F?o Contarini Carneiro Title - “Projeto de um Compilador de Estelle para VHDL” Advisor - Aloysio de Castro Pinto Pedroza IME and COPPE/PESC/UFRJ - 1996.

GTA-96-24 Master Science Thesis Author - Marcelo Ciminelli Mendes Title - “Uma Ferramenta para a S?tese de Protocolos em Estelle” Advisor - Aloysio de Castro Pinto Pedroza IME and COPPE/PESC/UFRJ - 1996.

GTA-95-01 Faerman, M., Albuquerque, C. V. N., and Duarte, O. C. M. B. - “Implementation of Policing Mechanisms for High Speed Networks”, January 1995. English, A4 size, 16 p., file: AFD95b.ps.gz

GTA-95-02 Faerman, M., Albuquerque, C. V. N., and Duarte, O. C. M. B. - “Time Slotted Moving Window - a New Policing Mechanism for High Speed Networks”, January 1995. English, A4 size, 12 p., file: FAD95b.ps.gz

GTA-95-03 de Lima, H. M. and Duarte, O. C. M. B. - “An Effective Selective Repeat ARQ Strategy for High Speed Point-to-Multipoint Communications”, IEEE Global Telecommunications Conference GLOBECOM’96, pp. 1059-1063, Londres, November 1996. English, A4 size, 18 p., file: LiDu95a.ps.gz

GTA-95-04 Graduation Final Project Author - Jo?Carlos Mendes Lu? Title - “Implementaç?de Protocolos de Comunicaç? em Rede usando o Mecanismo de STREAMS do Unix” Advisor - Otto Carlos Muniz Bandeira Duarte DEL/EE/UFRJ - March 1995 Portuguese, A4 size, 79 p., Directory: Luis95

GTA-95-05 Master Science Thesis Author - Miguel Grimm do Nascimento Title - “Uma interface de programaç?de aplicaç?? para sistemas distribuidos" Advisor - Jorge Lopes de Souza Le? COPPE/PEE/UFRJ - April 1995

GTA-95-06 Master Science Thesis Author - Agliberto Pessoa da Silva Title - “Escalamento Heur?tico baseado em L??a Difusa” Advisor - Jorge Lopes de Souza Le? COPPE/PEE/UFRJ - April 1995

GTA-95-07 Master Science Thesis Author - C?o Vinicius Neves de Albuquerque Title - “Arquitetura de Implementaç?de Alto Desempenho para Sistemas de Comunicaç?Multim?ia" Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - April 1995 Portuguese - A4 size - 127 p. - Directory: Albuquerque95

GTA-95-08 Master Science Thesis Author - Marcelo Dias Nunes Title - “Protocolo Mutidestinat?o com Configuraç? de Facilidades e Gerenciamento de Acesso" Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - April 1995 Portuguese - A4 size - 103 p. - Directory: Nunes95

GTA-95-09 Master Science Thesis Author - Rog?o Leone Texeira da Cunha Title - “Camada Rede e Gerenciamento de Mem?? de um Sistema de Comunicaç?Industrial de Alto Desempenho” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - April 1995

GTA-95-10 Sobral, J. B. M., Le? J. L. S., and Pedroza, A. C. P. - “Especificando protocolos em Z com agentes, 13^{∘} Simp?? Brasileiro de Redes de Computadores, pp. 41-60, Belo Horizonte, MG, May 1995. Portuguese

GTA-95-11 de Lima, H. M. and Duarte, O. C. M. B. - “Protocolos Ponto-a-Multiponto com Retrans mis??de M??plas C??s Adaptados ao Ambiente de Sat?te a Alta Velocidade”, 13^{∘} Simp?? Brasileiro de Redes de Computadores SBRC’95, pp. 181-198, Belo Horizonte, MG, May 1995. Portuguese, A4 size, 18 p., file: LiDu95b.ps.gz

GTA-95-12 Albuquerque, C. V. M., Faerman M., and Duarte, O. C. - “Implementaç?de Mecanismos de Policiamento em Redes de Alta Velocidade”, 13^{∘} Simp?? Brasileiro de Redes de Computadores SBRC’95, pp. 103-119, Belo Horizonte, MG, May 1995. Portuguese, A4 size, 17 p., file: AFD95a.ps.gz

GTA-95-13 Vidal, M. T. L., Duarte, O. C., and Pedroza, A. C. P. - “Serviço de Transfer?ia de Dados Multiponto-a-Multiponto ”, 13^{∘} Simp?? Brasileiro de Redes de Computadores SBRC’95, pp. 229-245, Belo Horizonte, MG, May 1995. Portuguese, A4 size, 17 p., file: VDP95.ps.gz

GTA-95-14 Pirmez, L., Pedroza, A. C. P., and Mesquita, A. C. - “Uma Metodologia para a Implementaç?de Sistemas Distribu?os em Hardware a partir de uma Descriç? Formal”, 13^{∘} Simp?? Brasileiro de Redes de Computadores SBRC’95, pp. 463-482, Belo Horizonte, MG, May 1995.

GTA-95-15 Nunes, M. D. and Duarte, O. C. M. B. - “Proposta e Implementaç?de um Protocolo Multidestinat?o Configur?l”, VII Simp?? Brasileiro de Arquitetura de Computadores e Processamento de Alto Desempenho, pp. 229-242, Canela, RS, July 1995. Portuguese, A4 size, 17 p., file: NuDu95.ps.gz

GTA-95-16 Master Science Thesis Author - Marcio Faerman Title - “Monitoraç?e Policiamento de Tr?go em Sistemas de Comunicaç?Multim?ia" Advisor - Otto Carlos M. B. Duarte COPPE/PEE/UFRJ - August 1995 Portuguese - A4 size - 74 p. - Directory: Faerman95

GTA-95-17 Fialho, S. V., Le? J. L. S., and Pedroza, A. C. P. - “A Formal Technique for the Specification and Verification of Distributed Systems and its Application in Manufacturing Automation”, 38th IEEE Midwest International Symposium on Circuit and Systems, pp. 27-30, Rio de Janeiro, RJ, August 1995.

GTA-95-18 de Lima, H. M. and Duarte, O. C. M. B. - “Performance of Repeated Retransmission GB(N) Schemes for Point-to-Multipoint Noisy Channels”, 38th IEEE Midwest International Symposium on Circuit and Systems, Rio de Janeiro, pp. 1-4, RJ, August 1995. English, Letter size, 4 p., file: LiDu95c.ps.gz GTA-95-19 Albuquerque, C. V. N., Nunes, M. D., and Duarte, O. C. M. B. - “An Efficient

Implementation Architecture for Layered Communication Systems”, 38th IEEE Midwest Symposium on Circuit and Systems, Rio de Janeiro, pp. 514-517, RJ, August 1995. English, Letter size, 4 p., file: AND95.ps.gz

GTA-95-20 Faerman, M., Albuquerque, C. V. N., and Duarte, O. C. M. B. - “Janela M?? Particionada: Proposta de um Novo Mecanismo de Policiamento de Tr?go”, 13^{∘} Simp?? Brasileiro de Telecomunicaç?? Águas de Lind?? SP, September 1995. Portuguse, A4 size, 14 p., file: FAD95a.ps.gz

GTA-95-21 Sobral, J. B. M., Le? J. L. S., and Pedroza, A. C. P. - “Especificando Software de Telecomunicaç??em Z com agentes”, 13^{∘} Simp?? Brasileiro de Telecomunicaç?? Águas de Lind?? SP, September 1995. Portuguese

GTA-95-22 Villela, D. A. M., Costa, L. H. M. K., Albuquerque, C. V. N., Nunes, M. D., and Duarte, O. C. M. B. - “Desempenho de uma Aplicaç?de Transfer?ia de Arquivos em um Sistema de Comunicaç?Industrial”, XI Congresso Brasileiro de Autom?ca, pp. 1745–175, S?Paulo, SP, September 1996. Portuguese, A4 size, 17 p., file: VCAN95.ps.gz

GTA-95-23 Pirmez, L., Carneiro, F. C., Pedroza, A. C., and Mesquita, A. C. - “A Methodology to Develop Integrated Circuits from ESTELLE Especifications”, 38th. IEEE/MIDWEST Symp., Rio de Janeiro, August 1995.

GTA-95-24 Pirmez, L., Pedroza, A. C., and Mesquita, A. C. - “A Methodology to the Implementation of Distributed Systems in Hardware from a Formal Description”, IFIP/PSTV’95, Warsaw, Poland, June 1995.

GTA-95-25 Master Science Thesis Author - Francisco Resende Loureiro Title - “Um Simulador para a Linguagem Estelle” Advisor - Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - 1995.

GTA-95-26 Master Science Thesis Author - Emmanuel Gomes Sanches Title - “Plataforma de Software para Desenvolvimento de Algoritmos de Busca” Advisors - Aloysio de Castro Pinto Pedroza and Edmundo A. de Souza e Silva COPPE/PESC/UFRJ - 1995.

GTA-94-01 Graduation Final Project Authors - Guilherme de M. B. Domingues and Marcio Faerman Title - “An?se de Desempenho do Protocolo FTAM” Advisor - Otto Carlos Muniz Bandeira Duarte DEL/EE/UFRJ - February 1994

GTA-94-02 Nunes, M. D. and Duarte, O. C. M. B. - “Protocolos de Alta Desempenho: Um Tutorial”, March 1994. Portuguese, A4 size, 17 p., file: NuDu94a.ps.gz

GTA-94-03 de Lima, H. M. and Duarte, O. C. M. B. - “Um Protocolo com Retransmiss??Multic??s para Comunicaç??via Sat?te a Alta Velocidade”, Revista da Sociedade Brasileira de Telecomunica?s, vol. 14, no. 2, pp. 56-63, December 1999. Portuguese, A4 size, 13 p., file: LiDu94.ps.gz

GTA-94-04 Doctor Science Thesis Author - Heliomar Medeiros de Lima Title - “Protocolos Ponto-a-Multiponto de Alto Desempenho” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - March 1994

GTA-94-05 Master Science Thesis Author - Maria Fernanda Belford Roxo Title - “Primata: Uma Processador Maciçamente Paralela" Advisor - Otto Carlos M. B. Duarte and Adriano Joaquim O. Cruz COPPE/PEE/UFRJ - March 1994 Portuguese - A4 size - 167 p. - Directory: Roxo94

GTA-94-06 Master Science Thesis Author - Marcelo Averbuch Title - “Aplicaç?do M?do de Tableau para L??a Temporal - Verificaç?de Programas Concorrentes" Advisor - Jorge Lopes de Souza Le? COPPE/PEE/UFRJ - April 1994

GTA-94-07 de Lima, H. M. and Duarte, O. C. M. B. - “A Go-Back-N Protocol with Multicopy Retransmission for High Speed Satellite Communications”, IEEE International Conference on Communications SUPERCOMM/ICC’94, pp. 859-863, New Orleans, EUA, May 1994. English, Letter size, 4 p., file: LiDu94a.ps.gz

GTA-94-08 de Lima, H. M. and Duarte, O. C. M. B. - “Um Protocolo SR ARQ Ponto a Multiponto com Reconhecimento Acumulativo para Comunicaç??a Alta Velocidades”, XII Simp?? Brasileiro de Redes de Computadores, pp. 321-339, Curitiba, PR, May 1994. Portuguese, A4 size, 19 p., file: LiDu94b.ps.gz

GTA-94-09 de Lima, H. M. and Duarte, O. C. M. B., “An Improved Point-to-Multipoint GB(N) ARQ Scheme for High Speed Satellite Communications”, IEEE International Symposium on Circuits and Systems ISCAS94, pp. 3.241-3.244, London, UK, May 1994. English, Letter size, 4 p., file: LiDu94c.ps.gz GTA-94-10 Albuquerque, C. V. N., Nunes, M. D., and Duarte, O. C. M. B. - “Performance

Measurements in a Manufacturing Communication System”, IEEE International Conference on Circuits and Systems ISCAS’94, pp. 3.69-3.72, London, England, May 1994. English, A4 size, 15 p., file: AND94.ps.gz

GTA-94-11 Albuquerque, C. V. N. and Duarte, O. C. M. B. - “Algoritmos de Controle de Tr?go para Garantia da Qualidade de Serviço em Aplicaç?? de Tempo Real”, July 1994. Portuguese, A4 size, 15 p., file: AlDu94.ps.gz

GTA-94-12 Nunes, M. D. and Duarte, O. C. M. B. - “An?se de Mecanismos para Protocolos de Alto Desempenho”, VI Simp?? Brasileiro de Arquitetura de Computadores e Processamento de Alto Desempenho, pp. 157-169, Caxambu, MG, August 1994. Portuguese, A4 size, 13 p., file: NuDu94b.ps.gz

GTA-94-13 Roxo, M. F. E. B., Cruz, A. J. O., Duarte, O. C. M. B., and Jorge, R. C. B. - “PRIMATA: Uma Arquitetura Maciçamente Paralela para Processamento de Imagens e Matrizes”, VI Simp?? Brasileiro de Arquitetura de Computadores Processamento de Alto Desempenho (SBAC-PAD), pp. 185-197, Caxambu, MG, August 1994. Portuguese, A4 size, 13 p., file: RCDJ94.ps.gz

GTA-94-14 de Lima, H. M. and Duarte, O. C. M. B. - “A Point-to-Multipoint ARQ Scheme with Multicopy Retransmission for High Speed Satellite Communications”, SBT/IEEE International Telecommunications Simposium ITS’94, pp.357-361, Rio de Janeiro, BR, September 1994. English, Letter size, 5 p., file: LiDu94d.ps.gz

GTA-94-15 Master Science Thesis Author - Marcos Tadeu von Lutzow Vidal Title - “Especificaç?Formal e Implementaç?de um Protocolo de Comunicaç?Multipontoa-Multiponto e Aplicaç?de Texto-confer?ia" Advisors - Otto Carlos M. B. Duarte and Aloysio de C. P. Pedroza COPPE/PEE/UFRJ - November 1994 Portuguese - A4 size - 192 p. - Directory: Vidal94

GTA-94-16 de Lima, H. M. and Duarte, O. C. M. B. - “An Accumulative Acknowledgement Protocol for Point-to-Multipoint High Speed Communications”, IEEE Global Telecommunication Conference GLOBECOM’94, pp. 181-185, San Francisco, USA, November 1994 English, Letter size, 5 p., file: LiDu94e.ps.gz

GTA-94-17 Abdalla, M. F. and Duarte, O. C. M. B. - “Caracteriza? e Modelagem de Tr?go Agregado em Redes ATM”, September 1994 Portuguese, A4 size, 23 p., file: AbDu94.ps.gz

GTA-94-18 Lehman Jr., E. O. and Pedroza, A. C. P. - “Especifica? e Verificac?do Protocolo CMIP para o Gerenciamento de Redes de Computadores”, XII Simp?? Brasileiro de Redes de Computadores, Curitiba, PR, May 1994.

GTA-94-19 Rodrigues, R. L., Pedroza, A. C. P., and Mesquita Filho, A. C.- “S?tese de alto n?el de protocolos de redes descritos em ESTELLE”, IX SBMICRO, Rio de Janeiro, RJ, August 1994.

GTA-94-20 Master Science Thesis Author - Ricardo Land Rodrigues Title - “S?tese de Alto N?el de Protocolos de Redes Especificados na Linguagem Estelle” Advisors - Aloysio de Castro Pinto Pedroza and Ant?io Carneiro de Mesquita Filho COPPE/PEE/UFRJ - 1994.

GTA-94-21 Master Science Thesis Author - Erny Otto Lehmann J??r Title - “Especifica? e Verifica? do Protocolo CMIP para Ger?ia de Redes” Advisor - Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - 1994.

GTA-94-22 Graduation Final Project Author - Gilberto Fernandes Taboada Title - “Verificador de Estelle para UNIX” Advisor - Aloysio de Castro Pinto Pedroza DEL/EE/UFRJ - 1994 GTA-93-01 Graduation Final Project

Author - C?o Vinicius Neves de Albuquerque Title - “Implementaç?do MHS no ambiente UNIX - SUN OS” Advisors - Otto Carlos Muniz Bandeira Duarte and Marcelo Luiz D. Lanza DEL/EE/UFRJ - March 1993

GTA-93-02 Graduation Final Project Author - Marcelo Dias Nunes Title - “Implementaç?do Protocolo X25” Advisors - Otto Carlos Muniz Bandeira Duarte and Marcelo Luiz D. Lanza DEL/EE/UFRJ - March 1993

GTA-93-03 Master Science Thesis Author - Calixto Damian Neto Title - “Implementaç?do Protocolo FTAM” Advisor - Otto Carlos Muniz Bandeira Duarte COPPE/PEE/UFRJ - April 1993 Portuguese - A4 size - 96 p.

GTA-93-04 Vidal, M. T., Barros, F. M. C., and Duarte, O. C. M. B. - “Servi? de Transfer?ia de Dados Ponto a Multiponto”, XI Simp?? Brasileiro de Redes de Computadores, pp. 97-116, Campinas, SP, May 1993. Portuguese, A4 size, 20 p., file: ViDu93.ps.gz

GTA-93-05 de Lima, H. M. and Duarte, O. C. M. B. - “Um Novo Protocolo Multiponto Para Comunica?s Via Sat?te a Altas Velocidades”, XI Simp?? Brasileiro de Redes de Computadores, pp. 616-630, Campinas, SP, Brasil, May 1993. Portuguese, A4 size, 13 p., file: LiDu93.ps.gz

GTA-93-06 Roxo, M. F. E. B., Cruz, A. J. O., and Duarte, O. C. M. B. - “PRIMATA: Desenvolvimento de um Elemento Processador para uma Arquitetura Massivamente Paralela”, V Simp?? Brasileiro de Arquitetura de Computadores Processamento de Alto Desempenho (SBAC-PAD), pp. 150-165, Florian??is, SC, September 1993. Portuguese, A4 size, 16 p., file: RCD93.ps.gz

GTA-93-07 Albuquerque, C. V. N., Lanza, M. L. D., and Duarte, O. C. M. B. - "Sistemas de Comunicaç?Abertos em Redes Locais: Uma Implementaç?e Medidas de Desempenho em Ambientes de Computadores Pessoais”, XIII Congresso da Sociedade Brasileira de Computaç?(SEMISH), pp. 327-339, Florian??is, SC, September 1993. Portuguese, A4 size, 13 p., file: ALD93.ps.gz

GTA-93-08 Graduation Final Project Author - Michel Ferreira Abdalla Title - “Um Ambiente Integrado para An?se de Redes de Petri” Advisors - Marcelo Luiz D. Lanza and Jo?Amaro B. Pereira DEL/EE/UFRJ - June 1993

GTA-93-09 Burrowes, F. B., Novaes, M. G., and Pedroza, A. C. P. - “Protocolo MMS Projeto e Aplica?es”, XI Simp?? Brasileiro de Redes de Computadores, Campinas, SP, May 1993.

GTA-93-10 Fialho, S. V., Courtiat, J. P., Pedroza, A. C. P., and Le? J. L. S. - “A Partial Reachability Graph Generation Technique for the Verification of ESTELLE Specifications”, XI Simp?? Brasileiro de Telecomunica?s, Natal, RN, September 1993.

GTA-93-11 Master Science Thesis Author - Eduardo Compasso Arbex Title - “Controle H?rido de Posi? de For?para um Manipulador Rob??o” Advisor - Jorge Lopes de Souza Le? COPPE/PEE/UFRJ - April 1993.

GTA-93-12 Master Science Thesis Author - Frederico Celestino Pontes Title - “LDHL - Uma Linguagem L??a para Verifica? de Circuitos Digitais” Advisor - Jorge Lopes de Souza Le? COPPE/PEE/UFRJ - October 1993.

GTA-93-13 Master Science Thesis Author - Luis Manuel S?hez y Bernal Title - “Um Compilador para Lotos” Advisor - Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - 1993.

GTA-93-14 Master Science Thesis Author - Helio Shor Title - “Projeto de uma C?la Flex?el de Montagem usando Servi? MMS” Advisor - Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - 1993.

GTA-93-15 Master Science Thesis Author - S?io Campos de Alc?ara Title - “Implementa? Semi-Autom?ca de Protocolos Descritos em Estelle” Advisor - Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - 1993.

GTA-93-16 Master Science Thesis Author - S?io Granato de Ara?? Title - “Estudo, Especifica? e Implementa? de Protocolos para o Plano de Controle da RDSI” Advisor - Aloysio de Castro Pinto Pedroza COPPE/PEE/UFRJ - 1993.